### v1

# Kronecker-product random matrices and a matrix least squares problem

Abstract.
We study the eigenvalue distribution and resolvent of a Kronecker-product random matrix model $A\otimes I_{n\times n}+I_{n\times n}\otimes B+\Theta\otimes\Xi\in\mathbb{C}^{n^{2}\times n^{2}}$, where $A,B$ are independent Wigner matrices and $\Theta,\Xi$ are deterministic and diagonal. For fixed spectral arguments, we establish a quantitative approximation for the Stieltjes transform by that of an approximating free operator, and a diagonal deterministic equivalent approximation for the resolvent. We further obtain sharp estimates in operator norm for the $n\times n$ resolvent blocks, and show that off-diagonal resolvent entries fall on two differing scales of $n^{-1/2}$ and $n^{-1}$ depending on their locations in the Kronecker structure.
Our study is motivated by consideration of a matrix-valued least-squares optimization problem $\min_{X\in\mathbb{R}^{n\times n}}\frac{1}{2}\|XA+BX\|_{F}^{2}+\frac{1}{2}\sum_{ij}\xi_{i}\theta_{j}x_{ij}^{2}$ subject to a linear constraint. For random instances of this problem defined by Wigner inputs $A,B$, our analyses imply an asymptotic characterization of the minimizer $X$ and its associated minimum objective value as $n\to\infty$.
1. Introduction
In recent years, high-dimensional probabilistic analyses have yielded important insights into the exact asymptotic behavior of many optimization problems with random data. We mention, as several examples, analyses of ridge regression [39, 16, 17] with possibly non-linear random features [19, 36, 46] and/or in kernelized domains [45, 65] using asymptotic random matrix theory, and analyses of optimization problems arising in contexts of non-linear regression [6, 21, 63, 18, 20, 62], classification [61, 48, 44, 15], and variational Bayesian inference [31, 58, 11] using Approximate Message Passing algorithms, Gaussian comparison and interpolation arguments, and cavity-method techniques. In most such examples, the behavior of the optimizer $\hat{\mathbf{x}}\in\mathbb{R}^{n}$ in the limit $n\to\infty$ is characterized by a system of scalar fixed-point equations, derived via mean-field approximations over an interaction matrix having a number of random elements much larger than the dimension $n$ of the optimization variable.
Our current work is motivated by the study of large-$n$ asymptotics for a different type of matrix-valued optimization problem, taking the form
$\min_{X\in\mathbb{R}^{n\times n}}\frac{1}{2}\|XA+BX\|_{F}^{2}+\frac{1}{2}\sum_{i,j=1}^{n}\xi_{i}\theta_{j}x_{ij}^{2}\qquad\text{ subject to }\frac{1}{n}\mathbf{v}^{*}X\mathbf{u}=1.$ (1)
Here, $\boldsymbol{\theta},\boldsymbol{\xi},\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ are deterministic ridge-regularization and linear constraint parameters, and we will study a setting of random inputs given by independent Wigner matrices $A,B\in\mathbb{R}^{n\times n}$. Notably, the optimization variable $X$ has dimension comparable to $A$ and $B$. This problem (1) may be written equivalently in terms of the vectorization $\mathbf{x}=\operatorname{vec}(X)\in\mathbb{R}^{n^{2}}$ and diagonal matrices $\Theta=\operatorname{diag}(\boldsymbol{\theta})$ and $\Xi=\operatorname{diag}(\boldsymbol{\xi})$, as
$\min_{\mathbf{x}\in\mathbb{R}^{n^{2}}}\frac{1}{2}\|(A\otimes I+I\otimes B)\mathbf{x}\|_{2}^{2}+\frac{1}{2}\,\mathbf{x}^{*}(\Theta\otimes\Xi)\mathbf{x}\quad\text{ subject to }\frac{1}{n}(\mathbf{u}\otimes\mathbf{v})^{*}\mathbf{x}=1.$
Such a problem is paradigmatic of a broader class of nonlinear problems/models having a Kronecker-product structure: For motivation, let us mention the matrix spin glass model [^1] We would like to thank Justin Ko for bringing a model similar to (2) to our attention.
$p(X)$ $=\frac{1}{\mathcal{Z}}\exp(\operatorname{Tr}AX^{*}X+\operatorname{Tr}BXX^{*})$ (2)
$=\frac{1}{\mathcal{Z}}\exp(\operatorname{vec}(X)^{*}(A\otimes I+I\otimes B)\operatorname{vec}(X)),\qquad X\in\{\pm 1\}^{n\times n}$
defined by independent GOE coupling matrices $A,B\in\mathbb{R}^{n\times n}$, and the optimization problem
$\min_{X\in\mathbb{R}^{n\times n}}\|XA+BX\|\quad\text{ subject to }\sum_{i=1}^{n}x_{ij}=\sum_{j=1}^{n}x_{ij}=1,\;x_{ij}\geq 0\text{ for all }i,j=1,\ldots,n$ (3)
defined by (possibly entrywise correlated) Wigner matrices $A,B\in\mathbb{R}^{n\times n}$. The former model (2) describes a disordered spin system on the lattice, with $A,B$ representing couplings for the row and column inner-products of $X$. The latter problem (3) for various choices of matrix norm $\|XA+BX\|$ corresponds to popular convex relaxations of combinatorial optimization problems over permutation matrices $X$ that arise in random graph matching [3, 66, 1].
As a step towards developing techniques and insight for asymptotic analyses of these types of Kronecker-structured models, in this work we carry out an analysis of the simpler linear problem (1) using random matrix theory methods. We will establish a deterministic approximation with $O_{\prec}(n^{-1/2})$ error for the value of the objective (1) at its minimizer $\widehat{X}$, as well as for the value of $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$ for arbitrary deterministic test vectors $\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$. These results are closely related to a deterministic equivalent approximation for the resolvent of a Kronecker-product random matrix
$Q=A\otimes I_{n\times n}+I_{n\times n}\otimes B+\Theta\otimes\Xi\in\mathbb{C}^{n^{2}\times n^{2}},$ (4)
which we will refer to as the “Kronecker deformed Wigner model” (in analogy with the deformed Wigner model $A+\Theta$ studied classically in random matrix theory [55, 10, 41, 42, 43, 40]). A second main focus of our work is to establish sharp quantitative estimates for the resolvent $G=(Q-zI)^{-1}$ of this model at global spectral scales, i.e. for fixed spectral parameters $z\in\mathbb{C}^{+}$. Writing $G_{ij}=(\mathbf{e}_{i}\otimes I)^{*}G(\mathbf{e}_{j}\otimes I)\in\mathbb{C}^{n\times n}$ and $G_{ij,\alpha\beta}=(\mathbf{e}_{i}\otimes\mathbf{e}_{\alpha})^{*}G(\mathbf{e}_{j}\otimes\mathbf{e}_{\beta})\in\mathbb{C}$ for the blocks and entries of this resolvent, we will show the operator-norm estimates
$\|G_{ii}^{-1}-G_{jj}^{-1}-(\theta_{i}-\theta_{j})\Xi\|_{\mathrm{op}}\prec n^{-1/2},\qquad\|G_{ij}\|_{\mathrm{op}}\prec n^{-1/2}\text{ for }i\neq j,$ (5)
the entrywise estimates
$G_{ii,\alpha\alpha}-[G_{0}]_{ii,\alpha\alpha}\prec n^{-1/2},\qquad G_{ii,\alpha\beta}\prec n^{-1/2}\text{ for }\alpha\neq\beta,\qquad G_{ij,\alpha\beta}\prec n^{-1}\text{ for }\alpha\neq\beta,i\neq j,$ (6)
and the bilinear-form estimates
$(\mathbf{u}\otimes\mathbf{v})^{*}[G-G_{0}](\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})\prec n^{-1/2}\text{ for deterministic unit vectors }\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{C}^{n},$ (7)
where $G_{0}\in\mathbb{C}^{n^{2}\times n^{2}}$ is a diagonal deterministic-equivalent matrix. Some high-level aspects of this model that may be suggestive of properties to be expected also in nonlinear models such as (2) and (3) include a characterization of the large-$n$ limit by a pair of fixed-point equations in an operator algebra rather than over the scalar field, and a “two-tiered” mean-field structure as reflected in (6), which arises from separate mean-field approximations over $A$ and $B$.
A special case of the optimization problem (1) with $\Theta=\Xi=\eta I_{n\times n}$ (and correlated Wigner inputs $A$ and $-B$) was previously analyzed in [30], in the context of the graph matching application. We review a central idea of this previous analysis in Appendix A, which, however, is special to a commutative setting and does not extend more generally. Here, we instead follow an alternative approach of a two-stage Schur-complement analysis of the resolvent, in each stage applying, in an operator-algebra setting, ideas around Dyson fixed-point equations and fluctuation averaging techniques developed for Wigner-type models in [26, 27, 23, 41] and for $\mathbb{C}^{k\times k}\otimes\mathbb{C}^{n\times n}$-valued Kronecker matrix models with fixed dimension $k$ in [4, 25]. The application of these methods in our setting of $k(n)=n$ seems to require new ideas, even to obtain optimal quantitative estimates for fixed spectral parameters on the global scale, and we discuss this further in Section 1.1 below.
Recently, a breakthrough line of work has obtained sharp operator norm estimates for polynomial matrix models in Kronecker-product spaces with growing first dimension $k\equiv k(n)$ [12, 52, 54, 5, 7, 53, 8]. The authors of [12] developed a new approach for estimating expected traces of smooth functions of polynomials in
$(A_{1}\otimes I_{n\times n},\ldots,A_{r}\otimes I_{n\times n},I_{k\times k}\otimes B_{1},\ldots,I_{k\times k}\otimes B_{s})$ (8)
for deterministic matrices $A_{1},\ldots,A_{r}\in\mathbb{C}^{k\times k}$ and independent GUE matrices $B_{1},\ldots,B_{s}\in\mathbb{C}^{n\times n}$, via an interpolation between the GUE and free semicircular variables and a differential calculus using Gaussian integration-by-parts and the semicircular Schwinger-Dyson equation. These results implied an estimate with $O(\frac{k^{2}}{n^{2}\operatorname{poly}(\Im z)})$-error for the expectation of the Stieltjes transform, and a strong convergence result (i.e. convergence in operator norm) when $k\ll n^{1/3}$. This was extended into a full asymptotic expansion in [54] and from GUE to Haar unitary matrices in [52, 53], with [53] deducing strong convergence for Kronecker-product polynomials in Haar unitary and deterministic matrices having first dimension $k\leq n/\operatorname{poly}(\log n)$. An analogous strong convergence statement in the Gaussian setting was shown as part of the general results of [5], using a different interpolation idea. The authors of [7] showed strong convergence for (8) when $k=n$ and $A_{1},\ldots,A_{r},B_{1},\ldots,B_{s}$ are all independent GUE matrices, via a mapping to unitary matrices, an extension of the asymptotic expansion in [54], and a precise analysis of the large-$|z|$ expansion of the expected Stieltjes transform. In [8], strong convergence for Kronecker-product polynomials of Haar-unitary matrices up to dimensions $k\leq\exp(n^{\alpha})$ was shown using a different non-backtracking high trace method.
The focus of our work is a bit different from the above, as we will not address this question of strong convergence for our model, but instead study the detailed structure of its resolvent in addition to its spectral measure. We carry out a more classical analysis using the resolvent calculus, and for our current purposes, we will also not separate the analysis of the expectation of the resolvent from its fluctuations. (In particular, we will not establish an estimate for the expected Stieltjes transform that is more precise than the $O_{\prec}(n^{-1})$ scale of its fluctuations, as would be needed to show strong convergence.) However, we highlight that our analyses apply directly to non-Gaussian Wigner matrices, and also yield sharp operator-norm estimates of the form (5) for the $n\times n$ resolvent blocks, which may be more difficult to obtain via arguments based solely on large-$|z|$ expansions and analytic continuation ideas.
1.1. Proof ideas
We present here the high-level ideas of the arguments, deferring to Sections 2 and 3 a more detailed description of the model and notations. Consider first the model
$Q=A\otimes I+I\otimes B+\Theta\otimes\Xi\in\mathbb{C}^{n^{2}\times n^{2}}$
and denote its resolvent $G=G(z)=(Q-zI)^{-1}$. The strategy will be to perform a two-stage Schur-complement analysis of this resolvent, first over the randomness of $A$, and then over $B$.
In the first stage, conditioning on $B$, the independence structure of $A\otimes I$ suggests a Schur-complement analysis at the level of its $n\times n$ blocks: Denoting $G_{ij}=(\mathbf{e}_{i}\otimes I)^{*}G(\mathbf{e}_{j}\otimes I)\in\mathbb{C}^{n\times n}$ and applying standard resolvent identities, we have
$G_{ii}=\Big{(}a_{ii}I+B+\theta_{i}\Xi-zI-\sum_{r,s}^{(i)}a_{ir}G_{rs}^{(i)}a_{si}\Big{)}^{-1},\qquad G_{ij}={-}G_{ii}\sum_{r}^{(i)}a_{ir}G_{rj}^{(i)}.$ (9)
Applying concentration of the quadratic form in the first expression and averaging over $i=1,\ldots,n$, we will eventually obtain an approximate fixed-point relation for the partial trace
$(n^{-1}\operatorname{Tr}\otimes I)G=\frac{1}{n}\sum_{i=1}^{n}G_{ii}=\sum_{i=1}^{n}\Big{(}B+\theta_{i}\Xi-zI-(n^{-1}\operatorname{Tr}\otimes I)G\Big{)}^{-1}+O_{\prec}(n^{-1}).$
Stability of this equation will imply
$(n^{-1}\operatorname{Tr}\otimes I)G=M_{B}+O_{\prec}(n^{-1}),\quad\text{ for the fixed point }M_{B}=\sum_{i=1}^{n}\Big{(}B+\theta_{i}\Xi-zI-M_{B}\Big{)}^{-1}.$ (10)
In the second stage, we then carry out the analysis over $B$. As the implicit dependence of $M_{B}$ on $B$ in (10) is rather complex and makes a direct Schur-complement analysis difficult, we introduce an explicit operator algebra representation and write
[…]
for a ($n$-dependent) von Neumann algebra $\mathcal{A}$ with trace $\tau$, containing a semicircular variable $\mathsf{a}$ and the subalgebra $\mathbb{C}^{n\times n}$ free of $\mathsf{a}$ (Lemma 4.3). We study the Stieltjes transform $n^{-2}\operatorname{Tr}G\approx n^{-1}\operatorname{Tr}M_{B}$ by first taking $1\otimes n^{-1}\operatorname{Tr}$ inside $\tau\otimes I$ in (11), and applying resolvent identities parallel to (9),
$\tilde{\mathsf{g}}_{\alpha\alpha}=\Big{(}b_{\alpha\alpha}1+\mathsf{a}+\xi_{\alpha}\Theta-z1-\sum_{\gamma,\delta}^{(\alpha)}b_{\alpha\gamma}\tilde{\mathsf{g}}_{\gamma\delta}^{(\alpha)}b_{\delta\alpha}\Big{)}^{-1},\qquad\tilde{\mathsf{g}}_{\alpha\beta}={-}\tilde{\mathsf{g}}_{\alpha\alpha}\sum_{\gamma}^{(\alpha)}b_{\alpha\gamma}\tilde{\mathsf{g}}_{\gamma\beta}^{(\alpha)},$ (12)
to analyze $(1\otimes n^{-1}\operatorname{Tr})\tilde{\mathsf{g}}$. This yields
$(1\otimes n^{-1}\operatorname{Tr})\tilde{\mathsf{g}}=\mathsf{m}_{a}+O_{\prec}(n^{-1}),\quad\text{ for the fixed point }\mathsf{m}_{a}=\sum_{\alpha=1}^{n}\Big{(}\mathsf{a}+\xi_{\alpha}\Theta-z1-\mathsf{m}_{a}\Big{)}^{-1},$ (13)
and applying this back to (11) gives an approximation for $n^{-2}\operatorname{Tr}G$. These arguments also yield, as direct consequences, estimates in operator norm for the resolvent blocks $G_{ii},G_{ij}\in\mathbb{C}^{n\times n}$.
In deducing (10) from (9) and (13) from (12), two difficulties arise that do not occur in usual scalar random matrix models:
  1. (1)
    Applying non-commutative concentration inequalities to analyze (9) and (12), several of the terms controlling the scale of fluctuations cannot be bounded spectrally by the resolvent. For example, a non-commutative Khintchine-type inequality gives for $G_{ij}$ in (9)
    $\mathbb{E}\|G_{ij}\|_{p}^{p}\prec\frac{1}{\sqrt{n}}\max\bigg{\{}\mathbb{E}\bigg{\|}\bigg{(}\sum_{r}G_{rj}^{(i)}G_{rj}^{(i)*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\,\mathbb{E}\bigg{\|}\bigg{(}\sum_{r}G_{rj}^{(i)*}G_{rj}^{(i)}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{\}}.$ (14)
    The second term may be directly controlled by $\|G^{*}G\|_{\mathrm{op}}$ and a spectral argument, but the first term is related instead to the spectrum of the partial transpose ${G^{\mathsf{t}}}=\sum_{i,j}E_{ji}\otimes G_{ij}$, for which a naive bound gives $\|G^{\mathsf{t}}\|_{\mathrm{op}}\leq n\|G\|_{\mathrm{op}}$ [64]. Applying this naive bound produces a trivial estimate $G_{ij}\prec 1$. A similar issue arises for the quadratic forms in the expressions of $G_{ii}$ and $\tilde{\mathsf{g}}_{\alpha\alpha}$ in (9) and (12).
  2. (2)
    In the infinite-dimensional context of (12), Khintchine-type inequalities only yield estimates in the $L^{p}$-norms $\|\cdot\|_{p}$ for $p<\infty$, rather than a dimension-free estimate in the operator norm, whereas stability of the fixed-point equation (13) is most readily established under perturbations that are bounded in operator norm.
We address these difficulties by first carrying out the analysis for $|z|$ sufficiently large for which the resolvent admits a convergent series expansion in $z^{-1}$, and furthermore the concentration of errors in (9) and (12) may be established by expanding into elementary tensors and applying scalar concentration term-by-term. Then, using a quantitative version of the maximum modulus principle, we obtain a weak high-probability estimate in the operator norm
$\big{\|}G_{ii}-\big{(}B+\theta_{i}\Xi-zI-(n^{-1}\operatorname{Tr}\otimes I)G\big{)}^{-1}\big{\|}_{\mathrm{op}}<n^{-\alpha}$
for any fixed $z\in\mathbb{C}^{+}$ and a small (sub-optimal) constant $\alpha>0$ depending on $\Im z$ (Lemma 4.9). Such an estimate and the stability of the fixed-point equation (10) is sufficient to deduce
$G_{ii}-\underbrace{(B+\theta_{i}\Xi-zI-M_{B})^{-1}}_{=M_{i}}\prec n^{-\alpha},\qquad G_{ij}\prec n^{-\alpha}.$ (15)
These estimates (15) now enable the application of non-commutative analogues of fluctuation averaging techniques [22], which we use in conjunction with an iterative bootstrapping argument to obtain the optimal estimates for $G_{ii},G_{ij}$ as follows: Writing the first non-spectral term of (14) as
$\frac{1}{n}\sum_{r}G_{rj}^{(i)}G_{rj}^{(i)*}=\frac{1}{n}\sum_{r}\mathbb{E}_{r}[G_{rj}^{(i)}G_{rj}^{(i)*}]+\frac{1}{n}\sum_{r}\mathcal{Q}_{r}[G_{rj}^{(i)}G_{rj}^{(i)*}]$
where $\mathbb{E}_{r}$ is the partial expectation over row/column $r$ of $A$ and $\mathcal{Q}_{r}=1-\mathbb{E}_{r}$, fluctuation averaging with (15) as input shows for the second term $n^{-1}\sum_{r}\mathcal{Q}_{r}[G_{rj}^{(i)}G_{rj}^{(i)*}]\prec n^{-3\alpha}$. Applying a resolvent expansion of the first term $n^{-1}\sum_{r}\mathbb{E}_{r}[G_{rj}^{(i)}G_{rj}^{(i)*}]$ then yields
$\mathcal{L}_{2}\bigg{(}\frac{1}{n}\sum_{r}G_{rj}^{(i)}G_{rj}^{(i)*}\bigg{)}\prec n^{-1}+n^{-3\alpha}$
for the linear operator $\mathcal{L}_{2}(X)=X-\frac{1}{n}\sum_{i}M_{i}XM_{i}^{*}$, with $M_{i}$ defined in (15). This type of operator associated to Dyson fixed-point equations has been studied previously in [2, 4], and we adapt a Perron-Frobenius argument of [2] to show quantitative invertibility of $\mathcal{L}_{2}$ for any fixed $z\in\mathbb{C}^{+}$. In our context, we require invertibility of $\mathcal{L}_{2}$ in the $L^{p}$-norm for each $p\in[1,\infty)$, and in infinite-dimensional settings where $\mathcal{L}_{2}$ may not have an exact Perron-Frobenius eigenvector — we thus carry out this analysis in the $L^{1}$-$L^{\infty}$ duality rather than the Hilbert-space setting of [2], and appeal to the Riesz-Thorin interpolation to obtain invertibility for all $p$ (Lemma 4.5). This shows $n^{-1}\sum_{r}G_{rj}^{(i)}G_{rj}^{(i)*}\prec n^{-1}+n^{-3\alpha}$, which applied to (14) shows the implication
$G_{ii}-M_{i}\text{ and }G_{ij}\prec n^{-\alpha}\Longrightarrow G_{ii}-M_{i}\text{ and }G_{ij}\prec\max(n^{-1/2},n^{-3\alpha/2}).$
Iterating this bound gives finally the optimal errors $G_{ii}-M_{i}\prec n^{-1/2}$ and $G_{ij}\prec n^{-1/2}$, and an additional fluctuation averaging step shows $n^{-1}\sum_{i}G_{ii}-M_{B}\prec n^{-1}$ for the partial trace, as claimed in (10).
These arguments extend to show quantitative approximations for bilinear forms of the resolvent $(\mathbf{u}\otimes\mathbf{v})^{*}G(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})$, with $O_{\prec}(n^{-1/2})$ error. To show the estimate $G_{ij,\alpha\beta}\prec n^{-1}$ for off-diagonal entries, we apply a similar idea as above, deriving from the resolvent identities and a fluctuation averaging argument, for each fixed $i\neq j$ and $\alpha\neq\beta$,
$\mathcal{L}_{1}\bigg{(}\sum_{k}G_{kj}\mathbf{e}_{\beta}\mathbf{e}_{\alpha}^{*}G_{ik}\bigg{)}\prec n^{-1/2}$
where $\mathcal{L}_{1}(X)=X-\frac{1}{n}\sum_{i=1}^{n}M_{i}XM_{i}$. Quantitative invertibility of $\mathcal{L}_{1}$ follows from differentiation of the fixed-point equation (10), yielding $\sum_{k}G_{kj}\mathbf{e}_{\beta}\mathbf{e}_{\alpha}^{*}G_{ik}\prec n^{-1/2}$, and we will deduce from this the estimate $G_{ij,\alpha\beta}\prec n^{-1}$ (Section 5). Finally, results for the least-squares problem (1) are obtained via a similar analysis of a linearized model (Section 6)
$\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\otimes(A\otimes I+I\otimes B)-i\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes\Theta\otimes\Xi-i\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\otimes I\otimes I\in\mathbb{C}^{2\times 2}\otimes\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}.$
1.2. Notation and conventions
$(\mathbf{e}_{1},\ldots,\mathbf{e}_{n})\in\mathbb{C}^{n}$ are the standard basis vectors, $E_{ij}=\mathbf{e}_{i}\mathbf{e}_{j}^{*}$ are the coordinate basis elements of $\mathbb{C}^{n\times n}$, and $I_{n\times n}$ is the identity matrix (omitting the subscript when the meaning is clear). For $M\in\mathbb{C}^{n\times n}$, $\|M\|_{F}=(\sum_{i,j}|m_{ij}|^{2})^{1/2}$ and $\operatorname{Tr}M=\sum_{i}m_{ii}$ are the Frobenius norm and (unnormalized) trace. We write $U=\operatorname{diag}(\mathbf{u})$ for the diagonal matrix with $\mathbf{u}\in\mathbb{C}^{n}$ on its diagonal, and $\mathbf{u}=\operatorname{diag}(U)$ for the vector $(U_{11},\ldots,U_{nn})$ on the diagonal of $U\in\mathbb{C}^{n\times n}$.
For a von Neumann algebra $\mathcal{X}$, we denote its unit as $1_{\mathcal{X}}$ (omitting the subscript when the meaning is clear). Scalars $z\in\mathbb{C}$ are identified implicitly as elements of $\mathcal{X}$ via $z\mapsto z1_{\mathcal{X}}$. We write $\mathsf{x}\geq 0$ if $\mathsf{x}\in\mathcal{X}$ is self-adjoint and has nonnegative spectrum. For $\mathsf{x},\mathsf{y}\in\mathcal{X}$ self-adjoint, $\mathsf{x}\geq\mathsf{y}$ means $\mathsf{x}-\mathsf{y}\geq 0$. We write
$\Re\mathsf{x}=\tfrac{1}{2}(\mathsf{x}+\mathsf{x}^{*}),\qquad\Im\mathsf{x}=\tfrac{1}{2i}(\mathsf{x}-\mathsf{x}^{*}),\qquad|\mathsf{x}|=(\mathsf{x}^{*}\mathsf{x})^{1/2}$
for the operator real and imaginary parts and absolute value. We denote
$\mathcal{X}^{+}=\{\mathsf{x}\in\mathcal{X}:\Im\mathsf{x}\geq\epsilon\text{ forsome }\epsilon>0\}$
as the elements with strictly positive imaginary part (not to be confused with the positive cone $\{\mathsf{x}\in\mathcal{X}:\mathsf{x}\geq 0\}$). $\|\mathsf{x}\|_{\mathrm{op}}$ is the operator norm, and $\|\mathsf{x}\|_{p}=\phi(|\mathsf{x}|^{p})^{1/p}$ is the $L^{p}$-norm for a given trace $\phi$.
2. Model and main results
2.1. Kronecker deformed Wigner model
Assumption 2.1.
  1. (a)
    $A,B\in\mathbb{C}^{n\times n}$ are independent Wigner matrices (defined on an underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$) satisfying $A=A^{*}$ and $B=B^{*}$, and having independent entries $(a_{ij},b_{ij})_{i\leq j}$ such that, for all $1\leq i<j\leq n$,
    $\mathbb{E}a_{ii}=\mathbb{E}b_{ii}=\mathbb{E}a_{ij}=\mathbb{E}b_{ij}=0,\qquad\mathbb{E}|a_{ij}|^{2}=\mathbb{E}|b_{ij}|^{2}=1/n$
    and for all $1\leq i,j\leq n$, each $p\geq 2$ and a constant $C_{p}>0$,
    $\mathbb{E}|a_{ii}|^{p},\mathbb{E}|a_{ij}|^{p},\mathbb{E}|b_{ii}|^{p},\mathbb{E}|b_{ij}|^{p}\leq C_{p}n^{-p/2}.$ (16)
    (This includes the special case where $A,B$ are real and symmetric.)
  2. (b)
    $\Theta=\operatorname{diag}(\theta_{1},\ldots,\theta_{n})\in\mathbb{R}^{n\times n}$ and $\Xi=\operatorname{diag}(\xi_{1},\ldots,\xi_{n})\in\mathbb{R}^{n\times n}$ are deterministic diagonal matrices satisfying $\|\Theta\|_{\mathrm{op}},\|\Xi\|_{\mathrm{op}}\leq\upsilon$ for a constant $\upsilon>0$.
We study the Kronecker deformed Wigner model
$Q=A\otimes I_{n\times n}+I_{n\times n}\otimes B+\Theta\otimes\Xi\in\mathbb{C}^{n^{2}\times n^{2}}.$ (17)
For spectral arguments $z\in\mathbb{C}^{+}$, define the resolvent and Stieltjes transform of $Q$ by
$G(z)$ $=(Q-z\,I_{n\times n}\otimes I_{n\times n})^{-1},$
$m(z)$ $=n^{-2}\operatorname{Tr}G(z)=(n^{-1}\operatorname{Tr}\,\otimes\,n^{-1}\operatorname{Tr})\Big{[}(Q-z\,I_{n\times n}\otimes I_{n\times n})^{-1}\Big{]}.$
We will use Roman indices for the first tensor factor, Greek indices for the second, and write the blocks and entries of $G(z)$ as
$G_{ij}=(\mathbf{e}_{i}\otimes I)^{*}G(\mathbf{e}_{j}\otimes I)\in\mathbb{C}^{n\times n},\qquad G_{\alpha\beta}=(I\otimes\mathbf{e}_{\alpha})^{*}G(I\otimes\mathbf{e}_{\beta})\in\mathbb{C}^{n\times n},$
$G_{ij,\alpha\beta}=(\mathbf{e}_{i}\otimes\mathbf{e}_{\alpha})^{*}G(\mathbf{e}_{j}\otimes\mathbf{e}_{\beta})\in\mathbb{C}.$
Our main result will establish deterministic equivalent approximations for $G(z)$ and $m(z)$. These may be defined through a free probability construction, in the following setting:
Assumption 2.2.
  1. (a)
    $\mathcal{A}$ is a von Neumann algebra with unit $1_{\mathcal{A}}$, operator norm $\|\cdot\|_{\mathrm{op}}$, and faithful, normal, tracial state $\tau:\mathcal{A}\to\mathbb{C}$.
  2. (b)
    $\mathcal{A}$ contains as a von Neumann subalgebra
    $\mathcal{M}\equiv\mathbb{C}^{n\times n}$
    where $I_{n\times n}\in\mathcal{M}\subset\mathcal{A}$ coincides with $1_{\mathcal{A}}$, and $\tau|_{\mathcal{M}}$ and $\|\cdot\|_{\mathrm{op}}|_{\mathcal{M}}$ restrict to the normalized matrix trace $\frac{1}{n}\operatorname{Tr}$ and matrix operator norm on $\mathcal{M}$.
  3. (c)
[…]
Let $\mathcal{D}\subset\mathcal{M}$ be the subalgebra of diagonal matrices, generated by the diagonal basis elements $\{E_{ii}\}_{i=1}^{n}$, and denote by $\tau^{\mathcal{D}}:\mathcal{A}\to\mathcal{D}$ the diagonal projection
$\tau^{\mathcal{D}}(\mathsf{x})=\sum_{i=1}^{n}\frac{\tau(\mathsf{x}E_{ii})}{\tau(E_{ii})}E_{ii}=\sum_{i=1}^{n}n\,\tau(\mathsf{x}E_{ii})E_{ii}.$ (18)
(This is the unique $\tau$-preserving conditional expectation onto $\mathcal{D}$, in the sense of Lemma D.2.) We implicitly identify $\mathcal{A}$ with its representation on an underlying Hilbert space $\mathcal{H}$, and denote by $\mathcal{A}\otimes\mathcal{A}$ the von Neumann tensor product acting on $\mathcal{H}\otimes\mathcal{H}$. We denote by $\tau\otimes\tau:\mathcal{A}\otimes\mathcal{A}\to\mathbb{C}$, $\tau\otimes 1:\mathcal{A}\otimes\mathcal{A}\to\mathcal{A}$, and $1\otimes\tau:\mathcal{A}\otimes\mathcal{A}\to\mathcal{A}$ the unique bounded linear maps satisfying
$(\tau\otimes\tau)(\mathsf{x}\otimes\mathsf{y})=\tau(\mathsf{x})\tau(\mathsf{y}),\qquad(\tau\otimes 1)(\mathsf{x}\otimes\mathsf{y})=\tau(\mathsf{x})\mathsf{y},\qquad(1\otimes\tau)(\mathsf{x}\otimes\mathsf{y})=\tau(\mathsf{y})\mathsf{x},$
and similarly for $\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}}$, $\tau^{\mathcal{D}}\otimes 1$, and $1\otimes\tau^{\mathcal{D}}$.
Identifying $\Theta,\Xi\in\mathcal{M}$ as elements of $\mathcal{A}$ (which are free of $\mathsf{a},\mathsf{b}$ by assumption), in the limit $n\to\infty$, an approximation of our matrix of interest $Q$ in the tracial sense is given by
$\mathsf{q}=\mathsf{a}\otimes 1+1\otimes\mathsf{b}+\Theta\otimes\Xi\in\mathcal{A}\otimes\mathcal{A}.$
For spectral arguments $z\in\mathbb{C}^{+}$, we define its (deterministic, $n$-dependent) resolvent and Stieltjes transform by
[…]
$m_{0}(z)$ $=\tau\otimes\tau[\mathsf{g}(z)]\in\mathbb{C}^{+}.$
The deterministic equivalent matrix $G_{0}(z)$ for $G(z)$ is then given by
$G_{0}(z)=(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[\mathsf{g}(z)]\in\mathcal{D}\otimes\mathcal{D}.$
We remark that $G_{0}(z)$ is a deterministic diagonal matrix in $\mathbb{C}^{n^{2}\times n^{2}}$. (We refer to [32, Theorem 4.4] for a previous example of this type of construction, in a different model.)
In contrast to more classical random matrix models, the above Stieltjes transform $m_{0}(z)$ in general does not seem to admit a simple characterization in terms of scalar-valued fixed-point equations. However, it may be characterized via a pair of fixed-point equations in the operator algebra $\mathcal{A}$, and $G_{0}(z)$ may be constructed from this characterization in the following way.
Proposition 2.3.
In the setting of Assumption 2.2, set $\mathcal{A}^{+}=\{\mathsf{x}\in\mathcal{A}:\Im\mathsf{x}\geq\epsilon\text{ forsome }\epsilon>0\}$. For any $z\in\mathbb{C}^{+}$, there exist unique elements $\mathsf{m}_{a}(z),\mathsf{m}_{b}(z)\in\mathcal{A}^{+}$ satisfying the fixed-point equations
$\mathsf{m}_{a}(z)$ $=\frac{1}{n}\sum_{\alpha=1}^{n}(\mathsf{a}+\xi_{\alpha}\Theta-z-\mathsf{m}_{a}(z))^{-1},$ (19)
$\mathsf{m}_{b}(z)$ $=\frac{1}{n}\sum_{i=1}^{n}(\mathsf{b}+\theta_{i}\Xi-z-\mathsf{m}_{b}(z))^{-1}.$ (20)
We have
$m_{0}(z)$ $=\tau[\mathsf{m}_{a}(z)]=\tau[\mathsf{m}_{b}(z)],$
$G_{0}(z)$ $=\sum_{\alpha=1}^{n}\tau^{\mathcal{D}}\big{[}(\mathsf{a}+\xi_{\alpha}\Theta-\mathsf{z}-\mathsf{m}_{a}(z))^{-1}\big{]}\otimes E_{\alpha\alpha}=\sum_{i=1}^{n}E_{ii}\otimes\tau^{\mathcal{D}}\big{[}(\mathsf{b}+\theta_{i}\Xi-z-\mathsf{m}_{b}(z))^{-1}\big{]}.$
Our main results for this model are the following estimates for its Stieltjes transform $m(z)$ and resolvent $G(z)$. Here, the stochastic domination notation $\xi\prec\zeta$ (c.f. [22]) means $\mathbb{P}[|\xi|>n^{\epsilon}\zeta]\leq n^{-D}$ for any fixed $\epsilon,D>0$ and all $n\geq n_{0}(\epsilon,D)$; we review this definition in Section 3.1.
Theorem 2.4.
Fix any $\upsilon,\delta>0$. Under Assumptions 2.1 and 2.2, uniformly over $z\in\mathbb{C}^{+}$ with $|z|\leq\upsilon$ and $\Im z\geq\delta$, and over $i,j,\alpha,\beta\in\{1,\ldots,n\}$ with $i\neq j$ and $\alpha\neq\beta$, we have:
  1. (a)
    (Stieltjes transform)
    $|m(z)-m_{0}(z)|\prec n^{-1}.$ (21)
  2. (b)
    (Resolvent blocks)
    $\Big{\|}G_{ii}^{-1}-G_{jj}^{-1}-(\theta_{i}-\theta_{j})\Xi\Big{\|}_{\mathrm{op}}$ $\prec n^{-1/2},$ (22)
    $\Big{\|}G_{\alpha\alpha}^{-1}-G_{\beta\beta}^{-1}-(\xi_{\alpha}-\xi_{\beta})\Theta\Big{\|}_{\mathrm{op}}$ $\prec n^{-1/2},$ (23)
    $\|G_{ij}\|_{\mathrm{op}},\|G_{\alpha\beta}\|_{\mathrm{op}}$ $\prec n^{-1/2}.$ (24)
  3. (c)
    (Resolvent entries)
    $G_{ii,\alpha\alpha}-(G_{0})_{ii,\alpha\alpha}$ $\prec n^{-1/2},$ (25)
    $G_{ii,\alpha\beta},G_{ij,\alpha\alpha}$ $\prec n^{-1/2},$ (26)
    $G_{ij,\alpha\beta}$ $\prec n^{-1}.$ (27)
  4. (d)
    (Bilinear forms) Uniformly over vectors $\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{C}^{n}$ with $\|\mathbf{u}\|_{2},\|\mathbf{v}\|_{2},\|\mathbf{u}^{\prime}\|_{2},\|\mathbf{v}^{\prime}\|_{2}\leq\upsilon$,
    $(\mathbf{u}\otimes\mathbf{v})^{*}\big{[}G-G_{0}\big{]}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})\prec n^{-1/2}.$ (28)
The estimates (25), (26), and (27) indicate that the entries of the resolvent fall on three scales of orders 1, $n^{-1/2}$, and $n^{-1}$, as depicted in the numerical simulation of Figure 1.
Figure 1. Entrywise modulus of the resolvent $G(z)=(A\otimes I+I\otimes B+\Theta\otimes\Xi-z)^{-1}$ at $z=i$, where $n=20$, $A,B$ are independent GOE matrices of size $n$, and $\Theta,\Xi$ have independent $\text{Uniform}(-1,1)$ diagonal entries.
2.2. Application to least-squares problem
In the same setting of Assumption 2.1 with real-valued Wigner matrices $A,B\in\mathbb{R}^{n\times n}$ and $\theta_{i},\xi_{j}>0$, consider now the optimization objective
$f(X)=\frac{1}{2}\|XA+BX\|_{F}^{2}+\frac{1}{2}\sum_{i,j=1}^{n}\xi_{i}\theta_{j}x_{ij}^{2}.$ (29)
Given vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$, define its minimizer under a linear constraint
$\widehat{X}=\operatorname*{arg\,min}_{X\in\mathbb{R}^{n\times n}}f(X)\qquad\text{ subject to }\qquad\frac{1}{n}\mathbf{v}^{*}X\mathbf{u}=1.$ (30)
Our main results for this optimization problem (30) are the following asymptotic characterization of the minimum objective value $f(\widehat{X})$ and the values of linear forms $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$ for deterministic test vectors $\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
Theorem 2.5.
Fix any $\upsilon,\delta>0$. Suppose Assumptions 2.1 and 2.2 hold, where $A,B\in\mathbb{R}^{n\times n}$ are real-valued. Associated to $\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$, denote $\mathsf{u}=\operatorname{diag}(\mathbf{u})$, $\mathsf{v}=\operatorname{diag}(\mathbf{v})$, $\mathsf{u}^{\prime}=\operatorname{diag}(\mathbf{u}^{\prime})$, and $\mathsf{v}^{\prime}=\operatorname{diag}(\mathbf{v}^{\prime})$ as diagonal matrices in $\mathcal{D}$. Then, uniformly over $\Theta,\Xi,\mathbf{u},\mathbf{v}$ defining (30) such that $\Theta,\Xi\geq\delta I$ and $\upsilon^{-1}\sqrt{n}\leq\|\mathbf{u}\|_{2},\|\mathbf{v}\|_{2}\leq\upsilon\sqrt{n}$:
  1. (a)
    (Objective value)
    $f(\widehat{X})=\frac{1}{2}\cdot\frac{1}{(\tau\otimes\tau)[(\mathsf{u}\otimes\mathsf{v})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathsf{u}\otimes\mathsf{v})]}+O_{\prec}(n^{-1/2})$
  2. (b)
    (Linear projection) Uniformly over $\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$ with $\|\mathbf{u}^{\prime}\|_{2},\|\mathbf{v}^{\prime}\|_{2}\leq\upsilon\sqrt{n}$,
    $\frac{1}{n}\,{\mathbf{u}^{\prime}}^{*}\widehat{X}{\mathbf{v}^{\prime}}=\frac{(\tau\otimes\tau)[(\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathsf{u}\otimes\mathsf{v})]}{(\tau\otimes\tau)[(\mathsf{u}\otimes\mathsf{v})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathsf{u}\otimes\mathsf{v})]}+O_{\prec}(n^{-1/2}).$
We remark that the above value
$(\tau\otimes\tau)[(\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathsf{u}\otimes\mathsf{v})]$ (31)
may be understood as
$n^{-2}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathbf{u}\otimes\mathbf{v}),$ (32)
where $(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}\in\mathbb{R}^{n^{2}\times n^{2}}$ is a deterministic-equivalent approximation for the matrix $[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}$ arising in the vectorization of the objective (29), and these expressions (31) and (32) coincide because this deterministic-equivalent matrix is diagonal.
In an asymptotic setting, if the empirical distributions of coordinates of $\Theta,\Xi,\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}$ converge to deterministic limits, then a qualitative implication of Theorem 2.5 is a characterization of the almost-sure limit values of $f(\widehat{X})$ and $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$, as summarized in the following corollary.
Corollary 2.6.
Asymptotically as $n\to\infty$, if the empirical distributions of coordinates of $\boldsymbol{\theta}=\operatorname{diag}(\Theta)$, $\boldsymbol{\xi}=\operatorname{diag}(\Xi)$, and $\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}$ satisfy
$\frac{1}{n}\sum_{i=1}^{n}\delta_{\theta_{i},u_{i},u_{i}^{\prime}}\Rightarrow\mathcal{P},\qquad\frac{1}{n}\sum_{\alpha=1}^{n}\delta_{\xi_{\alpha},v_{\alpha},v_{\alpha}^{\prime}}\Rightarrow\mathcal{Q}$
weakly for two joint laws $\mathcal{P},\mathcal{Q}$ on $\mathbb{R}^{3}$, then there exist almost-sure limit values
$T(\mathcal{P},\mathcal{Q})$ $=\lim_{n\to\infty}(\tau\otimes\tau)[(\mathsf{u}\otimes\mathsf{v})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathsf{u}\otimes\mathsf{v})],$
$T^{\prime}(\mathcal{P},\mathcal{Q})$ $=\lim_{n\to\infty}(\tau\otimes\tau)[(\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathsf{u}\otimes\mathsf{v})]$
depending only on $\mathcal{P},\mathcal{Q}$, and almost surely as $n\to\infty$,
$f(\widehat{X})\to\frac{1}{2T(\mathcal{P},\mathcal{Q})},\qquad\frac{1}{n}{\mathbf{u}^{\prime}}^{*}\widehat{X}\mathbf{v}^{\prime}\to\frac{T^{\prime}(\mathcal{P},\mathcal{Q})}{T(\mathcal{P},\mathcal{Q})}.$
2.2.1. Numerical computation
Since operator fixed-point equations of the form in Proposition 2.3 are not directly amenable to numerical computation, we provide in this section a procedure for numerically approximating the values of $T(\mathcal{P},\mathcal{Q})$ and $T^{\prime}(\mathcal{P},\mathcal{Q})$ in Corollary 2.6 using a moment expansion. (This procedure applies equally to approximate the finite-$n$ value of (31) rather than its limit $T^{\prime}(\mathcal{P},\mathcal{Q})$, upon replacing the expectations in (3334) below by averages under the empirical measures of coordinates $\frac{1}{n}\sum_{i=1}^{n}\delta_{\theta_{i},u_{i},u_{i}^{\prime}}$ and $\frac{1}{n}\sum_{\alpha=1}^{n}\delta_{\xi_{\alpha},v_{\alpha},v_{\alpha}^{\prime}}$.)
For each even integer $m\geq 2$, let $\operatorname{NC}_{2,2}(m)$ denote the set of ordered pairs $(\rho_{a},\rho_{b})$ where $\rho_{a},\rho_{b}$ are non-crossing pairings of two disjoint even-cardinality subsets of $\{1,2,\ldots,m\}$ whose union is all of $\{1,2,\ldots,m\}$. (The pairings $\rho_{a}$ and $\rho_{b}$ may cross each other, and either pairing may be empty.) We associate to each $(\rho_{a},\rho_{b})\in\operatorname{NC}_{2,2}(m)$ a value $\mathrm{val}(\rho_{a},\rho_{b})$ in the following way:
  1. (1)
    Let $w_{a}$ be the word in the letters $\{\mathsf{a},\mathsf{d}\}$ that is obtained by traversing the elements of $\{1,\ldots,m\}$ in sequential order, and writing $\mathsf{d}\mathsf{a}$ for odd elements in $\rho_{a}$, $\mathsf{a}\mathsf{d}$ for even elements in $\rho_{a}$, and $\mathsf{d}$ for all elements in $\rho_{b}$.
    Similarly, let $w_{b}$ be the word obtained by writing $\mathsf{d}\mathsf{b}$ for odd elements in $\rho_{b}$, $\mathsf{b}\mathsf{d}$ for even elements in $\rho_{b}$, and $\mathsf{d}$ for all elements in $\rho_{a}$.
  2. (2)
    Define the complement $K(\rho_{a})$ of $\rho_{a}$ in $w_{a}$ as the coarsest non-crossing partition of its letters $\mathsf{d}$, for which $\rho_{a}\cup K(\rho_{a})$ forms a non-crossing partition of all letters of $w_{a}$. Define similarly the complement $K(\rho_{b})$ of $\rho_{b}$ in $w_{b}$.
  3. (3)
    Finally, let $S_{1}$ be the block of $K(\rho_{a})$ that contains the first letter of $w_{a}$, let $T_{1}$ be the block of $K(\rho_{b})$ that contains the first letter of $w_{b}$, and set
    $\mathrm{val}(\rho_{a},\rho_{b})$ $=\mathbb{E}\Big{[}\mathsf{U}\mathsf{U}^{\prime}\theta^{-(|S_{1}|+2)/2}\Big{]}\prod_{S\in K(\rho_{a})\setminus S_{1}}\mathbb{E}\Big{[}\theta^{-|S|/2}\Big{]}\times$
    $\hskip 72.26999pt\mathbb{E}\Big{[}\mathsf{V}\mathsf{V}^{\prime}\xi^{-(|T_{1}|+2)/2}\Big{]}\prod_{T\in K(\rho_{b})\setminus T_{1}}\mathbb{E}\Big{[}\xi^{-|T|/2}\Big{]}$ (33)
    where the expectations are over $(\theta,\mathsf{U},\mathsf{U}^{\prime})\sim\mathcal{P}$ and $(\xi,\mathsf{V},\mathsf{V}^{\prime})\sim\mathcal{Q}$.
For $m=0$, we take $\operatorname{NC}_{2,2}(0)$ to consist of the single pair $(\rho_{a},\rho_{b})$ with $\rho_{a}=\rho_{b}=\emptyset$ both as the empty pairing, and set
$\mathrm{val}(\emptyset,\emptyset)=\mathbb{E}[\mathsf{U}\mathsf{U}^{\prime}\theta^{-1}]\mathbb{E}[\mathsf{V}\mathsf{V}^{\prime}\xi^{-1}].$ (34)
To illustrate this in an example, suppose $m=6$, $\rho_{a}=\{(1,3)\}$, and $\rho_{b}=\{(2,6),(4,5)\}$. Then
$w_{a}$ $=\underbrace{\mathsf{d}\mathsf{a}}_{1}\underbrace{\mathsf{d}}_{2}\underbrace{\mathsf{d}\mathsf{a}}_{3}\underbrace{\mathsf{d}}_{4}\underbrace{\mathsf{d}}_{5}\underbrace{\mathsf{d}}_{6}=\mathsf{d}\mathsf{a}\mathsf{d}\mathsf{d}\mathsf{a}\mathsf{d}\mathsf{d}\mathsf{d}$
$w_{b}$ $=\underbrace{\mathsf{d}}_{1}\underbrace{\mathsf{b}\mathsf{d}}_{2}\underbrace{\mathsf{d}}_{3}\underbrace{\mathsf{b}\mathsf{d}}_{4}\underbrace{\mathsf{d}\mathsf{b}}_{5}\underbrace{\mathsf{b}\mathsf{d}}_{6}=\mathsf{d}\mathsf{b}\mathsf{d}\mathsf{d}\mathsf{b}\mathsf{d}\mathsf{d}\mathsf{b}\mathsf{b}\mathsf{d}$
Re-numbering the letters of $w_{a}$ as $\{1,\ldots,8\}$, $\rho_{a}$ now corresponds to the pairing $\{(2,5)\}$ of the letters $\mathsf{a}$, and $K_{a}$ is the partition $\{(1,6,7,8),(3,4)\}$ of the letters $\mathsf{d}$. Similarly, re-numbering the letters of $w_{b}$ as $\{1,\ldots,10\}$, $\rho_{b}$ is now the pairing $\{(2,9),(5,8)\}$ of the letters $\mathsf{b}$, and $K_{b}$ is the partition $\{(1,10),(3,4),(6,7)\}$ of the letters $\mathsf{d}$. Then $K_{a}$ has blocks of sizes $4,2$, $K_{b}$ has blocks of sizes $2,2,2$, so
$\mathrm{val}(\rho_{a},\rho_{b})=\mathbb{E}[\mathsf{U}\mathsf{U}^{\prime}\theta^{-3}]\mathbb{E}[\theta^{-1}]\mathbb{E}[\mathsf{V}\mathsf{V}^{\prime}\xi^{-2}]\mathbb{E}[\xi^{-1}]\mathbb{E}[\xi^{-1}].$
The values $T(\mathcal{P},\mathcal{Q})$ and $T^{\prime}(\mathcal{P},\mathcal{Q})$ then admit the following series approximations.
Proposition 2.7.
Let $(\theta,\mathsf{U},\mathsf{U}^{\prime})\sim\mathcal{P}$ and $(\xi,\mathsf{V},\mathsf{V}^{\prime})\sim\mathcal{Q}$ where $\mathcal{P},\mathcal{Q}$ are the joint laws on $\mathbb{R}^{3}$ of Corollary 2.6. Denote $\|\mathsf{U}\|_{\infty}=\max\{|x|:x\in\mathrm{supp}(\mathsf{U})\}$ where $\mathrm{supp}(\mathsf{U})$ is the support of $\mathsf{U}$. Set $\eta=\min\{\sqrt{x_{a}x_{b}}:x_{a}\in\mathrm{supp}(\theta),\,x_{b}\in\mathrm{supp}(\xi)\}$. Then for any $M\geq 1$ and $z>1$,
$T^{\prime}(\mathcal{P},\mathcal{Q})$ $=\mathop{\sum_{m=0}^{M-1}}_{m\text{ is even}}\left(\frac{(-1)^{m/2}}{z(z-1)^{m}}\sum_{k=m}^{M-1}\binom{k}{m}\left(\frac{z-1}{z}\right)^{k}\right)\sum_{(\rho_{a},\rho_{b})\in\operatorname{NC}_{2,2}(m)}\mathrm{val}(\rho_{a},\rho_{b})+r_{M}$
where
$|r_{M}|\leq\|\mathsf{U}\|_{\infty}\|\mathsf{V}\|_{\infty}\|\mathsf{U}^{\prime}\|_{\infty}\|\mathsf{V}^{\prime}\|_{\infty}\cdot\eta^{-2}\left(\frac{\sqrt{(z-1)^{2}+16\eta^{-2}}}{z}\right)^{M}.$ (35)
The same result holds for $T(\mathcal{P},\mathcal{Q})$ upon replacing $\mathsf{U}^{\prime},\mathsf{V}^{\prime}$ in (33), (34), and (35) by $\mathsf{U},\mathsf{V}$.
In particular, choosing $z=1+16\eta^{-2}$, the remainder is bounded as
$|r_{M}|\leq\|\mathsf{U}\|_{\infty}\|\mathsf{V}\|_{\infty}\|\mathsf{U}^{\prime}\|_{\infty}\|\mathsf{V}^{\prime}\|_{\infty}\cdot\eta^{-2}\Big{(}\frac{z-1}{z}\Big{)}^{M/2},$
so the series in $m$ converges geometrically (with faster convergence for larger values of the regularization $\eta$). Then $T(\mathcal{P},\mathcal{Q})$ and $T^{\prime}(\mathcal{P},\mathcal{Q})$ may be approximated to high numerical accuracy by computing $\sum_{(\rho_{a},\rho_{b})\in\operatorname{NC}_{2,2}(m)}\mathrm{val}(\rho_{a},\rho_{b})$ for a small number of terms $m=0,2,\ldots,M$. A numerical illustration is provided in Figure 2.
Figure 2. Values of $1/(2f(\widehat{X}))$ obtained from solving (30) across 10 independent realizations (solid dots, with vertical lines indicating 1 standard deviation) versus the theoretical prediction $T(\mathcal{P},\mathcal{Q})$ computed from Proposition 2.7 with $M=12$ (dashed lines). Here $A,B$ are GOE matrices of size $n=1000$, we take $\theta_{i}^{-1},\xi_{\alpha}^{-1},u_{i}^{2},v_{\alpha}^{2}\sim\text{Uniform}(0.05,0.5)/k$ with $k\in\{1,1.3,3\}$, and the horizontal axis indicates the correlation between coordinate pairs $(\theta_{i}^{-1},u_{i}^{2})$ and between $(\xi_{\alpha}^{-1},v_{\alpha}^{2})$.
3. Preliminaries
In the remainder of the paper, we prove the preceding results. We summarize in this section the technical tools needed for our analyses. Section 4 contains the core of our main argument for analyzing a deformed Wigner-type matrix model in an operator-algebra setting. Section 5 completes the analyses for the Kronecker deformed Wigner model (17), and Section 6 completes the analyses for the least-squares problem (30).
Throughout this section and Section 4, $\mathcal{X}$ is a von Neumann algebra acting on a (complex) Hilbert space $\mathcal{H}$, having unit $1_{\mathcal{X}}$, operator norm $\|\cdot\|_{\mathrm{op}}$, and a faithful, normal, tracial state $\phi:\mathcal{X}\to\mathbb{C}$. We recall that this means $\phi$ is a linear functional satisfying
$|\phi(\mathsf{x})|\leq\|\mathsf{x}\|_{\mathrm{op}},\qquad\phi(1_{\mathcal{X}})=1,\qquad\phi(\mathsf{x}\mathsf{y})=\phi(\mathsf{y}\mathsf{x})\text{ for all }\mathsf{x},\mathsf{y}\in\mathcal{X},$
$\phi(\mathsf{x})\geq 0\text{ for all }\mathsf{x}\geq 0,\qquad\phi(\mathsf{x})=0\text{ only if }\mathsf{x}=0,$
and $\phi(\sup\mathsf{x}_{i})=\lim\phi(\mathsf{x}_{i})$ for any bounded increasing net $\{\mathsf{x}_{i}\}$ of elements $\mathsf{x}_{i}\geq 0$ in $\mathcal{X}$. We will be working with $\mathcal{X}$-valued random variables, understood as strongly (i.e. Bochner) measurable functions from the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$ to the Banach space $(\mathcal{X},\|\cdot\|_{\mathrm{op}})$. For a $\mathcal{X}$-valued random variable $\mathsf{x}$ satisfying $\mathbb{E}\|\mathsf{x}\|_{\mathrm{op}}<\infty$, we denote by $\mathbb{E}\mathsf{x}\in\mathcal{X}$ its expectation and $\mathbb{E}[\mathsf{x}\mid\mathscr{G}]$ its conditional expectation with respect to a sub-sigma-field $\mathscr{G}\subset\mathscr{F}$ (c.f. [57, Chapter 1]).
3.1. Stochastic domination
For $p\in[1,\infty)$, we denote the $L^{p}$-norms with respect to $\phi$ as
$\|\mathsf{x}\|_{p}=\phi\big{(}|\mathsf{x}|^{p}\big{)}^{1/p},\qquad|\mathsf{x}|=(\mathsf{x}^{*}\mathsf{x})^{1/2}.$
Relevant properties of these norms and their associated non-commutative $L^{p}$-spaces are reviewed in Appendix D. Throughout, we will write $\mathsf{x}\prec\zeta$ to mean stochastic domination of the $L^{p}$-norm for each fixed $p\in[1,\infty)$, in the following sense.
Definition 3.1 (Stochastic domination).
Let $\mathsf{x}=\{\mathsf{x}(u):u\in U\}$ be a $n$-dependent family of $\mathcal{X}$-valued random variables, and $\zeta=\{\zeta(u):u\in U\}$ a corresponding family of (positive) scalar-valued random variables, where $U$ is a $n$-dependent parameter set. We say that
$\mathsf{x}\prec\zeta,\qquad\mathsf{x}=O_{\prec}(\zeta)$
uniformly over $u\in U_{n}$ if, for any fixed $p\in[1,\infty)$ and $\epsilon,D>0$, there exists $n_{0}(p,\epsilon,D)>0$ such that for all $n\geq n_{0}$,
$\sup_{u\in U_{n}}\mathbb{P}\left[\|\mathsf{x}(u)\|_{p}\geq n^{\epsilon}\zeta(u)\right]\leq n^{-D}.$
In the scalar setting of a $\mathbb{C}$-valued random variable $\mathsf{x}$, this means $\sup_{u\in U_{n}}\mathbb{P}\left[|\mathsf{x}(u)|\geq n^{\epsilon}\zeta(u)\right]\leq n^{-D}$ for some $n_{0}(\epsilon,D)>0$ and all $n\geq n_{0}(\epsilon,D)$.
We will often use implicitly the following basic properties of $\prec$.
Lemma 3.2.
  1. (a)
    If $\mathsf{x}(u,v)\prec\zeta(u,v)$ uniformly over $u\in U$ and $v\in V$, and $|V|\leq n^{C}$ for some constant $C>0$, then uniformly over $u\in U$,
    $\sum_{v\in V}\mathsf{x}(u,v)\prec\sum_{v\in V}\zeta(u,v)$
  2. (b)
    If $\mathsf{x}_{1}\prec\zeta_{1}$ and $\mathsf{x}_{2}\prec\zeta_{2}$ uniformly over $u\in U$, then also $\mathsf{x}_{1}\mathsf{x}_{2}\prec\zeta_{1}\zeta_{2}$ uniformly over $u\in U$.
  3. (c)
    Suppose $\mathsf{x}\prec\zeta$ uniformly over $u\in U$, where $\zeta$ is deterministic, $\zeta>n^{-C}$, and $\mathbb{E}[\|\mathsf{x}\|_{p}^{k}]\leq n^{C_{p,k}}$ for all $p,k\in[1,\infty)$ and some constants $C,C_{p,k}>0$. Then $\mathbb{E}[\mathsf{x}\mid\mathscr{G}]\prec\zeta$ uniformly over $u\in U$ and over all sub-sigma-fields $\mathscr{G}\subseteq\mathscr{F}$ of the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$.
Proof.
The argument is similar to the scalar setting (c.f. [28, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.3),
$\bigg{\|}\sum_{v\in V}\mathsf{x}(u,v)\bigg{\|}_{p}\leq\sum_{v\in V}\|\mathsf{x}(u,v)\|_{p},\qquad\|\mathsf{x}_{1}\mathsf{x}_{2}\|_{p}\leq\|\mathsf{x}_{1}\|_{2p}\|\mathsf{x}_{2}\|_{2p}.$
Statements (a) and (b) then follow from a union bound. For (c), by the given assumptions, for any $p,k\in[1,\infty)$, $\epsilon>0$, and all $n\geq n_{0}(\epsilon,k,p)$,
$\mathbb{E}[\|\mathsf{x}\|_{p}^{k}]$ $\leq\mathbb{E}[\|\mathsf{x}\|_{p}^{k}\mathbf{1}\{\|\mathsf{x}\|_{p}^{k}\leq n^{\epsilon/2}\zeta^{k}\}]+\mathbb{E}[\|\mathsf{x}\|_{p}^{k}\mathbf{1}\{\|\mathsf{x}\|_{p}^{k}>n^{\epsilon/2}\zeta^{k}\}]$
$\leq n^{\epsilon/2}\zeta^{k}+\mathbb{E}[\|\mathsf{x}\|_{p}^{2k}]^{1/2}\mathbb{P}[\|\mathsf{x}\|_{p}^{k}>n^{\epsilon/2}\zeta^{k}]^{1/2}<n^{\epsilon}\zeta^{k}.$ (36)
The triangle inequality for $\|\cdot\|_{p}$ implies for $\lambda\in[0,1]$ that $\|\lambda\mathsf{x}+(1-\lambda)\mathsf{y}\|_{p}\leq\lambda\|\mathsf{x}\|_{p}+(1-\lambda)\|\mathsf{y}\|_{p}$, so $\mathsf{x}\mapsto\|\mathsf{x}\|_{p}$ is continuous and convex. Then (c.f. [57, Proposition 1.12])
$\|\mathbb{E}[\mathsf{x}\mid\mathscr{G}]\|_{p}\leq\mathbb{E}[\|\mathsf{x}\|_{p}\mid\mathscr{G}].$ (37)
So for any $\mathscr{G}\subseteq\mathscr{F}$, $p\in[1,\infty)$, and $\epsilon,D>0$, fixing $k\geq 1$ such that $(k-1)\epsilon>D$ and choosing $n\geq n_{0}(p,\epsilon,D)$ large enough so that (36) holds,
$\mathbb{P}[\|\mathbb{E}[\mathsf{x}\mid\mathscr{G}]\|_{p}>n^{\epsilon}\zeta]\leq\frac{\mathbb{E}[\|\mathbb{E}[\mathsf{x}\mid\mathscr{G}]\|_{p}^{k}]}{n^{k\epsilon}\zeta^{k}}\leq\frac{\mathbb{E}[\mathbb{E}[\|\mathsf{x}\|_{p}\mid\mathscr{G}]^{k}]}{n^{k\epsilon}\zeta^{k}}\leq\frac{\mathbb{E}[\|\mathsf{x}\|_{p}^{k}]}{n^{k\epsilon}\zeta^{k}}<n^{-(k-1)\epsilon}<n^{-D}.$
Remark 3.3.
For the finite-dimensional matrix algebra $(\mathbb{C}^{n\times n},\|\cdot\|_{\mathrm{op}},\frac{1}{n}\operatorname{Tr})$, we have $\|X\|_{\mathrm{op}}\leq(\operatorname{Tr}(X^{*}X)^{p/2})^{1/p}=n^{1/p}\|X\|_{p}$. Thus if $X\prec\zeta$, then for any $\epsilon,D>0$ and all $n\geq n_{0}(\epsilon,D)$, choosing $p=\max(1,2/\epsilon)$,
$\mathbb{P}[\|X\|_{\mathrm{op}}\geq n^{\epsilon}\zeta]\leq\mathbb{P}[n^{1/p}\|X\|_{p}\geq n^{\epsilon}\zeta]\leq\mathbb{P}[\|X\|_{p}\geq n^{\epsilon/2}\zeta]<n^{-D}$
so this implies the operator norm bound $\|X\|_{\mathrm{op}}\prec\zeta$. However, we caution that this implication does not hold in infinite-dimensional settings, where our notation $\mathsf{x}\prec\zeta$ has a weaker meaning than $\|\mathsf{x}\|_{\mathrm{op}}\prec\zeta$.
3.2. Khintchine-type inequalities
The following statements may be deduced from the non-commutative Rosenthal inequalities of [38]. In the setting of Rademacher variables, similar Khintchine inequalities have been shown in [56, Eq. (8.4.11)] and [59, Theorem 6.22].
Lemma 3.4.
Let $(\mathsf{x}_{i},\mathsf{y}_{ij}:i,j=1,\ldots,n)$ be (deterministic) elements in $\mathcal{X}$, and denote
$\mathsf{Y}=\sum_{i,j=1}^{n}E_{ij}\otimes\mathsf{y}_{ij}\in\mathbb{C}^{n\times n}\otimes\mathcal{X},\qquad\mathsf{Y}^{\mathsf{t}}=\sum_{i,j=1}^{n}E_{ji}\otimes\mathsf{y}_{ij}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ (38)
where $\mathsf{Y}^{\mathsf{t}}$ is the partial transpose of $\mathsf{Y}$ in its first tensor factor. We equip $\mathbb{C}^{n\times n}\otimes\mathcal{X}$ with the trace $n^{-1}\operatorname{Tr}\otimes\phi$ and its $L^{p}$-norm $\|\mathsf{x}\|_{p}=((n^{-1}\operatorname{Tr}\otimes\phi)[|\mathsf{x}|^{p}])^{1/p}$.
Let $(\alpha_{i},\beta_{i}:i=1,\ldots,n)$ be independent $\mathbb{C}$-valued random variables, satisfying $\mathbb{E}\alpha_{i}=\mathbb{E}\beta_{i}=0$ and $\mathbb{E}[|\alpha_{i}|^{p}],\mathbb{E}[|\beta_{i}|^{p}]\leq C_{p}$ for every $p\geq 2$ and some constant $C_{p}>0$. Then for all $p\in[2,\infty)$, there are constants $C_{p}^{\prime},C_{p}^{\prime\prime},C_{p}^{\prime\prime\prime}>0$ such that
  1. (a)
    $\mathbb{E}\bigg{[}\bigg{\|}\sum_{i=1}^{n}\alpha_{i}\mathsf{x}_{i}\bigg{\|}_{p}^{p}\bigg{]}\leq C_{p}^{\prime}\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{\}}$
  2. (b)
    $\mathbb{E}\bigg{[}\bigg{\|}\sum_{i,j=1}^{n}\alpha_{i}\beta_{j}\mathsf{y}_{ij}\bigg{\|}_{p}^{p}\bigg{]}$ $\leq C_{p}^{\prime\prime}\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{i,j=1}^{n}\mathsf{y}_{ij}\mathsf{y}_{ij}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\bigg{\|}\bigg{(}\sum_{i,j=1}^{n}\mathsf{y}_{ij}^{*}\mathsf{y}_{ij}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},n\|\mathsf{Y}\|_{p}^{p},n\|\mathsf{Y}^{\mathsf{t}}\|_{p}^{p}\bigg{\}}$
  3. (c)
    Suppose $\mathsf{y}_{ii}=0$ for each $i=1,\ldots,n$. Then
    $\mathbb{E}\bigg{[}\bigg{\|}\sum_{1\leq i\neq j\leq n}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg{\|}_{p}^{p}\bigg{]}$ $\leq C_{p}^{\prime\prime\prime}\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{1\leq i\neq j\leq n}\mathsf{y}_{ij}\mathsf{y}_{ij}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\bigg{\|}\bigg{(}\sum_{1\leq i\neq j\leq n}\mathsf{y}_{ij}^{*}\mathsf{y}_{ij}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},n\|\mathsf{Y}\|_{p}^{p},n\|\mathsf{Y}^{\mathsf{t}}\|_{p}^{p}\bigg{\}}$
Proof.
See Appendix B. ∎
3.3. Fluctuation averaging
Let $\{\mathscr{G}_{i}:i=1,\ldots,n\}$ be a collection of sub-sigma-fields in the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$. For a $\mathcal{X}$-valued random variable $\mathsf{x}$ with $\mathbb{E}\|\mathsf{x}\|_{\mathrm{op}}<\infty$, define the projections
$\mathbb{E}_{i}[\mathsf{x}]=\mathbb{E}[\mathsf{x}\mid\mathscr{G}_{i}],\qquad\mathcal{Q}_{i}[\mathsf{x}]=(1-\mathbb{E}_{i})[\mathsf{x}]=\mathsf{x}-\mathbb{E}_{i}[\mathsf{x}].$
Supposing that $\{\mathbb{E}_{i},\mathcal{Q}_{i}:i=1,\ldots,n\}$ all commute, set
$\mathbb{E}_{S}=\prod_{i\in S}\mathbb{E}_{i},\qquad\mathcal{Q}_{S}=\prod_{i\in S}\mathcal{Q}_{i}$
with the conventions $\mathbb{E}_{\emptyset}[\mathsf{x}]=\mathcal{Q}_{\emptyset}[\mathsf{x}]=\mathsf{x}$.
The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [23, Theorem 4.7] and [29, Lemma A.2].
Lemma 3.5 (Fluctuation averaging).
Suppose that $\{\mathbb{E}_{i},\mathcal{Q}_{i}:i=1,\ldots,n\}$ commute. Let $\{\mathsf{x}_{i}\}_{i=1}^{n}$ and $\{\mathsf{x}_{ij}\}_{1\leq i\neq j\leq n}$ be $\mathcal{X}$-valued random variables such that for any $p,k\in[1,\infty)$ and some constants $C_{p,k}>0$, we have $\mathbb{E}[\|\mathsf{x}_{i}\|_{p}^{k}],\mathbb{E}[\|\mathsf{x}_{ij}\|_{p}^{k}]\leq n^{C_{p,k}}$ for all $i\neq j$.
  1. (a)
    Suppose $\mathbb{E}_{i}[\mathsf{x}_{i}]=0$, and for some $\alpha,\beta>0$ and each fixed $l\geq 1$, uniformly over $S\subseteq\{1,\ldots,n\}$ with $|S|\leq l$ and over $i\notin S$, we have
    $\left\|\mathcal{Q}_{S}[\mathsf{x}_{i}]\right\|_{l}\prec n^{-\alpha-\beta|S|}.$ (39)
    Denote $\beta^{\prime}=\min\{1/2,\beta\}$. Then uniformly over deterministic vectors $\mathbf{u}=(u_{1},\ldots,u_{n})\in\mathbb{C}^{n}$,
    $\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\prec n^{-\alpha-\beta^{\prime}}(n\|\mathbf{u}\|_{\infty})$
  2. (b)
    Suppose $\mathbb{E}_{i}[\mathsf{x}_{i}]=0$, and for some $\alpha>0$ and each fixed $l\geq 1$, uniformly over $S\subseteq\{1,\ldots,n\}$ with $|S|\leq l$ and over $i\notin S$, we have
    $\left\|\mathcal{Q}_{S}[\mathsf{x}_{i}]\right\|_{l}\prec n^{-\alpha-|S|/2}.$ (40)
    Then uniformly over deterministic vectors $\mathbf{u}=(u_{1},\ldots,u_{n})\in\mathbb{C}^{n}$,
    $\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\prec n^{-\alpha}\|\mathbf{u}\|_{2}$
  3. (c)
    Suppose $\mathbb{E}_{i}[\mathsf{x}_{ij}]=\mathbb{E}_{j}[\mathsf{x}_{ij}]=0$, and for some $\alpha>0$ and each fixed $l\geq 1$, uniformly over $S\subseteq\{1,\ldots,n\}$ with $|S|\leq l$ and over $i,j\notin S$ with $i\neq j$, we have
    $\left\|\mathcal{Q}_{S}[\mathsf{x}_{ij}]\right\|_{l}\prec n^{-\alpha-|S|/2}.$ (41)
    Then uniformly over $(u_{ij})_{i,j=1}^{n}\in\mathbb{C}^{n\times n}$,
    $\sum_{i\neq j}u_{ij}\mathsf{x}_{ij}\prec n^{-\alpha}\left(\sum_{i\neq j}|u_{ij}|^{2}\right)^{1/2}$
Proof.
See Appendix C. ∎
3.4. Minors and resolvent identities
Let $H\in\mathbb{C}^{n\times n}$ and $\mathsf{x}\in\mathcal{X}$ be self-adjoint, and let
$\mathsf{z}=\sum_{i=1}^{n}E_{ii}\otimes\mathsf{z}_{i}\in\mathcal{D}\otimes\mathcal{X},\qquad\mathsf{z}_{i}\in\mathcal{X}^{+}=\{\mathsf{x}\in\mathcal{X}:\Im\mathsf{x}\geq\epsilon\text{ for some }\epsilon>0\}.$
Consider the generalized resolvent
$R=(H\otimes\mathsf{x}-\mathsf{z})^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X},$
which exists by Lemma D.1 as ${-}\Im(H\otimes\mathsf{x}-\mathsf{z})=\Im\mathsf{z}\geq\epsilon$ if $\Im\mathsf{z}_{i}\geq\epsilon$ for every $i$. For an index set $S\subseteq\{1,\ldots,n\}$, we define $H^{(S)}\in\mathbb{C}^{n\times n}$ and $R^{(S)}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ by
$H_{ij}^{(S)}=\begin{cases}H_{ij}&\text{ if }i,j\notin S\\ 0&\text{ otherwise}\end{cases},\qquad R^{(S)}=(H^{(S)}\otimes\mathsf{x}-\mathsf{z})^{-1},$
with $H^{(\emptyset)}=H$ and $R^{(\emptyset)}=R$. We will use the indexing $R_{ij}=(\mathbf{e}_{i}\otimes 1)^{*}R(\mathbf{e}_{j}\otimes 1)$ for the $\mathcal{X}$-valued entries of $R$, and the notations
$\sum_{i}^{(S)}=\sum_{i\in\{1,...,n\}\setminus S},\qquad\frac{1}{R_{ii}^{(S)}}=(R_{ii}^{(S)})^{-1},\qquad iS=S\cup\{i\}.$
Lemma 3.6 (Resolvent identities).
Suppose $H\in\mathbb{C}^{n\times n}$ and $\mathsf{x}\in\mathcal{X}$ are self-adjoint, and $\mathsf{z}_{i}\in\mathcal{X}^{+}$ for each $i=1,\ldots,n$. For any $S\subseteq\{1,\ldots,n\}$:
  1. (a)
    For any $i\notin S$, $R_{ii}^{(S)}\in\mathcal{X}$ is invertible, and
    $\frac{1}{R_{ii}^{(S)}}=h_{ii}\mathsf{x}-\mathsf{z}_{i}-\mathsf{x}\left(\sum_{r,s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}\right)\mathsf{x}$
  2. (b)
    For any $i,j\notin S$ with $i\neq j$,
    $R_{ij}^{(S)}$ $={-}R_{ii}^{(S)}\mathsf{x}\left(\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\right)={-}\left(\sum_{s}^{(jS)}R_{is}^{(jS)}h_{sj}\right)\mathsf{x}R_{jj}^{(S)}$
    $={-}h_{ij}R_{ii}^{(S)}\mathsf{x}R_{jj}^{(iS)}+R_{ii}^{(S)}\mathsf{x}\left(\sum_{r,s}^{(ijS)}h_{ir}R_{rs}^{(ijS)}h_{sj}\right)\mathsf{x}R_{jj}^{(iS)}.$
  3. (c)
    For any $i,j,r\notin S$ (including $i=j$) with $r\notin\{i,j\}$,
    $R_{ij}^{(S)}$ $=R_{ij}^{(rS)}+R_{ir}^{(S)}\frac{1}{R_{rr}^{(S)}}R_{rj}^{(S)},$
    $\frac{1}{R_{ii}^{(S)}}$ $=\frac{1}{R_{ii}^{(rS)}}-\frac{1}{R_{ii}^{(S)}}R_{ir}^{(S)}\frac{1}{R_{rr}^{(S)}}R_{ri}^{(S)}\frac{1}{R_{ii}^{(rS)}}.$
Proof.
These follow from Schur-complement inversion identities applied to
$H^{(S)}-\mathsf{z}=\begin{pmatrix}h_{11}^{(S)}\mathsf{x}-\mathsf{z}_{1}&h_{12}^{(S)}\mathsf{x}&\cdots&h_{1n}^{(S)}\mathsf{x}\\ h_{21}^{(S)}\mathsf{x}&h_{22}^{(S)}\mathsf{x}-\mathsf{z}_{2}&\cdots&h_{2n}^{(S)}\mathsf{x}\\ \vdots&\vdots&\ddots&\vdots\\ h_{n1}^{(S)}\mathsf{x}&h_{n2}^{(S)}\mathsf{x}&\cdots&h_{nn}^{(S)}\mathsf{x}-\mathsf{z}_{n}\end{pmatrix}\in\mathbb{C}^{n\times n}\otimes\mathcal{X},$
which are purely algebraic and identical to those of the scalar setting, see e.g. [26, Lemma 4.2]. ∎
3.5. Maximum modulus principle
We will use the following quantitative version of the maximum modulus principle, following [60, Appendix A].
Lemma 3.7.
Fix any $a>0$. For each $r\in(-\infty,0)$, define the circle in $\mathbb{C}^{+}$
$S_{r}=\{z\in\mathbb{C}^{+}:|z-ia\tfrac{1+e^{2r}}{1-e^{2r}}|=\tfrac{2ae^{r}}{1-e^{2r}}\}.$
Let $f:\mathbb{C}^{+}\to\mathbb{C}$ be any analytic function not identically equal to 0, and set
$M(r)=\max_{z\in S_{r}}\log|f(z)|.$
Then $r\mapsto M(r)$ is increasing and convex over $r\in(-\infty,0)$.
Proof.
For any analytic function $g:\mathbb{D}\to\mathbb{C}$ (not identically 0) on the unit disk $\mathbb{D}=\{z:|z|<1\}$, the function
$r\mapsto\max_{z\in\mathbb{D}:|z|=e^{r}}\log|g(z)|$
is increasing and convex [35]. Fixing $a>0$, consider the conformal mapping $\psi_{a}:\mathbb{D}\rightarrow\mathbb{C}^{+}$ given by $\psi_{a}(z)=ia\frac{1+z}{1-z}$. For any $z\in\mathbb{D}$,
$\left|\psi_{a}(z)-ia\frac{1+|z|^{2}}{1-|z|^{2}}\right|=\left|\frac{2a(z-|z|^{2})}{(1-z)(1-|z|^{2})}\right|=\frac{2a|z|}{1-|z|^{2}}\left|\frac{1-\bar{z}}{1-z}\right|=\frac{2a|z|}{1-|z|^{2}},$
so $\psi_{a}$ maps each circle $\{z\in\mathbb{D}:|z|=e^{r}\}$ bijectively to $S_{r}\subset\mathbb{C}^{+}$. The result then follows from applying the above monotonicity and convexity to the function $g(z)=f(\psi_{a}(z))$. ∎
4. Analysis of a generalized resolvent
Much of the analysis for Theorem 2.4 may be stated at the level of a generalized resolvent
$R=(H\otimes\mathsf{x}-\mathsf{z})^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X},\qquad\mathsf{z}=\sum_{k=1}^{K}D_{k}\otimes\mathsf{x}_{k}\in\mathcal{D}\otimes\mathcal{X}$ (42)
where $\mathsf{x},\mathsf{x}_{1},\ldots,\mathsf{x}_{K}$ are elements of an abstract von Neumann algebra $\mathcal{X}$, and $H$ is a Wigner random matrix. We will then specialize to $\mathcal{X}=\mathbb{C}^{n\times n}$ and $\mathcal{X}=\mathcal{A}$ for the two stages of analysis of the Kronecker deformed Wigner model (17). We emphasize that it is important for our arguments to hold when $\mathcal{X}$ is infinite-dimensional, for the analysis in the second stage with $\mathcal{X}=\mathcal{A}$.
We collect here the assumptions of this general setting.
Assumption 4.1.
There exist constants $\gamma\geq 1$, $\upsilon\geq\delta>0$, and $K\geq 1$ such that:
  1. (a)
    $H\in\mathbb{C}^{n\times n}$ is a random Wigner matrix satisfying the conditions of Assumption 2.1, and $D_{1},\ldots,D_{K}\in\mathbb{C}^{n\times n}$ are deterministic and diagonal matrices.
  2. (b)
    $(\mathcal{A},\|\cdot\|_{\mathrm{op}},\tau)$ is the von Neumann algebra of Assumption 2.2 with trace $\tau$, containing the subalgebra $\mathcal{M}\equiv\mathbb{C}^{n\times n}$ and a semicircular element, denoted here as $\mathsf{h}$, that is free of $\mathcal{M}$.
  3. (c)
    $(\mathcal{X},\|\cdot\|_{\mathrm{op}},\phi)$ is a von Neumann algebra with faithful, normal, tracial state $\phi:\mathcal{X}\to\mathbb{C}$, and $\mathsf{x}\in\mathcal{X}$ is self-adjoint and invertible.
  4. (d)
    We have
    $\|\mathsf{x}\|_{\mathrm{op}},\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma,\quad\|\mathsf{z}\|_{\mathrm{op}}\leq\sum_{k=1}^{K}\|D_{k}\|_{\mathrm{op}}\|\mathsf{x}_{k}\|_{\mathrm{op}}\leq\upsilon,\quad\Im\mathsf{z}=\Im\sum_{k=1}^{K}D_{k}\otimes\mathsf{x}_{k}\geq\delta.$ (43)
We denote by $\mathcal{D}\subset\mathcal{M}$ the subalgebra of diagonal matrices, and $\tau^{\mathcal{D}}:\mathcal{A}\to\mathcal{D}$ the diagonal projection (18) onto $\mathcal{D}$. We remark that under this assumption, $\mathsf{z}$ admits an equivalent representation
$\mathsf{z}=\sum_{i=1}^{n}E_{ii}\otimes\mathsf{z}_{i},\qquad\mathsf{z}_{i}=\sum_{k=1}^{K}[D_{k}]_{ii}\mathsf{x}_{k}\in\mathcal{X}^{+},\qquad\Im\mathsf{z}_{i}\geq\delta,$
and we will use these representations of $\mathsf{z}$ interchangeably. We denote the limiting operator for the generalized resolvent $R$ by
$\mathsf{r}=\left(\mathsf{h}\otimes\mathsf{x}-\mathsf{z}\right)^{-1}\in\mathcal{A}\otimes\mathcal{X},$
its projection under $\tau^{\mathcal{D}}\otimes 1$ (the $\mathcal{D}\otimes\mathcal{X}$-valued Stieltjes transform of $\mathsf{h}\otimes\mathsf{x}$ evaluated at $\mathsf{z}$) by
$\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[\mathsf{r}]\in\mathcal{D}\otimes\mathcal{X},$
and its projection under $\tau\otimes 1$ by
$\mathsf{m}_{0}=(\tau\otimes 1)[\mathsf{r}]=(n^{-1}\operatorname{Tr}\otimes 1)[\mathsf{r}_{0}]\in\mathcal{X}.$
We use the indexing $R_{ij}=(\mathbf{e}_{i}\otimes 1)^{*}R(\mathbf{e}_{j}\otimes 1)$ and $(\mathsf{r}_{0})_{ij}=(\mathbf{e}_{i}\otimes 1)^{*}\mathsf{r}_{0}(\mathbf{e}_{j}\otimes 1)$, and write $\mathsf{x}\prec\zeta$ or $\mathsf{x}=O_{\prec}(\zeta)$ for stochastic domination in the $L^{p}$-norms for $p\in[1,\infty)$ as discussed in Section 3.1.
Our main result in this context is the following theorem.
Theorem 4.2.
Uniformly over $\mathsf{x},\mathsf{z}$ satisfying Assumption 4.1,
$(n^{-1}\operatorname{Tr}\,\otimes 1)[R]-\mathsf{m}_{0}\prec n^{-1}.$ (44)
Also uniformly over $i\neq j\in\{1,\ldots,n\}$,
$R_{ii}-(\mathsf{r}_{0})_{ii}$ $\prec n^{-1/2},$ (45)
$R_{ij}$ $\prec n^{-1/2},$ (46)
and uniformly over deterministic vectors $\mathbf{u},\mathbf{v}\in\mathbb{C}^{n}$ with $\|\mathbf{u}\|_{2},\|\mathbf{v}\|_{2}\leq\upsilon$,
$(\mathbf{u}\otimes 1)^{*}[R-\mathsf{r}_{0}](\mathbf{v}\otimes 1)\prec n^{-1/2}.$ (47)
In the remainder of this section, we prove Theorem 4.2. Section 4.1 discusses existence, uniqueness, and stability of the solution to the relevant operator-valued fixed-point equation. Section 4.2 proves a preliminary estimate in operator norm
$\|R_{ii}-(\mathsf{r}_{0})_{ii}\|_{\mathrm{op}}\prec n^{-\alpha},\qquad\|R_{ij}\|_{\mathrm{op}}\prec n^{-\alpha}\text{ for }i\neq j$
for some $\alpha>0$, by conducting the analysis for sufficiently large $|z|$ and extending the result to all fixed $z\in\mathbb{C}^{+}$ using the maximum modulus principle. Section 4.3 improves this to the optimal estimate of $n^{-1/2}$ in the $L^{p}$-norms for $p\in[1,\infty)$, using the non-commutative Khintchine-type inequalities, fluctuation averaging techniques, and an iterative bootstrapping argument. Section 4.4 concludes the proof of Theorem 4.2.
All stochastic domination statements of this section are implicitly uniform over $\mathsf{x},\mathsf{z}$ satisfying Assumption 4.1, indices $i,j,k,\ldots\in\{1,\ldots,n\}$, subsets $S\subset\{1,\ldots,n\}$ having cardinality $|S|\leq l$ for any fixed ($n$-independent) value $l\geq 1$, and unit vectors $\mathbf{u},\mathbf{v}\in\mathbb{C}^{n}$ satisfying $\|\mathbf{u}\|_{2},\|\mathbf{v}\|_{2}\leq\upsilon$.
4.1. Operator fixed-point equations
Lemma 4.3.
Let $\mathcal{D}\subset\mathcal{M}\subset\mathcal{A}$ and $\mathcal{X}$ be as in Assumption 4.1. Let $\mathsf{x}\in\mathcal{X}$ be self-adjoint, and let $\mathsf{h}\in\mathcal{A}$ be a semicircular element free of $\mathcal{M}$.
  1. (a)
    ($\mathcal{D}\otimes\mathcal{X}$-valued fixed point) For any $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}=\{\mathsf{z}\in\mathcal{D}\otimes\mathcal{X}:\Im\mathsf{z}\geq\epsilon\text{ for some }\epsilon>0\}$, there exists a unique solution $\mathsf{s}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ to the fixed-point equation
    $\mathsf{s}=\left({-}\mathsf{z}-I_{n\times n}\otimes[\mathsf{x}\left(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]\right)\mathsf{x}]\right)^{-1}.$ (48)
    This solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$.
  2. (b)
    ($\mathcal{X}$-valued fixed point) For any $\mathsf{z}_{1},\ldots,\mathsf{z}_{n}\in\mathcal{X}^{+}$, there exists a unique solution $\mathsf{m}\in\mathcal{X}^{+}$ to the fixed-point equation
    $\mathsf{m}=\frac{1}{n}\sum_{i=1}^{n}\left({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}\mathsf{x}\right)^{-1}.$ (49)
    Setting $\mathsf{z}=\sum_{i=1}^{n}E_{ii}\otimes\mathsf{z}_{i}\in\mathcal{D}\otimes\mathcal{X}$, this solution is given by $\mathsf{m}_{0}=(\tau\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$.
Proof.
For existence and uniqueness in part (a), we apply the general result of [37, Theorem 2.1]: Let $\mathcal{B}$ be any $C^{*}$-algebra and denote its right operator half-plane
=$\mathcal{B}^{\prime}=\{\mathsf{a}\in\mathcal{B}:\Re\mathsf{a}=\tfrac{1}{2}(\mathsf{a}+\mathsf{a}^{*})\geq\epsilon\text{ for some }\epsilon>0\}.$
Then for any $\mathsf{b}\in\mathcal{B}^{\prime}$ and any analytic mapping $\eta:\mathcal{B}^{\prime}\to\mathcal{B}^{\prime}$ that is bounded on bounded domains of $\mathcal{B}^{\prime}$, there exists a unique solution $\mathsf{s}\in\mathcal{B}^{\prime}$ to the equation $\mathsf{s}=[\mathsf{b}+\eta(\mathsf{s})]^{-1}$.
Fixing $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, we take $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$, $\mathsf{b}={-}i\mathsf{z}/2$, and $\eta(\mathsf{s})={-}i\mathsf{z}/2+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$, noting that $\Re[\mathsf{b}]=\Im\mathsf{z}/2\geq\epsilon>0$ and $\Re\eta(\mathsf{s})=\Im z/2+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\Re s])\mathsf{x}\geq\epsilon>0$ for some $\epsilon>0$ whenever $\Re s\geq 0$ and $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, because $\mathsf{x}$ is self-adjoint and $n^{-1}\operatorname{Tr}\otimes 1$ is a positive map (Lemma D.2). Thus, there exists a unique solution $\mathsf{s}^{\prime}\in(\mathcal{D}\otimes\mathcal{X})^{\prime}$ to
$\mathsf{s}^{\prime}=\big{(}{-}i\mathsf{z}+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}^{\prime}])\mathsf{x}\big{)}^{-1}.$
Multiplying by $i$, there exists a unique solution $\mathsf{s}=i\mathsf{s}^{\prime}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ to (48).
To show that this solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$, we claim that the map $\mathsf{s}\mapsto I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$ is the $\mathcal{D}\otimes\mathcal{X}$-valued R-transform of $\mathsf{h}\otimes\mathsf{x}$, and (48) is the Cauchy-R relation over $\mathcal{D}\otimes\mathcal{X}$. We recall a few relevant notions from operator-valued free probability theory [47, Chapter 9.2]: Let $\operatorname{NC}(m)$ be the lattice of non-crossing partitions of $\{1,\ldots,m\}$. Associated to any von Neumann subalgebra $\mathcal{B}$ of $\mathcal{Y}=\mathcal{A}\otimes\mathcal{X}$ and its conditional expectation $\tau^{\mathcal{B}}:\mathcal{Y}\to\mathcal{B}$ (c.f. Lemma D.2) is a system of $\mathcal{B}$-valued cumulants $(\kappa_{\pi}^{\mathcal{B}})_{\pi\in\operatorname{NC}(m)}$, which are $\mathbb{C}$-multilinear maps $\kappa_{\pi}^{\mathcal{B}}:\mathcal{Y}^{m}\to\mathcal{B}$ satisfying the free moment-cumulant relation
$\tau^{\mathcal{B}}(\mathsf{y}_{1}\mathsf{y}_{2}\ldots\mathsf{y}_{m})=\sum_{\pi\in\operatorname{NC}(m)}\kappa_{\pi}^{\mathcal{B}}(\mathsf{y}_{1},\mathsf{y}_{2},\ldots,\mathsf{y}_{m})$ (50)
and the bimodule properties
$\kappa_{\pi}^{\mathcal{B}}(\mathsf{b}\mathsf{y}_{1},\mathsf{y}_{2},\ldots,\mathsf{y}_{m-1},\mathsf{y}_{m}\mathsf{b}^{\prime})$ $=\mathsf{b}\kappa_{\pi}^{\mathcal{B}}(\mathsf{y}_{1},\ldots,\mathsf{y}_{m-1},\mathsf{y}_{m})\mathsf{b}^{\prime},$ (51)
$\kappa_{\pi}^{\mathcal{B}}(\mathsf{y}_{1},\ldots,\mathsf{y}_{r}\mathsf{b},\mathsf{y}_{r+1},\ldots,\mathsf{y}_{m})$ $=\kappa_{\pi}^{\mathcal{B}}(\mathsf{y}_{1},\ldots,\mathsf{y}_{r},\mathsf{b}\mathsf{y}_{r+1},\ldots,\mathsf{y}_{m})$
for any $\mathsf{y}_{1},\ldots,\mathsf{y}_{m}\in\mathcal{Y}$ and $\mathsf{b},\mathsf{b}^{\prime}\in\mathcal{B}$. Fixing a self-adjoint element $\mathsf{y}\in\mathcal{Y}$, for any invertible $\mathsf{b}\in\mathcal{B}$ with $\|\mathsf{b}^{-1}\|_{\mathrm{op}}$ small enough, we may define the $\mathcal{B}$-valued Cauchy-transform
$G_{\mathsf{y}}^{\mathcal{B}}(\mathsf{b})=\tau^{\mathcal{B}}[(\mathsf{b}-\mathsf{y})^{-1}]=\sum_{m\geq 0}\tau^{\mathcal{B}}[\mathsf{b}^{-1}(\mathsf{y}\mathsf{b}^{-1})^{m}]\in\mathcal{B},$
and for any $\mathsf{b}\in\mathcal{B}$ with $\|\mathsf{b}\|_{\mathrm{op}}$ small enough, we may define the $\mathcal{B}$-valued R-transform
$\mathcal{R}_{\mathsf{y}}^{\mathcal{B}}(\mathsf{b})=\sum_{m\geq 1}\kappa_{m}^{\mathcal{B}}(\mathsf{y}\mathsf{b},\ldots,\mathsf{y}\mathsf{b},\mathsf{y})\in\mathcal{B}$
where $\kappa_{m}=\kappa_{\pi}$ for the partition $\pi=\{\{1,\ldots,m\}\}$. Then, for any invertible $\mathsf{b}\in\mathcal{B}$ with $\|\mathsf{b}^{-1}\|_{\mathrm{op}}$ small enough, these transforms satisfy the Cauchy-R relation
$G_{\mathsf{y}}^{\mathcal{B}}(\mathsf{b})=(\mathsf{b}-\mathcal{R}_{\mathsf{y}}^{\mathcal{B}}(G_{\mathsf{y}}^{\mathcal{B}}(\mathsf{b})))^{-1}.$ (52)
Specializing to our setting with $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$, we claim that for any $\mathsf{s}\in\mathcal{B}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
$\mathcal{R}_{\mathsf{h}\otimes\mathsf{x}}^{\mathcal{D}\otimes\mathcal{X}}(\mathsf{s})=I_{n\times n}\otimes\mathsf{x}\big{(}n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]\big{)}\mathsf{x}.$ (53)
Indeed, since $\mathsf{h}$ is free of $\mathcal{M}$ and hence of $\mathcal{D}\subset\mathcal{M}$, it is readily checked by definition (c.f. [47, Definition 9.5]) that $\mathsf{h}\otimes\mathsf{x}$ is free of $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$ with amalgamation over $1\otimes\mathcal{X}$ under the conditional expectation $\tau\otimes 1:\mathcal{A}\otimes\mathcal{X}\to\mathcal{X}\cong 1\otimes\mathcal{X}$. Then by [50, Theorem 3.6], for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
$\mathcal{R}_{\mathsf{h}\otimes\mathsf{x}}^{\mathcal{D}\otimes\mathcal{X}}(\mathsf{s})$ $=\sum_{m\geq 1}\kappa_{m}^{\mathcal{D}\otimes\mathcal{X}}((\mathsf{h}\otimes\mathsf{x})\mathsf{s},\ldots,(\mathsf{h}\otimes\mathsf{x})\mathsf{s},\mathsf{h}\otimes\mathsf{x})$
$=\sum_{m\geq 1}\kappa_{m}^{1\otimes\mathcal{X}}((\mathsf{h}\otimes\mathsf{x})(\tau\otimes 1)[\mathsf{s}],\ldots,(\mathsf{h}\otimes\mathsf{x})(\tau\otimes 1)[\mathsf{s}],\mathsf{h}\otimes\mathsf{x})$
$=\sum_{m\geq 1}\left(1\otimes\left[\mathsf{x}\left(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]\right)\right]^{m-1}\mathsf{x}\right)\kappa_{m}^{1\otimes\mathcal{X}}(\mathsf{h}\otimes 1,...,\mathsf{h}\otimes 1)$
where we used in the last step that $\tau\otimes 1[\mathsf{s}]=n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]$ for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$, that $\mathsf{h}\otimes 1$ commutes with $1\otimes\mathcal{X}$, and the bimodule properties (51). On the other hand, we have
$\kappa_{m}^{1\otimes\mathcal{X}}(\mathsf{h}\otimes 1,\ldots,\mathsf{h}\otimes 1)=\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})(1\otimes 1)$
where $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$ are the scalar-valued free cumulants of $\mathsf{h}$: This may be verified by expressing $\kappa_{m}^{1\otimes\mathcal{X}}(\mathsf{h}\otimes 1,\ldots,\mathsf{h}\otimes 1)$ in terms of moments under the conditional expectation $\tau\otimes 1$ using Möbius-inversion of (50) (c.f. [47, Eq. (9.19)]), applying the identity $(\tau\otimes 1)[\mathsf{a}\otimes 1]=\tau(\mathsf{a})(1\otimes 1)$ for each moment term, and re-applying (50) to express the result back in terms of the scalar cumulants $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$. Here, $\mathsf{h}\in\mathcal{A}$ is semi-circular, so $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})=1$ if $m=2$ and 0 otherwise (c.f. [51, Example 11.21]). Thus only the $m=2$ term above remains, and we obtain the claim (53).
Then, identifying $\mathsf{r}_{0}(\mathsf{z})=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]={-}G_{\mathsf{h}\otimes\mathsf{x}}^{\mathcal{D}\otimes\mathcal{X}}(\mathsf{z})$, the Cauchy-R relation (52) implies that
$\mathsf{r}_{0}(\mathsf{z})=\left({-}\mathsf{z}-I\otimes\mathsf{x}\left(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{r}_{0}(\mathsf{z})]\right)\mathsf{x}\right)^{-1}$
for all $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ with $\|\mathsf{z}^{-1}\|_{\mathrm{op}}$ small enough. Since the two sides of this equation are analytic over the operator half-plane $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ and are equal on an open subset of $(\mathcal{D}\otimes\mathcal{X})^{+}$, it follows from the identity principle they must be equal for all $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, showing part (a) that $\mathsf{r}_{0}$ is the unique solution of (48) for any such $\mathsf{z}$.
For part (b), observe that for any $\mathsf{m}\in\mathcal{X}^{+}$,
$(\mathsf{z}+I_{n\times n}\otimes\mathsf{x}\mathsf{m}\mathsf{x})^{-1}=\left(\sum_{i=1}^{n}E_{ii}\otimes(\mathsf{z}_{i}+\mathsf{x}\mathsf{m}\mathsf{x})\right)^{-1}=\sum_{i=1}^{n}E_{ii}\otimes(\mathsf{z}_{i}+\mathsf{x}\mathsf{m}\mathsf{x})^{-1}.$
Then setting $\mathsf{m}_{0}=n^{-1}\operatorname{Tr}\otimes 1[\mathsf{r}_{0}]=(\tau\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$ and taking $n^{-1}\operatorname{Tr}\otimes 1$ on both sides of (48) with $\mathsf{s}=\mathsf{r}_{0}$ shows that $\mathsf{m}_{0}$ solves (49). To see that $\mathsf{m}_{0}$ is the unique solution, observe that if $\mathsf{m}\in\mathcal{X}^{+}$ solves (49), then defining $\mathsf{s}_{i}=(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}\mathsf{x})^{-1}$ and $\mathsf{s}=\sum_{i}E_{ii}\otimes\mathsf{s}_{i}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, it follows from (49) that $\mathsf{m}=n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]$, so $\mathsf{s}$ solves (48). Then by the uniqueness claim of part (a), we must have $\mathsf{s}=\mathsf{r}_{0}$, so $\mathsf{m}=n^{-1}\operatorname{Tr}\otimes 1[\mathsf{r}_{0}]=\mathsf{m}_{0}$. Hence this solution $\mathsf{m}_{0}\in\mathcal{X}^{+}$ to (49) is unique. ∎
We deduce from the above the following stability statements for approximate solutions of these fixed-point equations.
Corollary 4.4.
Under Assumption 4.1:
  1. (a)
    Suppose $\mathsf{s}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ and $\Delta\in\mathcal{D}\otimes\mathcal{X}$ satisfy $\Im(\mathsf{z}+\Delta)\geq\delta/2$ and
    $\mathsf{s}=\left(-\mathsf{z}-\Delta-I_{n\times n}\otimes[\mathsf{x}\left(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]\right)\mathsf{x}]\right)^{-1}.$
    Then for any $p\in[1,\infty]$ (where $\|\cdot\|_{\infty}=\|\cdot\|_{\mathrm{op}}$),
    $\left\|\mathsf{s}-\mathsf{r}_{0}\right\|_{p}\leq 2\delta^{-2}\left\|\Delta\right\|_{p}.$
  2. (b)
    Suppose $\mathsf{m}\in\mathcal{X}^{+}$ and $\Delta\in\mathcal{X}$ satisfy $\Im(\mathsf{z}_{i}+\mathsf{x}\Delta\mathsf{x})\geq\delta/2$ for all $i=1,\ldots,n$ and
    $\mathsf{m}=\frac{1}{n}\sum_{i=1}^{n}\left(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}\mathsf{x}\right)^{-1}+\Delta.$ (54)
    Then for any $p\in[1,\infty]$,
    $\left\|\mathsf{m}-\mathsf{m}_{0}\right\|_{p}\leq(1+2\gamma^{2}\delta^{-2})\left\|\Delta\right\|_{p}$
Proof.
For part (a), defining $\mathsf{r}(\mathsf{z})=(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}$ and $\mathsf{r}_{0}(\mathsf{z})=(\tau^{\mathcal{D}}\otimes 1)[\mathsf{r}(\mathsf{z})]$, Lemma 4.3(a) implies that $\mathsf{s}=\mathsf{r}_{0}(\mathsf{z}+\Delta)$. Hence
$\left\|\mathsf{s}-\mathsf{r}_{0}\right\|_{p}$ $=\left\|\mathsf{r}_{0}(\mathsf{z}+\Delta)-\mathsf{r}_{0}(\mathsf{z})\right\|_{p}=\left\|(\tau^{\mathcal{D}}\otimes 1)[\mathsf{r}(\mathsf{z}+\Delta)\Delta\mathsf{r}(\mathsf{z})]\right\|_{p}$
$\leq\|\mathsf{r}(\mathsf{z}+\Delta)\|_{\mathrm{op}}\|\Delta\|_{p}\|\mathsf{r}(\mathsf{z})\|_{\mathrm{op}}\leq 2\delta^{-2}\left\|\Delta\right\|_{p}$
where we used $L^{p}$-contractivity of $\tau^{\mathcal{D}}\otimes 1$ (Lemma D.4), Hölder’s inequality (Lemma D.3), the given conditions $\Im\mathsf{z}\geq\delta$ and $\Im(\mathsf{z}+\Delta)\geq\delta/2$, and Lemma D.1.
Similarly for part (b), defining $\mathsf{m}_{0}(\mathsf{z})=(n^{-1}\operatorname{Tr}\otimes 1)[\mathsf{r}(\mathsf{z})]$ and setting $\mathsf{m}^{\prime}=\mathsf{m}-\Delta$, we have $\mathsf{m}^{\prime}=n^{-1}\sum_{i}({-}\mathsf{z}_{i}-\mathsf{x}\Delta\mathsf{x}-\mathsf{x}\mathsf{m}^{\prime}\mathsf{x})^{-1}$ so Lemma 4.3(a) implies that $\mathsf{m}^{\prime}=\mathsf{m}_{0}(\mathsf{z}+I\otimes\mathsf{x}\Delta\mathsf{x})$. Then, using also $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$,
$\|\mathsf{m}-\mathsf{m}_{0}\|_{p}$ $\leq\|\Delta\|_{p}+\|\mathsf{m}_{0}(\mathsf{z}+I\otimes\mathsf{x}\Delta\mathsf{x})-\mathsf{m}_{0}(\mathsf{z})\|_{p}$
$\leq\|\Delta\|_{p}+2\delta^{-2}\|I\otimes\mathsf{x}\Delta\mathsf{x}\|_{p}\leq(1+2\gamma^{2}\delta^{-2})\|\Delta\|_{p}.$
We close this section with an analysis of the quantitative invertibility of two linear operators $\mathcal{L}_{1},\mathcal{L}_{2}:\mathcal{X}\to\mathcal{X}$ in the $L^{p}$-norms, defined as
$\mathcal{L}_{1}(\mathsf{a})=\mathsf{a}-\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii},\qquad\mathcal{L}_{2}(\mathsf{a})=\mathsf{a}-\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}.$
The invertibility of $\mathcal{L}_{1}$ is more immediate, and follows from differentiating the preceding fixed-point equation. We state this in the following lemma.
Lemma 4.5.
Under Assumption 4.1, consider the linear operator $\mathcal{L}_{1}:\mathcal{X}\to\mathcal{X}$ given by
$\mathcal{L}_{1}(\mathsf{a})=\mathsf{a}-\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}.$
Then $\mathcal{L}_{1}$ is invertible, and for any $\mathsf{a}\in\mathcal{X}$ and $p\in[1,\infty]$ (where $\|\cdot\|_{\infty}=\|\cdot\|_{\mathrm{op}}$),
$\|\mathcal{L}_{1}^{-1}(\mathsf{a})\|_{p}\leq\gamma^{2}(\gamma^{2}+\delta^{-2})\|\mathsf{a}\|_{p}.$
Proof.
For any $\mathsf{b}\in\mathcal{X}$ with $\|\mathsf{b}\|_{\mathrm{op}}<\delta/2$ (which ensures $\Im(\mathsf{z}+1\otimes\mathsf{b})>\delta/2>0$), define
$f(\mathsf{b})=(\tau\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z}-1\otimes\mathsf{b})^{-1}],\qquad\omega(\mathsf{b})=\mathsf{x}^{-1}\mathsf{b}\mathsf{x}^{-1}+f(\mathsf{b}).$
Then Lemma 4.3(b) applied at $\mathsf{z}+1\otimes\mathsf{b}$ shows that $f(\mathsf{b})$ is the unique solution in $\mathcal{X}^{+}$ to the fixed-point equation $f(\mathsf{b})=n^{-1}\sum_{i}({-}\mathsf{z}_{i}-\mathsf{b}-\mathsf{x}f(\mathsf{b})\mathsf{x})^{-1}$, i.e.
$\mathsf{x}^{-1}\mathsf{b}\mathsf{x}^{-1}=\omega(\mathsf{b})-\frac{1}{n}\sum_{i=1}^{n}({-}\mathsf{z}_{i}-\mathsf{x}\,\omega(\mathsf{b})\mathsf{x})^{-1}.$ (55)
For all $\omega\in\mathcal{X}^{+}$, define
$g(\omega)=\omega-\frac{1}{n}\sum_{i=1}^{n}({-}\mathsf{z}_{i}-\mathsf{x}\,\omega\mathsf{x})^{-1}.$ (56)
Then for all $\mathsf{b}\in\mathcal{X}$ with $\|\mathsf{b}\|_{\mathrm{op}}<\delta/2$, since $\omega(\mathsf{b})$ satisfies (55), we have $g(\omega(\mathsf{b}))=\mathsf{x}^{-1}\mathsf{b}\mathsf{x}^{-1}$.
Let us write $D\omega(\mathsf{b}),Dg(\omega)$ for the Fréchet derivatives of $\omega(\mathsf{b})$ and $g(\omega)$ as linear maps on $\mathcal{X}$. Recalling $\mathsf{r}=(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}$ and differentiating $\omega(\mathsf{b})=\mathsf{x}^{-1}\mathsf{b}\mathsf{x}^{-1}+(\tau\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z}-1\otimes\mathsf{b})^{-1}]$ at $\mathsf{b}=0$, for any $\mathsf{a}\in\mathcal{X}$ and $p\in[1,\infty]$ we have
$\|D\omega(0)[\mathsf{a}]\|_{p}$ $=\|\mathsf{x}^{-1}\mathsf{a}\mathsf{x}^{-1}+(\tau\otimes 1)[\mathsf{r}(1\otimes\mathsf{a})\mathsf{r}]\|_{p}$
$\leq\gamma^{2}\|\mathsf{a}\|_{p}+\|\mathsf{r}\|_{\mathrm{op}}^{2}\|1\otimes\mathsf{a}\|_{p}\leq(\gamma^{2}+\delta^{-2})\|\mathsf{a}\|_{p}.$ (57)
Here, the second line uses Hölder’s inequality (Lemma D.3), $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$, contractivity of the conditional expectation $\tau\otimes 1$ in $L^{p}$ (Lemma D.4), and $\|\mathsf{r}\|_{\mathrm{op}}\leq\delta^{-1}$ by the assumption $\Im\mathsf{z}\geq\delta$ and Lemma D.1. On the other hand, differentiating (56) at $\omega_{0}=\omega(0)$ and using $({-}\mathsf{z}_{i}-\mathsf{x}\omega_{0}\mathsf{x})^{-1}=({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}=(\mathsf{r}_{0})_{ii}$ by Lemma 4.3(a), we have
$Dg(\omega_{0})[\mathsf{a}]=\mathsf{a}-\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}=\mathcal{L}_{1}(\mathsf{a}).$
Then differentiating both sides of the identity $g(\omega(\mathsf{b}))=\mathsf{x}^{-1}\mathsf{b}\mathsf{x}^{-1}$ at $\mathsf{b}=0$ and evaluating the derivative at $\mathsf{x}\mathsf{a}\mathsf{x}$ for any $\mathsf{a}\in\mathcal{X}$ gives
$\mathcal{L}_{1}(D\omega(0)[\mathsf{x}\mathsf{a}\mathsf{x}])=Dg(\omega_{0})[D\omega(0)[\mathsf{x}\mathsf{a}\mathsf{x}]]=\mathsf{x}^{-1}[\mathsf{x}\mathsf{a}\mathsf{x}]\mathsf{x}^{-1}=\mathsf{a}.$
The bound (57) for $p=\infty$ shows that $\mathsf{a}\mapsto D\omega(0)[\mathsf{x}\mathsf{a}\mathsf{x}]$ defines a bounded linear operator on $\mathcal{X}$, so $\mathcal{L}_{1}$ is invertible with inverse given explicitly by $\mathcal{L}_{1}^{-1}(\mathsf{a})=D\omega(0)[\mathsf{x}\mathsf{a}\mathsf{x}]$. Finally, (57) and the condition $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$ imply that $\|\mathcal{L}_{1}^{-1}(\mathsf{a})\|_{p}\leq\gamma^{2}(\gamma^{2}+\delta^{-2})\|\mathsf{a}\|_{p}$. ∎
We now turn to the invertibility of $\mathcal{L}_{2}$. We use an idea from [2, Section 4.2], which relies on the observation that $\mathcal{L}_{2}^{-1}$ may be controlled at a single positive element of $\mathcal{X}$ by taking imaginary parts of (49), and that the linear map $\mathsf{a}\mapsto(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}$ is positivity-preserving. For finite-dimensional matrix algebras, this implies that this map has a Perron-Frobenius eigenvector $\mathsf{a}\geq 0$ in the positive cone. The analyses of [2] construct a symmetrized version of this map that is self-adjoint, so that its $L^{2}\to L^{2}$ operator norm coincides with its spectral radius, and then apply a $L^{2}$-inner-product of the Perron-Frobenius eigenvector with the imaginary part of (49) to deduce a quantitative bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{2}\to L^{2}}$.
We adapt this idea to address two additional challenges in our setting: First, as the map $\mathsf{a}\mapsto(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}$ is non-compact and may have continuous spectrum, we can only guarantee the existence of an approximate Perron-Frobenius eigenvector (c.f. Lemma 4.6 below). For reasons that will be clear in the proof, the approximation error must be controlled in $L^{1}$ rather than $L^{2}$, and thus we implement a version of this argument in the $L^{1}$-$L^{\infty}$ duality rather than in a Hilbert space setting, to obtain a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{\infty}\to L^{\infty}}$. Second, as we will also require a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ for each $p\in[1,\infty)$, we carry out a dual version of this argument to obtain also a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{1}\to L^{1}}$, and hence deduce a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ via the Riesz-Thorin interpolation.
Lemma 4.6 (Approximate Perron-Frobenius eigenvector).
Let $(\mathcal{X},\|\cdot\|)$ be a real Banach space, and $\mathcal{K}\subset\mathcal{X}$ a closed convex cone such that $\mathcal{K}\cap(-\mathcal{K})=\{0\}$ ($\mathcal{K}$ is proper), $\mathcal{X}=\{x-y:x,y\in\mathcal{K}\}$ ($\mathcal{K}$ is generating), and for some $C>0$ we have $\|x\|\leq C\|x+y\|$ whenever $x,y\in\mathcal{K}$ ($\mathcal{K}$ is normal).
Let $T:\mathcal{X}\to\mathcal{X}$ be a bounded linear operator such that $T(\mathcal{K})\subseteq\mathcal{K}$, and let $r(T)$ be the spectral radius of its complexification $T_{\mathbb{C}}:\mathcal{X}+i\mathcal{X}\to\mathcal{X}+i\mathcal{X}$. Then $r(T)$ is an element of the spectrum of $T$. Furthermore, for any $\epsilon>0$, there exists an approximate eigenvector $x$ of $T$ such that
$x\in\mathcal{K},\qquad\|x\|=1,\qquad\|T(x)-r(T)x\|<\epsilon.$
Proof.
See [33, Lemma 3.5]. ∎
Lemma 4.7.
Under Assumption 4.1, consider the linear operator $\mathcal{L}_{2}:\mathcal{X}\to\mathcal{X}$ given by
$\mathcal{L}_{2}(\mathsf{a})=\mathsf{a}-\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}.$
Then $\mathcal{L}_{2}$ is invertible, and for any $\mathsf{a}\in\mathcal{X}$ and $p\in[1,\infty]$ (where $\|\cdot\|_{\infty}=\|\cdot\|_{\mathrm{op}}$),
$\|\mathcal{L}_{2}^{-1}(\mathsf{a})\|_{p}\leq\frac{2\gamma^{4}(\upsilon+\gamma^{2}\delta^{-1})^{2}}{\delta^{2}}\|\mathsf{a}\|_{p}.$ (58)
Proof.
Denote $L^{p}\equiv L^{p}(\mathcal{X})$ for the non-commutative $L^{p}$-spaces associated to $\mathcal{X}$ (Appendix D) and define $\mathcal{F}:L^{1}\to L^{1}$ and $\mathcal{F}^{\prime}:L^{\infty}\to L^{\infty}$ by
$\mathcal{F}(\mathsf{b})=\frac{1}{n}\sum_{i=1}^{n}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}\mathsf{b}(\mathsf{r}_{0})_{ii}\mathsf{x},\qquad\mathcal{F}^{\prime}(\mathsf{a})=\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}.$
Here $\mathcal{F}$ and $\mathcal{F}^{\prime}$ are bounded linear operators on the Banach spaces $(L^{1},\|\cdot\|_{1})$ and $(L^{\infty},\|\cdot\|_{\mathrm{op}})$, by Hölder’s inequality and the bounds $\|(\mathsf{r}_{0})_{ii}\|_{\mathrm{op}}\leq\|\mathsf{r}_{0}\|_{\mathrm{op}}\leq\delta^{-1}$ and $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$. Identifying the dual $(L^{1})^{*}$ with $L^{\infty}$ via the isometry $\mathsf{a}\in L^{\infty}\mapsto\ell_{\mathsf{a}}\in(L^{1})^{*}$ where $\ell_{\mathsf{a}}(\mathsf{b})=\phi(\mathsf{b}\mathsf{a})$ (Lemma D.3), for any $\mathsf{a}\in L^{\infty}$ and $\mathsf{b}\in L^{1}$ we have $\ell_{\mathsf{a}}(\mathcal{F}(\mathsf{b}))=\phi(\mathcal{F}(\mathsf{b})\mathsf{a})=\phi(b\mathcal{F}^{\prime}(\mathsf{a}))=\ell_{\mathcal{F}^{\prime}(\mathsf{a})}(\mathsf{b})$, so $\mathcal{F}^{\prime}$ is the Banach space adjoint of $\mathcal{F}$.
Let $r(\mathcal{F})$ be the spectral radius of $\mathcal{F}$ as an operator on $L^{1}$. Note that $\mathcal{F}$ restricts to a bounded linear operator on the real Banach space of self-adjoint elements $L^{1}_{\text{sa}}=\{\mathsf{a}\in L^{1}:\mathsf{a}=\mathsf{a}^{*}\}$ (whose complexification is $\mathcal{F}$ itself), which furthermore preserves the positive cone $L^{1}_{\text{pos}}=\{\mathsf{a}\in L^{1}_{\text{sa}}:\mathsf{a}\geq 0\}$. Any $\mathsf{a}\in L^{1}_{\text{sa}}$ may be decomposed as $\mathsf{a}=\mathsf{a}_{+}-\mathsf{a}_{-}$ with $\mathsf{a}_{+},\mathsf{a}_{-}\in L^{1}_{\text{pos}}$ via $\mathsf{a}_{+}=(|\mathsf{a}|+\mathsf{a})/2$ and $\mathsf{a}_{-}=(|\mathsf{a}|-\mathsf{a})/2$, and we have $\|\mathsf{a}\|_{1}=\tau(\mathsf{a})\leq\tau(\mathsf{a}+\mathsf{b})=\|\mathsf{a}+\mathsf{b}\|_{1}$ for all $\mathsf{a},\mathsf{b}\in L^{1}_{\text{pos}}$ by positivity of the trace. Thus $L^{1}_{\text{pos}}$ is a proper, generating, and normal cone in $L^{1}_{\text{sa}}$, so Lemma 4.6 ensures the existence of some $\mathsf{v}\in L^{1}$ satisfying
$\mathsf{v}=\mathsf{v}^{*},\qquad\mathsf{v}\geq 0,\qquad\|\mathsf{v}\|_{1}=1,\qquad\mathcal{F}(\mathsf{v})=r(\mathcal{F})\mathsf{v}+\Delta\text{ where }\|\Delta\|_{1}<\epsilon.$ (59)
Lemma 4.3 shows that $\mathsf{m}_{0}$ satisfies the fixed-point equation $\mathsf{m}_{0}=n^{-1}\sum_{i=1}^{n}({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}=n^{-1}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}$. Taking imaginary parts, this gives $\Im\mathsf{m}_{0}=n^{-1}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}(\Im\mathsf{z}_{i}+\mathsf{x}[\Im\mathsf{m}_{0}]\mathsf{x})(\mathsf{r}_{0})_{ii}^{*}$, which may be rearranged as
$\mathcal{L}_{2}[\Im\mathsf{m}_{0}]=(\mathrm{Id}-\mathcal{F}^{\prime})[\Im\mathsf{m}_{0}]=\mathsf{u},\qquad\text{ for }\mathsf{u}=\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}[\Im\mathsf{z}_{i}](\mathsf{r}_{0})_{ii}^{*}.$ (60)
Multiplying on the left by the above approximate eigenvector $\mathsf{v}$ and taking the trace,
$\phi(\mathsf{v}\mathsf{u})=\phi(\mathsf{v}\Im\mathsf{m}_{0})-\phi(\mathsf{v}\mathcal{F}^{\prime}[\Im\mathsf{m}_{0}])=\phi(\mathsf{v}\Im\mathsf{m}_{0})-\phi(\mathcal{F}[\mathsf{v}]\Im\mathsf{m}_{0})=(1-r(\mathcal{F}))\phi(\mathsf{v}\Im\mathsf{m}_{0})-\phi(\Delta\Im\mathsf{m}_{0}).$
Thus
$1-r(\mathcal{F})=\frac{\phi(\mathsf{v}\mathsf{u})+\phi(\Delta\Im\mathsf{m}_{0})}{\phi(\mathsf{v}\Im\mathsf{m}_{0})}.$
By Hölder’s inequality and the bound $\|\Im\mathsf{m}_{0}\|_{\mathrm{op}}\leq\|\mathsf{m}_{0}\|_{\mathrm{op}}\leq\delta^{-1}$, we have
$0\leq\phi(\mathsf{v}\Im\mathsf{m}_{0})\leq\delta^{-1},\qquad|\phi(\Delta\Im\mathsf{m}_{0})|\leq\delta^{-1}\epsilon.$ (61)
We have also $\Im\mathsf{z}_{i}\geq\delta$ and $\|(\mathsf{r}_{0})_{ii}^{-1}\|_{\mathrm{op}}=\|\mathsf{z}_{i}+\mathsf{x}\mathsf{m}_{0}\mathsf{x}\|_{\mathrm{op}}\leq\|\mathsf{z}\|_{\mathrm{op}}+\|\mathsf{x}\|_{\mathrm{op}}^{2}\|\mathsf{m}_{0}\|_{\mathrm{op}}\leq\upsilon+\gamma^{2}\delta^{-1}$, implying that
$\mathsf{u}\geq\delta\cdot\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}(\mathsf{r}_{0})_{ii}^{*}\geq c(\gamma,\upsilon,\delta)\qquad\text{ for }c(\gamma,\upsilon,\delta)=\delta(\upsilon+\gamma^{2}\delta^{-1})^{-2}.$ (62)
Then by positivity of $\mathsf{v}$, positivity of the trace, and the normalization $\|\mathsf{v}\|_{1}=1$, we have $\phi(\mathsf{v}\mathsf{u})=\phi(\mathsf{v}^{1/2}\mathsf{u}\mathsf{v}^{1/2})\geq c(\gamma,\upsilon,\delta)\phi(\mathsf{v})=c(\gamma,\upsilon,\delta)$. Applying these bounds above and then taking $\epsilon\to 0$, we obtain
$01-r(\mathcal{F})\geq\delta c(\gamma,\upsilon,\delta)>$ 0 . (63)
Since $\mathcal{F}^{\prime}$ is the adjoint of $\mathcal{F}$ and thus shares its spectrum, this shows $r(\mathcal{F}^{\prime})=r(\mathcal{F})<1$. Then $\mathcal{L}_{2}=\mathrm{Id}-\mathcal{F}^{\prime}$ is invertible as a bounded linear operator on $L^{\infty}$, and its inverse has the Neumann series representation
$\mathcal{L}_{2}^{-1}=\sum_{k=0}^{\infty}(\mathcal{F}^{\prime})^{k}$
which is convergent in the induced operator norm $\|\cdot\|_{L^{\infty}\to L^{\infty}}$ (by Gelfand’s formula $r(\mathcal{F}^{\prime})=\lim_{k\to\infty}\|{\mathcal{F}^{\prime}}^{k}\|_{L^{\infty}\to L^{\infty}}^{1/k}$). Thus, for any $\mathsf{a}\in L^{\infty}$, we have $\mathcal{L}_{2}^{-1}(\mathsf{a})=\sum_{k=0}^{\infty}(\mathcal{F}^{\prime})^{k}[\mathsf{a}]$ which is convergent under $\|\cdot\|_{\mathrm{op}}$. We note that $(\mathcal{F}^{\prime})^{k}$ is also positivity-preserving, i.e. if $\mathsf{a}\geq 0$, then $(\mathcal{F}^{\prime})^{k}[\mathsf{a}]\geq 0$. Thus if $\mathsf{a}\geq 0$, then $\mathcal{L}_{2}^{-1}(\mathsf{a})\geq 0$ since the positive cone is closed under $\|\cdot\|_{\mathrm{op}}$. To summarize, we have shown that $\mathcal{L}_{2}:L^{\infty}\to L^{\infty}$ is invertible, and
$\mathsf{a}\geq 0\Longrightarrow\mathcal{L}_{2}^{-1}(\mathsf{a})\geq 0.$ (64)
Now take any $\mathsf{b}\in L^{\infty}$ self-adjoint. Applying (62), we have
$\mathcal{L}_{2}(\mathsf{b})\leq\|\mathcal{L}_{2}(\mathsf{b})\|_{\mathrm{op}}\cdot 1_{\mathcal{X}}\leq\|\mathcal{L}_{2}(\mathsf{b})\|_{\mathrm{op}}\cdot c(\gamma,\upsilon,\delta)^{-1}\cdot\mathsf{u}$
so the monotonicity of $\mathcal{L}_{2}^{-1}$ in (64) and explicit form of $\mathcal{L}_{2}^{-1}[\mathsf{u}]$ in (60) imply
$\mathsf{b}\leq\|\mathcal{L}_{2}(\mathsf{b})\|_{\mathrm{op}}\cdot c(\gamma,\upsilon,\delta)^{-1}\mathcal{L}_{2}^{-1}[\mathsf{u}]=\|\mathcal{L}_{2}(\mathsf{b})\|_{\mathrm{op}}\cdot c(\gamma,\upsilon,\delta)^{-1}\Im\mathsf{m}_{0}.$
Similarly $\mathsf{b}\geq{-}\|\mathcal{L}_{2}(\mathsf{b})\|_{\mathrm{op}}\cdot c(K,\upsilon,\delta)^{-1}\Im\mathsf{m}_{0}$. Applying again $\|\Im\mathsf{m}_{0}\|_{\mathrm{op}}\leq\delta^{-1}$, this shows for every $\mathsf{b}\in L^{\infty}$ self-adjoint that $\|\mathsf{b}\|_{\mathrm{op}}\leq\delta^{-1}c(K,\upsilon,\delta)^{-1}\|\mathcal{L}_{2}(\mathsf{b})\|_{\mathrm{op}}$. Then for any (non-self-adjoint) $\mathsf{a}\in L^{\infty}$, noting that $(\|\Re\mathsf{a}\|_{\mathrm{op}}+\|\Im\mathsf{a}\|_{\mathrm{op}})/2\leq\|\mathsf{a}\|_{\mathrm{op}}\leq\|\Re\mathsf{a}\|_{\mathrm{op}}+\|\Im\mathsf{a}\|_{\mathrm{op}}$ and that $\Re\mathcal{L}_{2}(\mathsf{a})=\mathcal{L}_{2}(\Re\mathsf{a})$ and $\Im\mathcal{L}_{2}(\mathsf{a})=\mathcal{L}_{2}(\Im\mathsf{a})$, this implies
$\|\mathsf{a}\|_{\mathrm{op}}\leq 2\delta^{-1}c(\gamma,\delta,\upsilon)^{-1}\|\mathcal{L}_{2}(\mathsf{a})\|_{\mathrm{op}}=\frac{2(\upsilon+\gamma^{2}\delta^{-1})^{2}}{\delta^{2}}\|\mathcal{L}_{2}(\mathsf{a})\|_{\mathrm{op}},$ (65)
which implies (58) for $p=\infty$.
Next, we show (58) for $p=1$ using a dual argument: As $L^{1}$ is not reflexive, we reverse the roles of $\mathcal{F},\mathcal{F}^{\prime}$ and define the bounded linear operators $\mathcal{G}:L^{1}\to L^{1}$ and $\mathcal{G}^{\prime}:L^{\infty}\to L^{\infty}$ by
$\mathcal{G}(\mathsf{b})=\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{b}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*},\qquad\mathcal{G}^{\prime}(\mathsf{a})=\frac{1}{n}\sum_{i=1}^{n}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}\mathsf{a}(\mathsf{r}_{0})_{ii}\mathsf{x}.$
Then again $\mathcal{G}^{\prime}$ is the adjoint of $\mathcal{G}$. For any $\epsilon>0$, Lemma 4.6 shows there exists $\mathsf{v}\in L^{1}$ satisfying
$\mathsf{v}=\mathsf{v}^{*},\qquad\mathsf{v}\geq 0,\qquad\|\mathsf{v}\|_{1}=1,\qquad\mathcal{G}(\mathsf{v})=r(\mathcal{G})\mathsf{v}+\Delta\text{ where }\|\Delta\|_{1}<\epsilon.$
Now taking imaginary parts of the fixed-point equation $\mathsf{m}_{0}^{*}=n^{-1}\sum_{i=1}^{n}({-}\mathsf{z}_{i}^{*}-\mathsf{x}\mathsf{m}_{0}^{*}\mathsf{x})^{-1}$, we have $\Im(\mathsf{m}_{0}^{*})=n^{-1}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}^{*}(\Im(\mathsf{z}_{i}^{*})+\mathsf{x}[\Im(\mathsf{m}_{0}^{*})]\mathsf{x})(\mathsf{r}_{0})_{ii}$. Negating and applying $\Im(\mathsf{a}^{*})={-}\Im\mathsf{a}$, this gives
$\Im\mathsf{m}_{0}=\mathsf{x}^{-1}\mathcal{G}^{\prime}(\mathsf{x}[\Im\mathsf{m}_{0}]\mathsf{x})\mathsf{x}^{-1}+\frac{1}{n}\sum_{i=1}^{n}(\mathsf{r}_{0})_{ii}^{*}[\Im\mathsf{z}_{i}](\mathsf{r}_{0})_{ii}.$
Thus, in place of (60) we have
$(\mathrm{Id}-\mathcal{G}^{\prime})[\mathsf{x}(\Im\mathsf{m}_{0})\mathsf{x}]=\mathsf{w},\qquad\text{ for }\mathsf{w}=\frac{1}{n}\sum_{i=1}^{n}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}[\Im\mathsf{z}_{i}](\mathsf{r}_{0})_{ii}\mathsf{x}.$ (66)
In place of (61) and (62), we may apply $\phi(\mathsf{v}\mathsf{x}[\Im\mathsf{m}_{0}]\mathsf{x})\leq\gamma^{2}\delta^{-1}$ and $\mathsf{w}\geq\gamma^{-2}c(\gamma,\upsilon,\delta)$, where the first inequality uses $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$ and the second uses $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$ and $c(\gamma,\upsilon,\delta)$ as defined in (62). Then, multiplying (66) by $\mathsf{v}$, taking the trace, and then taking the limit $\epsilon\to 0$, we obtain similarly to (63) that $01-r(\mathcal{G})\geq\delta\gamma^{-4}c(\gamma,\upsilon,\delta)>$ 0. This implies that $\mathrm{Id}-\mathcal{G}^{\prime}:L^{\infty}\to L^{\infty}$ is invertible with positivity-preserving inverse, and repeating the preceding arguments gives, for any $\mathsf{a}\in L^{\infty}$,
$\|\mathsf{a}\|_{\mathrm{op}}\leq\frac{2\gamma^{4}(\upsilon+\gamma^{2}\delta^{-1})^{2}}{\delta^{2}}\|(\mathrm{Id}-\mathcal{G}^{\prime})[\mathsf{a}]\|_{\mathrm{op}},$
Since $(\mathrm{Id}-\mathcal{G}^{\prime})^{-1}$ is the adjoint of $\mathcal{L}_{2}^{-1}=(\mathrm{Id}-\mathcal{G})^{-1}:L^{1}\to L^{1}$, we have $\|(\mathrm{Id}-\mathcal{G}^{\prime})^{-1}\|_{L^{\infty}\to L^{\infty}}=\|\mathcal{L}_{2}^{-1}\|_{L^{1}\to L^{1}}$. Thus, this shows also for any $\mathsf{a}\in L^{1}$ that
$\|\mathsf{a}\|_{L^{1}}\leq\frac{2\gamma^{4}(\upsilon+\gamma^{2}\delta^{-1})^{2}}{\delta^{2}}\|\mathcal{L}_{2}(\mathsf{a})\|_{L^{1}},$ (67)
which is the desired result (58) for $p=1$.
Finally, the result (58) for general $p\in[1,\infty]$ follows from the bounds for $\|\mathcal{L}_{2}^{-1}\|_{L^{\infty}\to L^{\infty}}$ and $\|\mathcal{L}_{2}^{-1}\|_{L^{1}\to L^{1}}$ already shown, and the Riesz-Thorin interpolation (Lemma D.5). ∎
4.2. Weak estimates in operator norm
Lemma 4.8.
Under Assumption 4.1, there exists a constant $\alpha\in(0,1/2)$ depending only on $\gamma,\upsilon,\delta$ such that for any $l\geq 1$, $D>0$, and all $n\geq n_{0}(l,K,\gamma,\upsilon,\delta,D)$, with probability at least $1-n^{-D}$,
$\sup_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\sup_{i\notin S}\left\|\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}+\sum_{r\neq s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}\right\|_{\mathrm{op}}<n^{-\alpha},$ (68)
$\sup_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\sup_{i,j\notin S:i\neq j}\left\|\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\right\|_{\mathrm{op}}<n^{-\alpha}.$ (69)
Proof.
We present the argument for (68): Take any $S\subset\{1,\ldots,n\}$ with $|S|\leq l$, and any $i\notin S$. Let $H^{(iS)}$ be as defined in Section 3.4, and let $\mathbf{h}_{i}^{(iS)}\in\mathbb{C}^{n}$ be the $i^{\text{th}}$ column of $H^{(iS)}$, i.e. the vector with entries $(\mathbf{h}_{i}^{(iS)})_{j}=h_{ji}$ for $j\notin S\cup\{i\}$ and 0 otherwise. Let $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})$ be the Hilbert space on which $\mathcal{X}$ acts. For any unit vectors $\xi,\zeta\in\mathcal{H}$, define the linear functional $f_{i,\xi,\zeta}^{(S)}:\mathbb{C}^{n\times n}\otimes\mathcal{X}\to\mathbb{C}$ by
$f_{i,\xi,\zeta}^{(S)}(M)$ $=\left\langle\xi,\,\left(\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)M_{rr}+\sum_{r\neq s}^{(iS)}h_{ir}M_{rs}h_{si}\right)\zeta\right\rangle_{\mathcal{H}}$
$=\left\langle\xi,\,\left((\mathbf{h}_{i}^{(iS)}\otimes 1)^{*}M(\mathbf{h}_{i}^{(iS)}\otimes 1)-(n^{-1}\operatorname{Tr}^{(iS)}\otimes 1)M\right)\zeta\right\rangle_{\mathcal{H}}\quad\text{ for }\operatorname{Tr}^{(iS)}M:=\sum_{j}^{(iS)}M_{jj}$
where $M_{rs}=(\mathbf{e}_{r}\otimes 1)^{*}M(\mathbf{e}_{s}\otimes 1)$ is the $\mathcal{X}$-valued $(r,s)$ entry of $M$. Define $M^{(iS)}:\mathbb{C}^{+}\to\mathbb{C}^{n\times n}\otimes\mathcal{X}$ by
$M^{(iS)}(z)=\left(H^{(iS)}\otimes\mathsf{x}-\mathsf{z}+i(\delta/2)-z\right)^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}.$
Note that $\Im(\mathsf{z}-i(\delta/2)+z)>\delta/2$ for all $z\in\mathbb{C}^{+}$, so by Lemma D.1, this inverse is well-defined and
$\|M^{(iS)}(z)\|_{\mathrm{op}}\leq 2/\delta\text{ for all }z\in\mathbb{C}^{+}.$ (70)
By definition we have $R^{(iS)}=M^{(iS)}(i\delta/2)$, so the operator norm to be bounded in (68) is
$\left\|\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}+\sum_{r\neq s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}\right\|_{\mathrm{op}}=\sup_{\xi,\zeta\in\mathcal{H}:\|\xi\|_{\mathcal{H}}^{2}=\|\zeta\|_{\mathcal{H}}^{2}=1}\,f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(i\delta/2).$
Set $C_{0}=3\gamma+(3/2)\upsilon$, $\mathcal{D}=\{z\in\mathbb{C}^{+}:|z|\geq 2C_{0}\}$, and $\epsilon=0.1$, and define the event
$\mathcal{E}=\bigcap_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\bigcap_{i\notin S}\;\bigcap_{\xi,\zeta:\|\xi\|_{\mathcal{H}}^{2}=\|\zeta\|_{\mathcal{H}}^{2}=1}\left\{\sup_{z\in\mathcal{D}}|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|\leq n^{-1/2+\epsilon}\right\}\cap\{\|\mathbf{h}_{i}^{(iS)}\|_{2}\leq 3\}.$
Noting that $f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)$ is an analytic function of $z\in\mathbb{C}^{+}$, we apply Lemma 3.7 on this event $\mathcal{E}$: Let $a=3C_{0}$, and set $r_{0}<r_{1}<r_{2}<0$ such that
$a\frac{1-e^{r_{0}}}{1+e^{r_{0}}}=2C_{0},\qquad a\frac{1-e^{r_{1}}}{1+e^{r_{1}}}=\delta/2,\qquad a\frac{1-e^{r_{2}}}{1+e^{r_{2}}}=\delta/4.$
Then defining $S_{r}$ as in Lemma 3.7, we have $S_{r_{0}}\subset\mathcal{D}$ and $i\delta/2\in S_{r_{1}}$. On $\mathcal{E}$, the inclusion $S_{r_{0}}\subset\mathcal{D}$ implies $|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|\leq n^{-1/2+\epsilon}$ for all $z\in S_{r_{0}}$, and the bounds (70) and $\|\mathbf{h}_{i}^{(iS)}\|_{2}\leq 3$ imply $|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|\leq 20/\delta$ for all $z\in\mathbb{C}^{+}$. Thus Lemma 3.7 shows
$\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(i\delta/2)|$ $\leq\sup_{z\in S_{r_{1}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|$
$\leq\frac{|r_{1}-r_{0}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{0}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|+\frac{|r_{1}-r_{2}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{2}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|$
$\leq-\alpha\log n$
for some constant $\alpha\in(0,1/2)$ depending only on $\gamma,\upsilon,\delta$, and for all $n\geq n_{0}(\gamma,\upsilon,\delta)$. Thus (68) holds on the event $\mathcal{E}$ for all such $n$.
We now check that $\mathbb{P}[\mathcal{E}]\geq 1-n^{-D}$ for all $n\geq n_{0}(l,K,D)$: By a standard tail bound for the operator norm (see e.g. [23, Theorem 7.3]), the event
$\mathcal{E}_{0}$ $=\bigcap_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\bigcap_{i\notin S}\left\{\|H^{(iS)}\|_{\mathrm{op}}\leq 3\right\}\subset\{\|H\|_{\mathrm{op}}\leq 3\}$
holds with probability at least $1-n^{-D}$ for all $n\geq n_{0}(D)$. On $\mathcal{E}_{0}$, we have also $\|\mathbf{h}_{i}^{(iS)}\|_{2}\leq 3$. Set for notational convenience $D_{0}={-}H^{(iS)}$ (which depends implicitly on $i$ and $S$), $\mathsf{x}_{0}=\mathsf{x}$, $D_{K+1}={-}i\delta/2$, and $\mathsf{x}_{K+1}=1$ , so that
$H^{(iS)}\otimes\mathsf{x}-\mathsf{z}+i(\delta/2)={-}\sum_{k=0}^{K+1}D_{k}\otimes\mathsf{x}_{k}.$
By the assumptions (43), we have on $\mathcal{E}_{0}$ that
$\left\|H^{(iS)}\otimes\mathsf{x}-\mathsf{z}+i(\delta/2)\right\|_{\mathrm{op}}\leq\sum_{k=0}^{K+1}\|D_{k}\|_{\mathrm{op}}\|\mathsf{x}_{k}\|_{\mathrm{op}}\leq 3\gamma+(3/2)\upsilon=C_{0}.$
Then, since $|z|\geq 2C_{0}$ for $z\in\mathcal{D}$, the series expansion of $M^{(iS)}(z)$ in $z^{-1}$ is absolutely convergent in operator norm, and
$M^{(iS)}(z)={-}\sum_{t=0}^{T(n)}z^{-(t+1)}\left({-}\sum_{k=0}^{K+1}D_{k}\otimes\mathsf{x}_{k}\right)^{t}+J^{(iS)}(z),\qquad\|J^{(iS)}(z)\|_{\mathrm{op}}\leq n^{-1}$ (71)
for some $T(n)\leq C\log n$ and an absolute constant $C>0$. On $\mathcal{E}_{0}$, using again $\|\mathbf{h}_{i}^{(iS)}\|_{2}\leq 3$, we have $|f_{i,\xi,\zeta}^{(S)}(M)|\leq 10\|M\|_{\mathrm{op}}$ for any $M\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$. Thus, for the remainder of (71),
$|f_{i,\xi,\zeta}^{(S)}\circ J^{(iS)}(z)|\leq 10\|J^{(iS)}(z)\|_{\mathrm{op}}\leq 10n^{-1}.$ (72)
For the leading terms of (71), let us expand
$\left(\sum_{k=0}^{K+1}D_{k}\otimes\mathsf{x}_{k}\right)^{t}=\sum_{w\in\mathcal{W}_{t}}\underbrace{w(D_{0},\ldots,D_{K+1})}_{:=w(D^{(iS)})}\otimes\underbrace{w(\mathsf{x}_{0},\ldots,\mathsf{x}_{K+1})}_{:=w(\mathsf{x})}$ (73)
where $\mathcal{W}_{t}$ denotes the set of all length-$t$ words (i.e. non-commutative degree-$t$ monic monomials) in $K+2$ variables. For $t=0$, we use the convention $\mathcal{W}_{t}=\{o\}$ where $o$ is the word of length 0, with $o(D^{(iS)})=I$ and $o(\mathsf{x})=1$. Then for each $w\in\mathcal{W}_{t}$, by the definition of $f_{i,\xi,\zeta}^{(S)}$, we have the factorization
$f_{i,\xi,\zeta}^{(S)}\big{(}w(D^{(iS)})\otimes w(\mathsf{x})\big{)}$ $=\left(\mathbf{h}_{i}^{(iS)*}w(D^{(iS)})\mathbf{h}_{i}^{(iS)}-n^{-1}\operatorname{Tr}^{(iS)}w(D^{(iS)})\right)\cdot\langle\xi,w(\mathsf{x})\zeta\rangle_{\mathcal{H}}.$ (74)
Since $\mathbb{E}[|h_{ji}|^{2}]=n^{-1}$, the scalar version of Lemma 3.4(a,c) (i.e. with $\mathcal{X}=\mathbb{C}$) implies that uniformly over deterministic matrices $M\in\mathbb{C}^{n\times n}$,
$\mathbf{h}_{i}^{(iS)*}M\mathbf{h}_{i}^{(iS)}-n^{-1}\operatorname{Tr}^{(iS)}M\prec n^{-1}\left(\sum_{r,s}^{(iS)}|M_{rs}|^{2}\right)^{1/2}\prec n^{-1/2}\|M\|_{\mathrm{op}}.$ (75)
Note that $w(D^{(iS)})$ is independent of $\mathbf{h}_{i}^{(iS)}$, and |$|\{S\subset\{1,\ldots,n\}:|S|\leq l\}|\leq n^{l}$ and $|\mathcal{W}_{t}|=(K+2)^{t}\leq n^{C\log(K+2)}$ for all $t\leq T(n)\leq C\log n$. Then, fixing $\epsilon^{\prime}=0.05$, applying (75) to each matrix $M=w(D^{(iS)})$ conditional on $H^{(iS)}$, and taking a union bound, the event
$\mathcal{E}_{1}=\bigcap_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\bigcap_{i\notin S}\bigcap_{t=0}^{T(n)}\bigcap_{w\in\mathcal{W}_{t}}\left\{\left|\mathbf{h}_{i}^{(iS)*}w(D^{(iS)})\mathbf{h}_{i}^{(iS)}-n^{-1}\operatorname{Tr}^{(iS)}w(D^{(iS)})\right|\leq n^{-1/2+\epsilon^{\prime}}\|w(D^{(iS)})\|_{\mathrm{op}}\right\}$
holds with probability at least $1-n^{-D}$ for all $n\geq n_{0}(l,K,D)$. On $\mathcal{E}_{1}$, applying (74) to (73) gives, for each $t=0,\ldots,T(n)$,
$\left|f_{i,\xi,\zeta}^{(S)}\left(\left(\sum_{k=0}^{K+1}D_{k}\otimes\mathsf{x}_{k}\right)^{t}\right)\right|$ $\leq n^{-1/2+\epsilon^{\prime}}\sum_{w\in\mathcal{W}_{t}}\|w(D^{(iS)})\|_{\mathrm{op}}\|w(\mathsf{x})\|_{\mathrm{op}}$
$\leq n^{-1/2+\epsilon^{\prime}}\sum_{w\in\mathcal{W}_{t}}w(\|D_{0}\|_{\mathrm{op}}\|\mathsf{x}_{0}\|_{\mathrm{op}},\ldots,\|D_{K+1}\|_{\mathrm{op}}\|\mathsf{x}_{K+1}\|_{\mathrm{op}})$
$=n^{-1/2+\epsilon^{\prime}}\Big{(}\|D_{0}\|_{\mathrm{op}}\|\mathsf{x}_{0}\|_{\mathrm{op}}+\ldots+\|D_{K+1}\|_{\mathrm{op}}\|\mathsf{x}_{K+1}\|_{\mathrm{op}}\Big{)}^{t}$
$\leq n^{-1/2+\epsilon^{\prime}}C_{0}^{t}.$ (76)
Then, applying (76) and (72) to (71), we have on $\mathcal{E}_{0}\cap\mathcal{E}_{1}$ for all $z\in\mathcal{D}$ that
$|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|\leq\sum_{t=0}^{T(n)}|z|^{-(t+1)}n^{-1/2+\epsilon^{\prime}}C_{0}^{t}+10n^{-1}\leq n^{-1/2+\epsilon},$
the final inequality holding for our preceding choices of $\epsilon=0.1$, $\epsilon^{\prime}=0.05$, and all $n\geq n_{0}$ since $|z|\geq 2C_{0}$. So $\mathcal{E}_{0}\cap\mathcal{E}_{1}\subseteq\mathcal{E}$, implying that $\mathbb{P}[\mathcal{E}]\geq 1-n^{-D}$ for $n\geq n_{0}(l,K,D)$, as claimed.
This shows that (68) holds with probability $1-n^{-D}$. The proof for (69) is the same, applying these arguments with the function
$f_{i,j,\xi,\zeta}^{(S)}(M)=\left\langle\xi,\left(\sum_{r}^{(iS)}h_{ir}M_{rj}\right)\zeta\right\rangle=\left\langle\xi,\left((\mathbf{h}_{i}^{(iS)}\otimes 1)^{*}M(\mathbf{e}_{j}\otimes 1)\right)\zeta\right\rangle$
in place of $f_{i,\xi,\zeta}^{(S)}$. ∎
Lemma 4.9.
Under Assumption 4.1, there exists a constant $\alpha\in(0,1/2)$ depending only on $\gamma,\upsilon,\delta$ such that for any $l\geq 1$, $D>0$, and all $n\geq n_{0}(l,K,\gamma,\upsilon,\delta,D)$, with probability at least $1-n^{-D}$,
$\sup_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\sup_{i\notin S}\|R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\|_{\mathrm{op}}<n^{-\alpha},\qquad\sup_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\sup_{i,j\notin S:i\neq j}\|R_{ij}^{(S)}\|_{\mathrm{op}}<n^{-\alpha}.$ (77)
Proof.
Let $\alpha\in(0,1/2)$ be as in Lemma 4.8. Fixing $l\geq 1$, let $\mathcal{E}$ be the event on which the statements (6869) of Lemma 4.8 hold, and in addition, $\sup_{i=1}^{n}|h_{ii}|<n^{-\alpha}$. Lemma 4.8 and Assumption 2.1 for $H$ imply that $\mathcal{E}$ holds with probability at least $1-n^{-D}$ for all $n\geq n_{0}(l,K,\gamma,\upsilon,\delta,D)$.
Take any $S\subset\{1,\ldots,n\}$ with $|S|\leq l$, and any $i,j\notin S$ with $i\neq j$. By Lemma 3.6(b), on $\mathcal{E}$,
$\|R_{ij}^{(S)}\|_{\mathrm{op}}\leq\|R_{ii}^{(S)}\|_{\mathrm{op}}\|\mathsf{x}\|_{\mathrm{op}}\left\|\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\right\|_{\mathrm{op}}\leq\gamma\delta^{-1}n^{-\alpha}$ (78)
where we have used also $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma D.1 and the assumption $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$. Adjusting the value of $\alpha$ yields the second statement of (77).
For the first statement of (77), take any $S\subset\{1,\ldots,n\}$ with $|S|\leq l$. Define $\widetilde{R}_{ii}^{(S)}=R_{ii}^{(S)}$ for $i\notin S$, and
$\widetilde{R}_{ii}^{(S)}=\bigg{(}{-}\mathsf{z}_{i}-\mathsf{x}\bigg{[}\frac{1}{n}\sum_{j}^{(S)}R_{jj}^{(S)}\bigg{]}\mathsf{x}\bigg{)}^{-1}\text{ for }i\in S.$ (79)
By this definition and Lemma 3.6(a), for each $i=1,\ldots,n$ we have
$\widetilde{R}_{ii}^{(S)}=\bigg{(}{-}\mathsf{z}_{i}-\Delta_{i}^{(S)}-\mathsf{x}\bigg{[}\frac{1}{n}\sum_{j=1}^{n}\widetilde{R}_{jj}^{(S)}\bigg{]}\mathsf{x}\bigg{)}^{-1}$ (80)
where
$\Delta_{i}^{(S)}$ $={-}h_{ii}\mathsf{x}+\mathsf{x}\left(\sum_{r,s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}-\frac{1}{n}\sum_{j=1}^{n}\widetilde{R}_{jj}^{(S)}\right)\mathsf{x}$
$={-}h_{ii}\mathsf{x}+\mathsf{x}\Bigg{[}\underbrace{\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}+\sum_{r\neq s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}}_{:=\mathrm{I}}-\underbrace{\frac{1}{n}\sum_{j\in S\cup\{i\}}\widetilde{R}_{jj}^{(S)}}_{:=\mathrm{II}}+\underbrace{\frac{1}{n}\sum_{j}^{(iS)}(R_{jj}^{(iS)}-R_{jj}^{(S)})}_{:=\mathrm{III}}\Bigg{]}\mathsf{x}$
[…]
and
$\Delta_{i}^{(S)}={-}\mathsf{x}\bigg{[}\underbrace{\frac{1}{n}\sum_{j\in S}\widetilde{R}_{jj}^{(S)}}_{:=\mathrm{IV}}\bigg{]}\mathsf{x}\qquad\text{ for all }i\in S.$ (82)
On $\mathcal{E}$, we have $\|\mathrm{I}\|_{\mathrm{op}}\leq n^{-\alpha}$ by (68). For $\mathrm{II}$ and $\mathrm{IV}$, Lemma D.1 implies $\|\widetilde{R}_{ii}^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ for both $i\in S$ and $i\notin S$, the latter because $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ and the former because $\Im R^{(S)}\geq 0$ so $\Im R_{jj}^{(S)}\geq 0$ and $\Im\mathsf{z}_{i}\geq\delta$ in the definition (79). Then
$\|\mathrm{II}\|_{\mathrm{op}},\|\mathrm{IV}\|_{\mathrm{op}}\leq\frac{(|S|+1)\delta^{-1}}{n}.$ (83)
For $\mathrm{III}$, we have by Lemma 3.6(b–c) that
$\mathrm{III}={-}\frac{1}{n}\sum_{j}^{(iS)}R_{ji}^{(S)}\frac{1}{R_{ii}^{(S)}}R_{ij}^{(S)}=\frac{1}{n}\sum_{j}^{(iS)}R_{ji}^{(S)}\mathsf{x}\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}.$ (84)
Thus on $\mathcal{E}$, $\|\mathrm{III}\|_{\mathrm{op}}\leq\gamma^{2}\delta^{-1}n^{-2\alpha}$ by (69) and (78). Collecting these bounds and applying also $|h_{ii}|\leq n^{-\alpha}$ on $\mathcal{E}$, we have for a constant $C(\gamma,\delta)>0$ and every $i=1,\ldots,n$ that
$\|\Delta_{i}^{(S)}\|_{\mathrm{op}}\leq C(\gamma,\delta)n^{-\alpha}\text{ on }\mathcal{E}.$ (85)
Taking the Kronecker product with $E_{ii}$ on both sides of (80), summing over $i=1,\ldots,n$, recalling $\mathsf{z}=\sum_{i}E_{ii}\otimes\mathsf{z}_{i}$, and setting $\Delta^{(S)}=\sum_{i}E_{ii}\otimes\Delta_{i}^{(S)}$, we have
$\sum_{i=1}^{n}E_{ii}\otimes\widetilde{R}_{ii}^{(S)}=\left(-\mathsf{z}-\Delta^{(S)}-I_{n\times n}\otimes\mathsf{x}\left(\frac{1}{n}\operatorname{Tr}\otimes 1\left[\sum_{i=1}^{n}E_{ii}\otimes\widetilde{R}_{ii}^{(S)}\right]\right)\mathsf{x}\right)^{-1}.$ (86)
On $\mathcal{E}$, by the estimate $\|\Delta^{(S)}\|_{\mathrm{op}}=\max_{i}\|\Delta_{i}^{(S)}\|_{\mathrm{op}}\leq C(\gamma,\delta)n^{-\alpha}$ from (85), for all $n\geq n_{0}(\gamma,\upsilon,\delta)$ this implies $\Im(\mathsf{z}+\Delta^{(S)})\geq\delta/2$, so Corollary 4.4(a) shows
$\max_{i=1}^{n}\|\widetilde{R}_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\|_{\mathrm{op}}=\left\|\sum_{i=1}^{n}E_{ii}\otimes\widetilde{R}_{ii}^{(S)}-\mathsf{r}_{0}\right\|_{\mathrm{op}}\leq 2\delta^{-2}\|\Delta^{(S)}\|_{\mathrm{op}}\leq 2\delta^{-1}C(\gamma,\delta)n^{-\alpha}.$ (87)
Then specializing this to $i\notin S$ and adjusting the value of $\alpha$ yields the first statement of (77). ∎
4.3. Iterative bootstrapping
We now improve the preceding estimates of Lemma 4.9 in the $L^{p}$-norms for $p<\infty$, using the following bootstrapping lemma.
Lemma 4.10.
Suppose Assumption 4.1 holds. Suppose also that, for some $\alpha\in(0,1/2)$ and any fixed $l\geq 1$, uniformly over $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and over $i,j\notin S$ with $i\neq j$, we have
$R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\prec n^{-\alpha},\qquad R_{ij}^{(S)}\prec n^{-\alpha}.$ (88)
Set $\alpha^{\prime}=\min(\frac{3\alpha}{2},\frac{1}{2})$. Then for any fixed $l\geq 1$, uniformly over $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and over $i,j\notin S$ with $i\neq j$,
$R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\prec n^{-\alpha^{\prime}},\qquad R_{ij}^{(S)}\prec n^{-\alpha^{\prime}}.$
Before proving this lemma, we derive an estimate (Lemma 4.12 below) that we will use in conjunction with the fluctuation averaging result of Lemma 3.5. Define
$\mathbb{E}_{i}[\cdot]=\mathbb{E}[\;\cdot\mid H^{(i)}],\qquad\mathcal{Q}_{i}=1-\mathbb{E}_{i},$ (89)
where $\mathbb{E}_{i}$ is the expectation over the entries in only row and column $i$ of $H$. Note that $\mathbb{E}_{i}\mathbb{E}_{j}=\mathbb{E}_{j}\mathbb{E}_{i}$ for all $i\neq j$, so $\{\mathbb{E}_{i},\mathcal{Q}_{i}:i=1,\ldots,n\}$ form a commuting system of projections. Set $\mathcal{Q}_{T}=\prod_{i\in T}\mathcal{Q}_{i}$.
Lemma 4.11.
Under Assumption 4.1, for any fixed $l\geq 1$, uniformly over $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and $i\notin S$,
$(R_{ii}^{(S)})^{-1}\prec 1.$ (90)
Proof.
Let $\mathbf{h}_{i}^{(iS)}$ be the $i^{\text{th}}$ column of $H^{(iS)}$, having entries $(\mathbf{h}_{i}^{(iS)})_{j}=h_{ji}$ if $j\notin S\cup\{i\}$ and 0 otherwise. Lemma 3.6(a) gives
$\|(R_{ii}^{(S)})^{-1}\|_{\mathrm{op}}$ $\leq|h_{ii}|\left\|\mathsf{x}\right\|_{\mathrm{op}}+\left\|\mathsf{z}_{i}\right\|_{\mathrm{op}}+\|(\mathbf{h}_{i}^{(iS)}\otimes\mathsf{x})^{*}R^{(iS)}(\mathbf{h}_{i}^{(iS)}\otimes\mathsf{x})\|_{\mathrm{op}}$
$\leq|h_{ii}|\left\|\mathsf{x}\right\|_{\mathrm{op}}+\left\|\mathsf{z}_{i}\right\|_{\mathrm{op}}+\|R^{(iS)}\|_{\mathrm{op}}\left\|\mathsf{x}\right\|_{\mathrm{op}}^{2}\|\mathbf{h}_{i}^{(iS)}\|_{2}^{2}.$
We have $\|\mathsf{x}\|_{\mathrm{op}},\|\mathsf{z}_{i}\|_{\mathrm{op}}\prec 1$ by (43), $h_{ii}\prec n^{-1/2}$ by (16), and $\|\mathbf{h}_{i}^{(iS)}\|_{2}\prec 1$ by Lemma 3.4(a) in the scalar case of $\mathcal{X}=\mathbb{C}$, so the result follows. ∎
Lemma 4.12.
For any $S\subset\{1,\ldots,n\}$ and $r,j\notin S$ with $r\neq j$, define
$\mathcal{Z}_{rj}^{(S)}=\mathcal{Q}_{r}[R_{rj}^{(S)}R_{rj}^{(S)*}].$
Fix any $l\geq 1$. Under the assumptions of Lemma 4.10, uniformly over subsets $S\subset\{1,\ldots,n\}$ with $|S|\leq l$, $r,j\notin S$ with $r\neq j$, and $T\subset\{1,\ldots,n\}\setminus(S\cup\{r,j\})$ with $|T|\leq l$,
$\mathcal{Q}_{T}[\mathcal{Z}_{rj}^{(S)}]\prec n^{-(2+|T|)\alpha}.$
Proof.
For $|T|=0$, we have $R_{rj}^{(S)}R_{rj}^{(S)*}\prec n^{-2\alpha}$ by the assumption (88) and Hölder’s inequality. Then also $\mathcal{Z}_{rj}^{(S)}=R_{rj}^{(S)}R_{rj}^{(S)*}-\mathbb{E}_{r}[R_{rj}^{(S)}R_{rj}^{(S)*}]\prec n^{-2\alpha}$ by Lemma 3.2(c), so the assertion holds for $|T|=0$.
If $|T|\geq 1$, suppose $i_{1}\in T$. Then Lemma 3.6(c) gives
$R_{rj}^{(S)}R_{rj}^{(S)*}$ $=\underbrace{R_{rj}^{(i_{1}S)}R_{rj}^{(i_{1}S)*}}_{=L(\{i_{1}\})}$
$\quad+\underbrace{\Big{(}R_{ri_{1}}^{(S)}\frac{1}{R_{i_{1}i_{1}}^{(S)}}R_{i_{1}j}^{(S)}\Big{)}R_{rj}^{(i_{1}S)*}+R_{rj}^{(i_{1}S)}\Big{(}R_{ri_{1}}^{(S)}\frac{1}{R_{i_{1}i_{1}}^{(S)}}R_{i_{1}j}^{(S)}\Big{)}^{*}+\Big{(}R_{ri_{1}}^{(S)}\frac{1}{R_{i_{1}i_{1}}^{(S)}}R_{i_{1}j}^{(S)}\Big{)}\Big{(}R_{ri_{1}}^{(S)}\frac{1}{R_{i_{1}i_{1}}^{(S)}}R_{i_{1}j}^{(S)}\Big{)}^{*}}_{=P(\{i_{1}\})}.$
Here, the first term $L(\{i_{1}\})$ is independent of the entries of row and column $i_{1}$ of $H$, so $\mathcal{Q}_{i_{1}}[L(\{i_{1}\})]=0$. The remaining terms constituting $P(\{i_{1}\})$ each have at least 3 off-diagonal resolvent factors, i.e. factors of the form $R_{pq}^{(S^{\prime})}$ or $R_{pq}^{(S^{\prime})*}$ for some $p\neq q$ and $p,q\notin S^{\prime}$, so (90) and (88) imply $P(\{i_{1}\})\prec n^{-3\alpha}$.
Now if $|T|\geq 2$ and $i_{2}\in T$ with $i_{2}\neq i_{1}$, we apply Lemma 3.6(c) again to expand each factor of $P(\{i_{1}\})$ over $i_{2}$, yielding
$R_{rj}^{(S)}R_{rj}^{(S)*}=L(\{i_{1},i_{2}\})+P(\{i_{1},i_{2}\})$
where
$L(\{i_{1},i_{2}\})$ $=L(\{i_{1}\})+\Big{(}R_{ri_{1}}^{(i_{2}S)}\frac{1}{R_{i_{1}i_{1}}^{(i_{2}S)}}R_{i_{1}j}^{(i_{2}S)}\Big{)}R_{rj}^{(i_{1}i_{2}S)*}+R_{rj}^{(i_{1}i_{2}S)}\Big{(}R_{ri_{1}}^{(i_{2}S)}\frac{1}{R_{i_{1}i_{1}}^{(i_{2}S)}}R_{i_{1}j}^{(i_{2}S)}\Big{)}^{*}$
$\hskip 72.26999pt+\Big{(}R_{ri_{1}}^{(i_{2}S)}\frac{1}{R_{i_{1}i_{1}}^{(i_{2}S)}}R_{i_{1}j}^{(i_{2}S)}\Big{)}\Big{(}R_{ri_{1}}^{(i_{2}S)}\frac{1}{R_{i_{1}i_{1}}^{(i_{2}S)}}R_{i_{1}j}^{(i_{2}S)}\Big{)}^{*}$
and $P(\{i_{1},i_{2}\})$ collects all remaining terms in the expansion of $P(\{i_{1}\})$. Each term of $L(\{i_{1},i_{2}\})$ is independent of the entries of either row and column $i_{1}$ or $i_{2}$ of $H$, so $Q_{\{i_{1},i_{2}\}}[L(\{i_{1},i_{2}\})]=0$. Each term of $P(\{i_{1},i_{2}\})$ has at least 4 off-diagonal resolvent factors, so $P(\{i_{1},i_{2}\})\prec n^{-4\alpha}$. Inductively applying Lemma (3.6)(c) to expand $P(\{i_{1},\ldots,i_{k}\})$ in each successive index $i_{k+1}$ of $T$, this shows
$R_{rj}^{(S)}R_{rj}^{(S)*}=L(T)+P(T)$
where
  • Each term in $L(T)$ is independent of the entries of row and column $i$ of $H$ for at least one index $i\in T$, so $\mathcal{Q}_{T}[L(T)]=0$.
  • $P(T)$ is a sum of at most $C_{l}$ summands, each summand a product of at most $C_{l}$ factors, for a constant $C_{l}>0$ depending only on the given upper bound $l$ for $|T|$.
  • Each term of $P(T)$ has at least $2+|T|$ off-diagonal resolvent factors, and hence by (88) is of size $O_{\prec}(n^{-(2+|T|)\alpha})$.
Then $P(T)\prec n^{-(2+|T|)\alpha}$, uniformly over $T\subset\{1,\ldots,n\}\setminus(S\cup\{r,j\})$ with $|T|\leq l$. This implies $\mathcal{Q}_{T\cup\{r\}}[L(T)]=0$ and $\mathcal{Q}_{T\cup\{r\}}[P(T)]\prec n^{-(2+|T|)\alpha}$ by Lemma 3.2(c), so $\mathcal{Q}_{T}[\mathcal{Z}_{rj}^{(T)}]=\mathcal{Q}_{T\cup\{r\}}[R_{rj}^{(T)}R_{rj}^{(T)*}]\prec n^{-(2+|T|)\alpha}$ as desired. ∎
Proof of Lemma 4.10.
In view of Lemma 3.6(b) for the form of $R_{ij}^{(S)}$, we consider first the quantity
$\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}$
for $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and $i,j\notin S$ with $i\neq j$. Recalling the assumption $\mathbb{E}[(\sqrt{n}|h_{ij}|)^{p}]\leq C_{p}$ and applying Lemma 3.4(a) conditional on $H^{(iS)}$, for any $p\in[2,\infty)$ and a constant $C_{p}>0$,
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\bigg{\|}_{p}^{p}\bigg{]}\leq C_{p}n^{-p/2}\max\bigg{\{}\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r}^{(iS)}R_{rj}^{(iS)}R_{rj}^{(iS)*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]},\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r}^{(iS)}R_{rj}^{(iS)*}R_{rj}^{(iS)}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}\bigg{\}}.$ (91)
The second bound of (91) may be controlled spectrally: Using that $R^{(iS)}_{rj}=0$ for $r\in S\cup\{i\}$ and $j\notin S\cup\{i\}$,
$\bigg{\|}\sum_{r}^{(iS)}R_{rj}^{(iS)*}R_{rj}^{(iS)}\bigg{\|}_{\mathrm{op}}=\Big{\|}(\mathbf{e}_{j}\otimes 1)R^{(iS)*}R^{(iS)}(\mathbf{e}_{j}\otimes 1)\Big{\|}_{\mathrm{op}}\leq\|R^{(iS)}\|_{\mathrm{op}}^{2}\leq\delta^{-2}.$
Thus, applying monotonicity of the $L^{p}$-norm in $p$ (Lemma D.3),
$\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r}^{(iS)}R_{rj}^{(iS)*}R_{rj}^{(iS)}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}\leq\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}R_{rj}^{(iS)*}R_{rj}^{(iS)}\bigg{\|}_{\mathrm{op}}^{p/2}\bigg{]}\leq\delta^{-p}.$ (92)
For the first bound of (91), let us write
$\sum_{r}^{(iS)}R_{rj}^{(iS)}R_{rj}^{(iS)*}=R_{jj}^{(iS)}R_{jj}^{(iS)*}+\sum_{r}^{(ijS)}\mathbb{E}_{r}[R_{rj}^{(iS)}R_{rj}^{(iS)*}]+\sum_{r}^{(ijS)}\mathcal{Q}_{r}[R_{rj}^{(iS)}R_{rj}^{(iS)*}].$
Applying $\|R_{jj}^{(iS)}\|_{\mathrm{op}}\leq\delta^{-1}$ for the first term, and Lemma 3.5(a) with the estimates of Lemma 4.12 for the third term, this gives
$\sum_{r}^{(iS)}R_{rj}^{(iS)}R_{rj}^{(iS)*}=\sum_{r}^{(ijS)}\mathbb{E}_{r}[R_{rj}^{(iS)}R_{rj}^{(iS)*}]+O_{\prec}(1+n^{1-3\alpha}).$ (93)
We expand the resolvent for the remaining term of (93), applying Lemma 3.4(b) to write, for any $r\notin S\cup\{i,j\}$,
$R_{rj}^{(iS)}={-}R_{rr}^{(iS)}\mathsf{x}\sum_{s}^{(irS)}h_{rs}R_{sj}^{(irS)}.$
Here $R_{rj}^{(iS)}\prec n^{-\alpha}$ by the second statement of (88), so multiplying by $(R_{rr}^{(iS)})^{-1}$ and $\mathsf{x}^{-1}$ and applying (90) and $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$, also $\sum_{s}^{(irS)}h_{rs}R_{sj}^{(irS)}\prec n^{-\alpha}$. Then the first statement of (88) gives
$R_{rj}^{(iS)}={-}(\mathsf{r}_{0})_{rr}\mathsf{x}\sum_{s}^{(irS)}h_{rs}R_{sj}^{(irS)}+O_{\prec}(n^{-2\alpha}).$
Applying this and independence of $R_{sj}^{(irS)}$ with the variables in row/column $r$ of $H$,
$\mathbb{E}_{r}[R_{rj}^{(iS)}R_{rj}^{(iS)*}]$ $=(\mathsf{r}_{0})_{rr}\mathsf{x}\bigg{(}\sum_{s,t}^{(irS)}\mathbb{E}[h_{rs}h_{tr}]R_{sj}^{(irS)}R_{jt}^{(irS)*}\bigg{)}\mathsf{x}(\mathsf{r}_{0})_{rr}^{*}+O_{\prec}(n^{-3\alpha})$
$=\frac{1}{n}(\mathsf{r}_{0})_{rr}\mathsf{x}\bigg{(}\sum_{s}^{(irS)}R_{sj}^{(irS)}R_{js}^{(irS)*}\bigg{)}\mathsf{x}(\mathsf{r}_{0})_{rr}^{*}+O_{\prec}(n^{-3\alpha}).$
Recalling $r\notin S\cup\{i,j\}$, we have
$\sum_{s}^{(irS)}R_{sj}^{(irS)}R_{js}^{(irS)*}$ $=R_{jj}^{(irS)}R_{jj}^{(irS)*}+\sum_{s}^{(ijrS)}R_{sj}^{(irS)}R_{js}^{(irS)*}$
$=R_{jj}^{(irS)}R_{jj}^{(irS)*}+\sum_{s}^{(ijrS)}R_{sj}^{(iS)}R_{js}^{(iS)*}+O_{\prec}(n^{1-3\alpha})$
$=\sum_{s}^{(iS)}R_{sj}^{(iS)}R_{js}^{(iS)*}+O_{\prec}(1+n^{1-3\alpha}),$
the second line using $R_{sj}^{(irS)}=R_{sj}^{(iS)}+O_{\prec}(n^{-2\alpha})$ for $r\notin S\cup\{i,j,s\}$ by Lemma 3.6(c), (88), and (90), and the third line using $\|R_{jj}^{(irS)}\|_{\mathrm{op}}\leq\delta^{-1}$ and $\|R_{sj}^{(iS)}\|_{\mathrm{op}}\leq\delta^{-1}$ for $s\in\{j,r\}$. So
$\mathbb{E}_{r}[R_{rj}^{(iS)}R_{rj}^{(iS)*}]=\frac{1}{n}(\mathsf{r}_{0})_{rr}\mathsf{x}\bigg{(}\sum_{s}^{(iS)}R_{sj}^{(iS)}R_{js}^{(iS)*}\bigg{)}\mathsf{x}(\mathsf{r}_{0})_{rr}^{*}+O_{\prec}(n^{-1}+n^{-3\alpha}).$
Summing over $r\notin S\cup\{i,j\}$, this gives
$\sum_{r}^{(ijS)}\mathbb{E}_{r}[R_{rj}^{(iS)}R_{rj}^{(iS)*}]$ $=\sum_{r}^{(ijS)}\frac{1}{n}(\mathsf{r}_{0})_{rr}\mathsf{x}\bigg{(}\sum_{s}^{(iS)}R_{sj}^{(iS)}R_{js}^{(iS)*}\bigg{)}\mathsf{x}(\mathsf{r}_{0})_{rr}^{*}+O_{\prec}(1+n^{1-3\alpha})$
$=\sum_{r=1}^{n}\frac{1}{n}(\mathsf{r}_{0})_{rr}\mathsf{x}\bigg{(}\sum_{s}^{(iS)}R_{sj}^{(iS)}R_{js}^{(iS)*}\bigg{)}\mathsf{x}(\mathsf{r}_{0})_{rr}^{*}+O_{\prec}(1+n^{1-3\alpha}),$
where the second step applies the trivial bound $\|\sum_{s}^{(iS)}R_{sj}^{(iS)}R_{js}^{(iS)*}\|_{\mathrm{op}}\leq n\|R^{(iS)}\|_{\mathrm{op}}^{2}\leq n\delta^{-2}$ to include the summands for $r\in S\cup\{i,j\}$ with an additional $O_{\prec}(1)$ error. Applying this back to (93), we obtain
$\mathcal{L}_{2}\bigg{(}\sum_{r}^{(iS)}R_{rj}^{(iS)}R_{rj}^{(iS)*}\bigg{)}\prec 1+n^{1-3\alpha}$
where here $\mathcal{L}_{2}$ is the linear operator of Lemma 4.7. By the quantitative invertibility of $\mathcal{L}_{2}$ established in Lemma 4.7, this implies
$\sum_{r}^{(iS)}R_{rj}^{(iS)}R_{rj}^{(iS)*}\prec 1+n^{1-3\alpha}.$
Then, for any fixed $p\in[2,\infty)$, we get (from this and Lemma 3.2)
$\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r}^{(iS)}R_{rj}^{(iS)}R_{rj}^{(iS)*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}\prec 1+n^{(1-3\alpha)(p/2)},$ (94)
which controls the first bound of (91).
Then, applying (94) and (92) to (91), for any fixed $p\in[2,\infty)$ and all $n\geq n_{0}(p)$,
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\bigg{\|}_{p}^{p}\bigg{]}\leq n^{-p\alpha^{\prime}+0.1}$
where we have set $\alpha^{\prime}=\min(\frac{3\alpha}{2},\frac{1}{2})$. Then, for any fixed $q\in[1,\infty)$ and $\epsilon,D>0$, choosing $p\geq q$ large enough so that $p\epsilon>D+0.1$ and applying Markov’s inequality and monotonicity of $\|\cdot\|_{p}$ in $p$ (Lemma D.3), for all $n\geq n_{0}(q,\epsilon,D)$,
$\mathbb{P}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\bigg{\|}_{q}\geq n^{-\alpha^{\prime}+\epsilon}\bigg{]}\leq n^{p\alpha^{\prime}-p\epsilon}\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\bigg{\|}_{q}^{p}\bigg{]}\leq n^{p\alpha^{\prime}-p\epsilon}\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\bigg{\|}_{p}^{p}\bigg{]}<n^{-D}.$ (95)
This shows $\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\prec n^{-\alpha^{\prime}}$. Then by Lemma 3.6(b) and the bound $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$, we have the improved estimate (uniformly over $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and $i,j\notin S$ with $i\neq j$)
$R_{ij}^{(S)}\prec n^{-\alpha^{\prime}}.$ (96)
Now to show $R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\prec n^{-\alpha^{\prime}}$, recall from (80) that
$\widetilde{R}_{ii}^{(S)}=\bigg{(}{-}\mathsf{z}_{i}-\Delta_{i}^{(S)}-\mathsf{x}\bigg{[}\frac{1}{n}\sum_{j=1}^{n}\widetilde{R}_{jj}^{(S)}\bigg{]}\mathsf{x}\bigg{)}^{-1}$
where $\widetilde{R}_{ii}^{(S)}=R_{ii}^{(S)}$ for all $i\notin S$, and $\Delta_{i}^{(S)}$ is the error defined by (81) and (82). For the term $\mathrm{I}$ of (81), applying Lemma 3.4(a) conditional on $H^{(iS)}$ gives, for any $p\geq 2$ and a constant $C_{p}>0$,
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}\bigg{\|}_{p}^{p}\bigg{]}$
$\leq C_{p}n^{-p}\max\bigg{\{}\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r}^{(iS)}R_{rr}^{(iS)}R_{rr}^{(iS)*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]},\;\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r}^{(iS)}R_{rr}^{(iS)*}R_{rr}^{(iS)}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}\bigg{\}}$
$\leq C_{p}n^{-p}\max\bigg{\{}\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r,s=1}^{n}R_{rs}^{(iS)}R_{rs}^{(iS)*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]},\;\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r,s=1}^{n}R_{rs}^{(iS)*}R_{rs}^{(iS)}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}\bigg{\}}$
where the second inequality applies monotonicity of the operator square-root $0\leq\mathsf{x}\leq\mathsf{y}\Rightarrow\mathsf{x}^{1/2}\leq\mathsf{y}^{1/2}$ and monotonicity of the $L^{p}$-norm over the positive cone (Lemma D.3). Similarly applying Lemma 3.4(c) conditional on $H^{(iS)}$ for the summation over $r\neq s$, the term $\mathrm{I}$ of (81) is bounded as
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}+\sum_{r\neq s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}\bigg{\|}_{p}^{p}\bigg{]}$
$\leq C_{p}n^{-p}\max\bigg{\{}\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r,s=1}^{n}R_{rs}^{(iS)}R_{rs}^{(iS)*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]},\;\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r,s=1}^{n}R_{rs}^{(iS)*}R_{rs}^{(iS)}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]},$
$\hskip 144.54ptn\,\mathbb{E}[\|R^{(iS)}-\operatorname{diag}(R^{(iS)})\|_{p}^{p}],\;n\,\mathbb{E}[\|{R^{(iS)}}^{\mathsf{t}}-\operatorname{diag}(R^{(iS)})\|_{p}^{p}]\bigg{\}}$ (97)
where ${R^{(iS)}}^{\mathsf{t}}=\sum_{k,l=1}^{n}E_{kl}\otimes R^{(iS)}_{lk}$ is the partial transpose of $R^{(iS)}$ as defined in (38), and here $\operatorname{diag}(R^{(iS)})=\sum_{k=1}^{n}E_{kk}\otimes R_{kk}^{(iS)}$ is the operator that has only the diagonal entries of $R^{(iS)}$. The first three bounds on the right side of (97) may be controlled spectrally: For the third bound,
$n\,\mathbb{E}[\|R^{(iS)}-\operatorname{diag}(R^{(iS)})\|_{p}^{p}]\leq n\,\mathbb{E}[(\|R^{(iS)}\|_{\mathrm{op}}+\|\operatorname{diag}(R^{(iS)})\|_{\mathrm{op}})^{p}]\leq n(2\delta^{-1})^{p}.$ (98)
For the first and second bounds,
$\bigg{\|}\sum_{r,s=1}^{n}R_{rs}^{(iS)}R_{rs}^{(iS)*}\bigg{\|}_{\mathrm{op}}$ $=n\,\Big{\|}(n^{-1}\operatorname{Tr}\otimes 1)R^{(iS)}R^{(iS)*}\Big{\|}_{\mathrm{op}}\leq n\|R^{(iS)}\|_{\mathrm{op}}^{2}\leq n\delta^{-2},$
$\bigg{\|}\sum_{r,s=1}^{n}R_{rs}^{(iS)*}R_{rs}^{(iS)}\bigg{\|}_{\mathrm{op}}$ $=n\,\Big{\|}(n^{-1}\operatorname{Tr}\otimes 1)R^{(iS)*}R^{(iS)}\Big{\|}_{\mathrm{op}}\leq n\|R^{(iS)}\|_{\mathrm{op}}^{2}\leq n\delta^{-2}.$
Thus
$\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r,s=1}^{n}R_{rs}^{(iS)}R_{rs}^{(iS)*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]},\;\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{r,s=1}^{n}R_{rs}^{(iS)*}R_{rs}^{(iS)}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}\leq n^{p/2}\delta^{-p}.$ (99)
For the fourth bound $n\,\mathbb{E}[\|{R^{(iS)}}^{\mathsf{t}}-\operatorname{diag}(R^{(iS)})\|_{p}^{p}]$ of (97), we apply the following argument: For any $M=\sum_{j,k}E_{jk}\otimes M_{jk}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$, note that
$M^{\mathsf{t}}-\operatorname{diag}(M)=\sum_{1\leq j\neq k\leq n}E_{jk}\otimes M_{kj}=\sum_{j=1}^{n-1}M[j],\qquad\text{ for }M[j]=\sum_{k=1}^{n}E_{j+k,k}\otimes M_{k,j+k}$
where $j+k$ is interpreted modulo $n$. Here
$\|M[j]\|_{p}^{p}$ $=\|(M[j]M[j]^{*})^{1/2}\|_{p}^{p}=\bigg{\|}\sum_{k=1}^{n}E_{j+k,j+k}\otimes(M_{k,j+k}M_{k,j+k}^{*})^{1/2}\bigg{\|}_{p}^{p}$
$=(n^{-1}\operatorname{Tr}\otimes\tau)\sum_{k=1}^{n}E_{j+k,j+k}\otimes(M_{k,j+k}M_{k,j+k}^{*})^{p/2}=\frac{1}{n}\sum_{k=1}^{n}\|M_{k,j+k}\|_{p}^{p}\leq\max_{k=1}^{n}\|M_{k,j+k}\|_{p}^{p},$
so
$\|M^{\mathsf{t}}-\operatorname{diag}(M)\|_{p}\leq\sum_{j=1}^{n-1}\|M[j]\|_{p}\leq\sum_{j=1}^{n-1}\Big{(}\max_{k=1}^{n}\|M_{k,j+k}\|_{p}\Big{)}.$
Applying this to $M=R^{(iS)}$, observe that for each $j=1,\ldots,n-1$ we have $R^{(iS)}_{k,j+k}\prec n^{-\alpha^{\prime}}$ if $k,j+k\notin S\cup\{i\}$ as already shown in (96), and $R^{(iS)}_{k,j+k}=0$ by definition if $k\in S\cup\{i\}$ or $j+k\in S\cup\{i\}$ with $k\neq j+k$. Thus $\max_{k}\|R_{k,j+k}^{(iS)}\|_{p}\prec n^{-\alpha^{\prime}}$, so $\|{R^{(iS)}}^{\mathsf{t}}\|_{p}\prec n^{1-\alpha^{\prime}}$, and
$n\,\mathbb{E}[\|{R^{(iS)}}^{\mathsf{t}}\|_{p}^{p}]\prec n\cdot n^{p(1-\alpha^{\prime})}.$ (100)
Applying (98), (99), and (100) to (97), for any fixed $p\in[2,\infty)$ and all $n\geq n_{0}(p)$,
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}+\sum_{r\neq s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}\bigg{\|}_{p}^{p}\bigg{]}\leq n^{\max(1-p,-p/2,1-p\alpha^{\prime})+0.1}=n^{-p\alpha^{\prime}+1.1},$
the last equality using $\alpha^{\prime}\leq 1/2$. Then, for any fixed $q\in[1,\infty)$ and $\epsilon,D>0$, applying this with $p\geq q$ large enough so that $p\epsilon>D+1.1$, we obtain similarly to (95)
$\mathbb{P}\bigg{[}\bigg{\|}\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}+\sum_{r\neq s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}\bigg{\|}_{q}\geq n^{-\alpha^{\prime}+\epsilon}\bigg{]}<n^{-D}.$
Hence, for the term $\mathrm{I}$ defining $\Delta_{i}^{(S)}$ in (81), we have $\mathrm{I}\prec n^{-\alpha^{\prime}}$.
For the other terms of (81), we have $\mathrm{II},\mathrm{IV}\prec n^{-1}$ by (83), and $\mathrm{III}\prec n^{-2\alpha^{\prime}}$ by its form (84) and (96) and (90). Applying also $h_{ii}\prec n^{-1/2}$, this gives (uniformly over $i=1,\ldots,n$)
$\Delta_{i}^{(S)}\prec n^{-\alpha^{\prime}}.$ (101)
Then, setting $\Delta^{(S)}=\sum_{i}E_{ii}\otimes\Delta_{i}^{(S)}$, we have $\|\Delta^{(S)}\|_{p}^{p}=n^{-1}\sum_{i}\|\Delta_{i}^{(S)}\|_{p}^{p}$ for any $p\in[1,\infty)$ so also $\|\Delta^{(S)}\|_{p}\prec n^{-\alpha^{\prime}}$ for each fixed $p$. Recalling the weak estimate for $\|\Delta^{(S)}\|_{\mathrm{op}}$ from (85) on the high probability event $\mathcal{E}$, which ensures $\Im(\mathsf{z}+\Delta^{(S)})\geq\delta/2$, this implies by (86) and Corollary 4.4(a) applied now in the norm $\|\cdot\|_{p}$ that
$\bigg{\|}\sum_{i=1}^{n}E_{ii}\otimes\widetilde{R}_{ii}^{(S)}-\mathsf{r}_{0}\bigg{\|}_{p}\prec n^{-\alpha^{\prime}}.$
Thus, for any $i\notin S$, any $q\in[1,\infty)$, and any $\epsilon,D>0$, choosing $p\geq q$ large enough such that $p\epsilon>1.1$,
$\|R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\|_{q}^{p}\leq\sum_{i=1}^{n}\|\widetilde{R}_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\|_{p}^{p}=n\bigg{\|}\sum_{i=1}^{n}E_{ii}\otimes\widetilde{R}_{ii}^{(S)}-\mathsf{r}_{0}\bigg{\|}_{p}^{p}\leq n^{-\alpha^{\prime}p+1.1}<n^{(-\alpha^{\prime}+\epsilon)p}$
with probability at least $1-n^{-D}$ for $n\geq n_{0}(q,\epsilon,D)$. This shows $R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\prec n^{-\alpha^{\prime}}$. ∎
4.4. Proof of Theorem 4.2
We now prove the main result of this section, Theorem 4.2.
Proof of Theorem 4.2, (4546).
Lemma 4.9 implies $R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\prec n^{-\alpha}$ and $R_{ij}^{(S)}\prec n^{-\alpha}$ for some $\alpha>0$, uniformly over sets $S$ with $|S|\leq l$ and $i,j\notin S$ with $i\neq j$. Then, iterating Lemma 4.10 a constant number of times, we get $R_{ii}^{(S)}-(\mathsf{r}_{0})_{ii}\prec n^{-1/2}$ and $R_{ij}^{(S)}\prec n^{-1/2}$, implying for $S=\emptyset$ the statements (4546). ∎
For the remaining statements of Theorem 4.2, we collect here several estimates needed for additional applications of fluctuation averaging (Lemma 3.5).
Lemma 4.13.
Under Assumption 4.1, define
$\mathcal{Z}_{i}$ $=(\mathsf{r}_{0})_{ii}\mathcal{Q}_{i}[R_{ii}^{-1}](\mathsf{r}_{0})_{ii},$
$\mathcal{Z}_{ij}$ $=\mathcal{Q}_{i}\mathcal{Q}_{j}[R_{ij}]\text{ for }i\neq j,$
$\mathcal{K}_{r}^{(ij)}$ $=(\mathsf{r}_{0})_{ii}\mathsf{x}(\mathsf{r}_{0})_{rr}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathcal{Q}_{r}[R_{rj}^{(i)}]\text{ for distinct}r,i,j.$
Fix any $l\geq 1$.
  1. (a)
    Uniformly over $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and $i\notin S$, $\mathcal{Q}_{S}[\mathcal{Z}_{i}]\prec n^{-1/2-|S|/2}$.
  2. (b)
    Uniformly over $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and $i,j\notin S$ with $i\neq j$, $\mathcal{Q}_{S}[\mathcal{Z}_{ij}]\prec n^{-1/2-|S|/2}$.
  3. (c)
    Uniformly over $S\subset\{1,\ldots,n\}$ with $|S|\leq l$ and distinct $i,j,r\notin S$, $\mathcal{Q}_{S}[\mathcal{K}_{r}^{(ij)}]\prec n^{-1/2-|S|/2}$.
Proof.
For (a), since $\mathsf{r}_{0}$ does not depend on $H$, we have $\mathcal{Q}_{S}[\mathcal{Z}_{i}]=(\mathsf{r}_{0})_{ii}\mathcal{Q}_{S\cup\{i\}}[R_{ii}^{-1}](\mathsf{r}_{0})_{ii}$, so by Hölder’s inequality and the bound $\|(\mathsf{r}_{0})_{ii}\|_{\mathrm{op}}\leq\delta$ it suffices to show
$\mathcal{Q}_{S}[\mathcal{Q}_{i}[R_{ii}^{-1}]]\prec n^{-1/2-|S|/2}.$
We remark that for $S=\emptyset$, this follows from $\mathcal{Q}_{i}[R_{ii}^{-1}]=-\Delta_{i}+O_{\prec}(n^{-1})\prec n^{-1/2}$ (c.f. (106) and (103) below). For (c), similarly it suffices to show
$\mathcal{Q}_{S}[\mathcal{Q}_{r}[R_{rj}^{(i)}]]\prec n^{-1/2-|S|/2}.$
The argument for all three parts is then the same as in Lemma 4.12, applying Lemma 3.6(c) to iteratively expand $R_{ii}^{-1}$, $R_{ij}$, and $R_{rj}^{(i)}$ in the indices of $S$, and applying now the optimal estimates $R_{pq}^{(S^{\prime})}\prec n^{-1/2}$ from (46) to bound the terms of the expansion. We omit further details for brevity. ∎
Proof of Theorem 4.2, (44).
Let us specialize (81) to $S=\emptyset$, and write simply $\Delta_{i}=\Delta_{i}^{(S)}$. Then
$\Delta_{i}={-}h_{ii}\mathsf{x}+\mathsf{x}\Bigg{[}\sum_{r}^{(i)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(i)}+\sum_{r\neq s}^{(i)}h_{ir}R_{rs}^{(i)}h_{si}\Bigg{]}\mathsf{x}+O_{\prec}(n^{-1}),$ (102)
using the bounds $\mathrm{II}\prec n^{-1}$ from (83) and and $\mathrm{III}\prec n^{-1}$ from (84), (46), and (90). Furthermore, (101) from the final iteration of bootstrapping with $\alpha^{\prime}=1/2$ shows
$\Delta_{i}\prec n^{-1/2}.$ (103)
Define
$\widehat{R}_{ii}=\bigg{(}{-}\mathsf{z}_{i}-\mathsf{x}\bigg{[}\frac{1}{n}\sum_{j=1}^{n}R_{jj}\bigg{]}\mathsf{x}\bigg{)}^{-1}.$ (104)
By (45) already shown, we have $R_{ii}=(\mathsf{r}_{0})_{ii}+O_{\prec}(n^{-1/2})$ uniformly in $i\in\{1,\ldots,n\}$. Then averaging over $i$ gives
$\frac{1}{n}\sum_{i=1}^{n}R_{ii}=\frac{1}{n}\operatorname{Tr}\otimes 1[\mathsf{r}_{0}]+O_{\prec}(n^{-1/2})=\mathsf{m}_{0}+O_{\prec}(n^{-1/2}).$
Applying this, the identity $\mathsf{a}^{-1}-\mathsf{b}^{-1}=\mathsf{a}^{-1}(\mathsf{b}-\mathsf{a})\mathsf{b}^{-1}$, and $\|\widehat{R}_{ii}\|_{\mathrm{op}},\|(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma D.1, we may approximate $\widehat{R}_{ii}$ in (104) as
$\widehat{R}_{ii}=(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x}+O_{\prec}(n^{-1/2}))^{-1}=(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}+O_{\prec}(n^{-1/2})=(\mathsf{r}_{0})_{ii}+O_{\prec}(n^{-1/2}),$
the last identity using the characterization of $\mathsf{r}_{0}$ via the fixed-point equation of Lemma 4.3(a). Therefore, applying again $\mathsf{a}^{-1}-\mathsf{b}^{-1}=\mathsf{a}^{-1}(\mathsf{b}-\mathsf{a})\mathsf{b}^{-1}$ with $\widehat{R}_{ii}^{-1}=R_{ii}^{-1}+\Delta_{i}$ from (80), we have
$R_{ii}=\widehat{R}_{ii}+R_{ii}\Delta_{i}\widehat{R}_{ii}=\widehat{R}_{ii}+(\mathsf{r}_{0})_{ii}\Delta_{i}(\mathsf{r}_{0})_{ii}+O_{\prec}(n^{-1})$ (105)
where this applies (45) and (103) to bound the error.
Comparing the form (102) for $\Delta_{i}$ with the expansion of $R_{ii}^{-1}$ in Lemma 3.6(a), and noting that $\mathcal{Q}_{i}[\mathsf{z}_{i}]=0$ because $\mathsf{z}_{i}$ does not depend on $H$, we have
$\Delta_{i}={-}\mathcal{Q}_{i}[R_{ii}^{-1}]+O_{\prec}(n^{-1}).$ (106)
Thus, setting $\mathcal{Z}_{i}=(\mathsf{r}_{0})_{ii}\mathcal{Q}_{i}[R_{ii}^{-1}](\mathsf{r}_{0})_{ii}$ and averaging again over $i$,
$\frac{1}{n}\sum_{i=1}^{n}R_{ii}=\frac{1}{n}\sum_{i=1}^{n}\widehat{R}_{ii}-\frac{1}{n}\sum_{i=1}^{n}\mathcal{Z}_{i}+O_{\prec}(n^{-1}).$
By Lemma 3.5(a) (or (b)) applied with the estimates of Lemma 4.13(a), we have $n^{-1}\sum_{i}\mathcal{Z}_{i}\prec n^{-1}$. Thus, defining
$\widehat{\Delta}=\frac{1}{n}\sum_{i=1}^{n}R_{ii}-\frac{1}{n}\sum_{i=1}^{n}\widehat{R}_{ii}=\frac{1}{n}\sum_{i=1}^{n}R_{ii}-\frac{1}{n}\sum_{i=1}^{n}\bigg{(}{-}\mathsf{z}_{i}-\mathsf{x}\bigg{[}\frac{1}{n}\sum_{j=1}^{n}R_{jj}\bigg{]}\mathsf{x}\bigg{)}^{-1},$
this shows $\widehat{\Delta}\prec n^{-1}$. Observe also that
$\|\widehat{\Delta}\|_{\mathrm{op}}=\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}(R_{ii}-\widehat{R}_{ii})\bigg{\|}_{\mathrm{op}}=\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}R_{ii}\Delta_{i}\widehat{R}_{ii}\bigg{\|}_{\mathrm{op}}\leq\delta^{-2}\max_{i=1}^{n}\left\|\Delta_{i}\right\|_{\mathrm{op}}.$
By the weak bound (85) for $\|\Delta_{i}\|_{\mathrm{op}}$ on a high-probability event $\mathcal{E}$, this implies $\Im(\mathsf{z}_{i}+\mathsf{x}\widehat{\Delta}\mathsf{x})\geq\delta/2$ for all $i=1,\ldots,n$ and all $n\geq n_{0}(\gamma,\delta)$. On this event $\mathcal{E}$, Corollary 4.4(b) implies, for each $p\in[1,\infty)$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}R_{ii}-\mathsf{m}_{0}\right\|_{p}\leq(1+2\gamma^{2}\delta^{-2})\|\widehat{\Delta}\|_{p}.$
Thus $(n^{-1}\operatorname{Tr}\otimes 1)[R]-\mathsf{m}_{0}\prec n^{-1}$, showing (44). ∎
Proof of Theorem 4.2, (47).
By the polarization identity
$(\mathbf{u}\otimes 1)^{*}R(\mathbf{v}\otimes 1)$ $=\frac{1}{4}\Big{[}((\mathbf{v}+\mathbf{u})\otimes 1)^{*}R((\mathbf{v}+\mathbf{u})\otimes 1)-((\mathbf{v}-\mathbf{u})\otimes 1)^{*}R((\mathbf{v}-\mathbf{u})\otimes 1)$
$\qquad+i((\mathbf{v}+i\mathbf{u})\otimes 1)^{*}R((\mathbf{v}+i\mathbf{u})\otimes 1)-i((\mathbf{v}-i\mathbf{u})\otimes 1)^{*}R((\mathbf{v}-i\mathbf{u})\otimes 1)\Big{]}$
it suffices to prove the result uniformly over $\mathbf{u}=\mathbf{v}\in\mathbb{C}^{n}$ satisfying $\|\mathbf{v}\|_{2}\leq\sqrt{2}\upsilon$. We have
$(\mathbf{v}\otimes 1)^{*}R(\mathbf{v}\otimes 1)$ $=\sum_{i}|v_{i}|^{2}R_{ii}+\sum_{i\neq j}\bar{v}_{i}v_{j}R_{ij}$
$=\sum_{i}|v_{i}|^{2}(\mathsf{r}_{0})_{ii}+\sum_{i\neq j}\bar{v}_{i}v_{j}R_{ij}+O_{\prec}(n^{-1/2})$
$=(\mathbf{v}\otimes 1)^{*}\mathsf{r}_{0}(\mathbf{v}\otimes 1)+\sum_{i\neq j}\bar{v}_{i}v_{j}R_{ij}+O_{\prec}(n^{-1/2})$
where the second line applies $\|\mathbf{v}\|_{2}\prec 1$ and (45). It remains to show $\sum_{i\neq j}\bar{v}_{i}v_{j}R_{ij}\prec n^{-1/2}$.
Separating $R_{ij}$ into its conditional mean and fluctuations,
$\sum_{i\neq j}\bar{v}_{i}v_{j}R_{ij}=\sum_{i\neq j}\bar{v}_{i}v_{j}\mathbb{E}_{i}\mathbb{E}_{j}[R_{ij}]+\sum_{i\neq j}\bar{v}_{i}v_{j}\mathbb{E}_{i}\mathcal{Q}_{j}[R_{ij}]+\sum_{i\neq j}\bar{v}_{i}v_{j}\mathbb{E}_{j}\mathcal{Q}_{i}[R_{ij}]+\sum_{i\neq j}\bar{v}_{i}v_{j}\mathcal{Q}_{i}\mathcal{Q}_{j}[R_{ij}].$
We first examine $\mathbb{E}_{i}[R_{ij}]$. Recall $\widehat{R}_{ii}$ from (104). Observe that $R_{ii}=(\mathsf{r}_{0})_{ii}+O_{\prec}(n^{-1/2})$ by (45), whereas $\widehat{R}_{ii}=(\mathsf{r}_{0})_{ii}+O_{\prec}(n^{-1})$ with the smaller error $n^{-1}$ by its definition and the estimate (44) already shown. Then, applying Lemma 3.6(b) and (105) to expand $R_{ij}$,
${-}\mathbb{E}_{i}[R_{ij}]$ $=\mathbb{E}_{i}\left[R_{ii}\mathsf{x}\sum_{r}^{(i)}h_{ir}R_{rj}^{(i)}\right]=\mathbb{E}_{i}\left[\widehat{R}_{ii}\mathsf{x}\sum_{r}^{(i)}h_{ir}R_{rj}^{(i)}\right]+\mathbb{E}_{i}\left[R_{ii}\Delta_{i}\widehat{R}_{ii}\mathsf{x}\sum_{r}^{(i)}h_{ir}R_{rj}^{(i)}\right]$
$=\underbrace{\mathbb{E}_{i}\left[(\mathsf{r}_{0})_{ii}\mathsf{x}\sum_{r}^{(i)}h_{ir}R_{rj}^{(i)}\right]}_{=\mathrm{I}}+\underbrace{\mathbb{E}_{i}\left[(\mathsf{r}_{0})_{ii}\Delta_{i}(\mathsf{r}_{0})_{ii}\mathsf{x}\sum_{r}^{(i)}h_{ir}R_{rj}^{(i)}\right]}_{=\mathrm{II}}+\underbrace{\mathbb{E}_{i}\left[O_{\prec}(n^{-1})\cdot\mathsf{x}\sum_{r}^{(i)}h_{ir}R_{rj}^{(i)}\right]}_{=\mathrm{III}}.$
In these expressions, by (46), (90), and Lemma 3.6(b), we have
$\sum_{r}^{(i)}h_{ir}R_{rj}^{(i)}={-}\mathsf{x}^{-1}R_{ii}^{-1}R_{ij}\prec n^{-1/2}$ (107)
so $\mathrm{III}\prec n^{-3/2}$. Since $R^{(i)}_{rj}$ is a function only of $H^{(i)}$ (and $\mathsf{r}_{0},\mathsf{x}$ do not depend on $H$), it follows from $\mathbb{E}[h_{ir}]=0$ that $\mathrm{I}=0$. For $\mathrm{II}$, recall the form of $\Delta_{i}$ from (102). Substituting this expression of $\Delta_{i}$ into $\mathrm{II}$ gives
$\mathrm{II}$ $=-(\mathsf{r}_{0})_{ii}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}\sum_{r}^{(i)}\mathbb{E}_{i}[h_{ii}h_{ir}]R_{rj}^{(i)}+\sum_{t,r}^{(i)}\mathbb{E}_{i}\left[\left(|h_{it}|^{2}-\frac{1}{n}\right)h_{ir}\right](\mathsf{r}_{0})_{ii}\mathsf{x}R_{tt}^{(i)}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}R_{rj}^{(i)}$
$\qquad+\sum_{r}^{(i)}\sum_{t\neq s}^{(i)}\mathbb{E}_{i}[h_{it}h_{si}h_{ir}](\mathsf{r}_{0})_{ii}\mathsf{x}R_{ts}^{(i)}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}R_{rj}^{(i)}+O_{\prec}(n^{-3/2}).$
The first term is 0 since $r\neq i$. Similarly, the second term is 0 for summands $r\neq t$, and the third term is 0 since at least one of $r,s,t$ is distinct from the other two. Thus
$\mathrm{II}=\sum_{r}^{(i)}\mathbb{E}_{i}[|h_{ir}|^{2}h_{ir}](\mathsf{r}_{0})_{ii}\mathsf{x}R_{rr}^{(i)}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}R_{rj}^{(i)}+O_{\prec}(n^{-3/2}).$
Note that the single summand for $r=j$ is $O_{\prec}(n^{-3/2})$, that all summands for $r\neq j$ are $O_{\prec}(n^{-2})$ by (46), and that $R_{rr}^{(i)}=R_{rr}+O_{\prec}(n^{-1})=(\mathsf{r}_{0})_{rr}+O_{\prec}(n^{-1/2})$. Then we may further write this as
$\mathrm{II}$ $=\sum_{r}^{(ij)}\mathbb{E}_{i}[|h_{ir}|^{2}h_{ir}](\mathsf{r}_{0})_{ii}\mathsf{x}R_{rr}^{(i)}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}R_{rj}^{(i)}+O_{\prec}(n^{-3/2})$
$=\sum_{r}^{(ij)}\mathbb{E}_{i}[|h_{ir}|^{2}h_{ir}](\mathsf{r}_{0})_{ii}\mathsf{x}(\mathsf{r}_{0})_{rr}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}R_{rj}^{(i)}+O_{\prec}(n^{-3/2})$
$=\sum_{r}^{(ij)}\mathbb{E}_{i}[|h_{ir}|^{2}h_{ir}](\mathsf{r}_{0})_{ii}\mathsf{x}(\mathsf{r}_{0})_{rr}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}(\mathbb{E}_{r}+\mathcal{Q}_{r})[R_{rj}^{(i)}]+O_{\prec}(n^{-3/2}).$
By Lemma 3.5(a) applied with the estimates of Lemma 4.13(c), we have
$\sum_{r}^{(ij)}\mathbb{E}_{i}[|h_{ir}|^{2}h_{ir}](\mathsf{r}_{0})_{ii}\mathsf{x}(\mathsf{r}_{0})_{rr}\mathsf{x}(\mathsf{r}_{0})_{ii}\mathsf{x}\mathcal{Q}_{r}[R_{rj}^{(i)}]\prec n^{-3/2}.$
We now examine $\mathbb{E}_{r}[R_{rj}^{(i)}]$. Again by Lemmas 3.6(b) and 3.2(c),
${-}\mathbb{E}_{r}[R_{rj}^{(i)}]=\mathbb{E}_{r}\left[R_{rr}^{(i)}\mathsf{x}\sum_{s}^{(ir)}h_{rs}R_{sj}^{(ir)}\right]=(\mathsf{r}_{0})_{rr}\mathsf{x}\mathbb{E}_{r}\left[\sum_{s}^{(ir)}h_{rs}R_{sj}^{(ir)}\right]+O_{\prec}(n^{-1}).$
This first term is 0, so $\mathbb{E}_{r}[R_{rj}^{(i)}]\prec n^{-1}$. Combining the above gives $\mathrm{II}\prec n^{-3/2}$, and hence $\mathbb{E}_{i}[R_{ij}]\prec n^{-3/2}$. By symmetry, also $\mathbb{E}_{j}[R_{ij}]\prec n^{-3/2}$, so we conclude that
$\sum_{i\neq j}\bar{v}_{i}v_{j}R_{ij}=\sum_{i\neq j}\bar{v}_{i}v_{j}\mathcal{Q}_{i}\mathcal{Q}_{j}[R_{ij}]+\sum_{i\neq j}\bar{v}_{i}v_{j}\cdot O_{\prec}(n^{-3/2})=\sum_{i\neq j}\bar{v}_{i}v_{j}\mathcal{Q}_{i}\mathcal{Q}_{j}[R_{ij}]+O_{\prec}(n^{-1/2}),$
where the second equality applies $\sum_{i}|v_{i}|\leq\sqrt{n}\|\mathbf{v}\|_{2}\prec\sqrt{n}$. Finally, by Lemma 3.5(c) applied with the estimates of Lemma 4.13(b), we have
$\sum_{i\neq j}\bar{v}_{i}v_{j}\mathcal{Q}_{i}\mathcal{Q}_{j}[R_{ij}]\prec\frac{1}{\sqrt{n}}\left(\sum_{i\neq j}|v_{i}|^{2}|v_{j}|^{2}\right)^{1/2}\prec n^{-1/2}.$
So $\sum_{i\neq j}\bar{v}_{i}v_{j}R_{ij}\prec n^{-1/2}$ as desired, completing the proof. ∎
5. Analysis of the Kronecker deformed Wigner model
We now prove Proposition 2.3 and Theorem 2.4. For spectral arguments $z\in\mathbb{C}^{+}$, recall the following quantities from Section 2:
$Q$ $=A\otimes I+I\otimes B+\Theta\otimes\Xi\in\mathbb{C}^{n^{2}\times n^{2}},$
$\mathsf{q}$ $=\mathsf{a}\otimes 1+1\otimes\mathsf{b}+\Theta\otimes\Xi\in\mathcal{A}\otimes\mathcal{A},$
$G(z)$ $=(Q-z\,I\otimes I)^{-1},\qquad m(z)=n^{-2}\operatorname{Tr}G(z),$
[…]
We first show Proposition 2.3 and all statements of Theorem 2.4 except the estimate $G_{ij,\alpha\beta}\prec n^{-1}$ of (27) using the analyses of Section 4.
Proof of Proposition 2.3.
We apply Lemma 4.3(b) with $\mathcal{X}=\mathcal{A}$, $\mathsf{x}=1_{\mathcal{A}}$, and
$\mathsf{z}={-}I\otimes\mathsf{b}-\Theta\otimes\Xi+z\,I\otimes 1_{\mathcal{A}}\in\mathbb{C}^{n\times n}\otimes\mathcal{A}.$
Then $\Im\mathsf{z}=(\Im z)(I\otimes 1)$, so $\mathsf{z}_{i}={-}\mathsf{b}-\theta_{i}\Xi+z\in\mathcal{A}^{+}$ for each $i=1,\ldots,n$. Then Lemma 4.3(b) ensures that $\mathsf{m}_{b}(z)=(\tau\otimes 1)[\mathsf{g}(z)]$ is the unique fixed point in $\mathcal{A}^{+}$ to the fixed-point equation (20). Similarly $\mathsf{m}_{a}(z)=(1\otimes\tau)[\mathsf{g}(z)]$ is the unique fixed point in $\mathcal{A}^{+}$ to (19), and the identity $m_{0}(z)=\tau[\mathsf{m}_{a}(z)]=\tau[\mathsf{m}_{b}(z)]$ follows from taking a second trace $\tau$ for either $\mathsf{m}_{a}$ or $\mathsf{m}_{b}$. Furthermore, Lemma 4.3(a) shows that
$(\tau^{\mathcal{D}}\otimes 1)[\mathsf{g}(z)]=({-}\mathsf{z}-I\otimes\mathsf{m}_{b}(z))^{-1}=\sum_{i=1}^{n}E_{ii}\otimes(\mathsf{b}+\theta_{i}\Xi-z-\mathsf{m}_{b}(z))^{-1}.$
Then applying $1\otimes\tau^{\mathcal{D}}$ shows
$G_{0}=\sum_{i=1}^{n}E_{ii}\otimes\tau^{\mathcal{D}}\big{[}(\mathsf{b}+\theta_{i}\Xi-z-\mathsf{m}_{b}(z))^{-1}\big{]},$
and similarly
$G_{0}=\sum_{\alpha=1}^{n}\tau^{\mathcal{D}}\big{[}(\mathsf{a}+\xi_{\alpha}\Theta-z-\mathsf{m}_{a}(z))^{-1}\big{]}\otimes E_{\alpha\alpha}.$
Proof of Theorem 2.4, (2126) and (28).
Define
$\tilde{\mathsf{g}}(z)=(\mathsf{a}\otimes I+1_{\mathcal{A}}\otimes B+\Theta\otimes\Xi-z1_{\mathcal{A}}\otimes I)^{-1}\in\mathcal{A}\otimes\mathbb{C}^{n\times n}$ (108)
and abbreviate $G=G(z)$, $\tilde{\mathsf{g}}=\tilde{\mathsf{g}}(z)$, and $\mathsf{g}=\mathsf{g}(z)$. Fix any $\epsilon>0$ and consider the event =$\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$. We apply Theorem 4.2 conditional on $B$ and this event $\mathcal{E}$, with $\mathcal{X}=\mathbb{C}^{n\times n}$, $R=G$, $\mathsf{r}_{0}=\tilde{\mathsf{g}}$, $H=A$, $\mathsf{h}=\mathsf{a}$, $\mathsf{x}=I$, and
$\mathsf{z}={-}I\otimes B-\Theta\otimes\Xi+z\,I\otimes I.$
Then Assumption 4.1 holds (for $\gamma=1$ and modified constants $\upsilon,\delta>0$). Conditional on $B$ and the event $\mathcal{E}$, Theorem 4.2 shows over the randomness of $A$, uniformly in $i\neq j$ and $\mathbf{u},\mathbf{u}^{\prime}\in\mathbb{C}^{n}$ with $\|\mathbf{u}\|_{2},\|\mathbf{u}^{\prime}\|_{2}\leq\upsilon$,
$\begin{gathered}(n^{-1}\operatorname{Tr}\otimes I)[G]-(\tau\otimes I)[\tilde{\mathsf{g}}]\prec n^{-1},\\ G_{ii}-(\tau^{\mathcal{D}}\otimes I)[\tilde{\mathsf{g}}]_{ii}\prec n^{-1/2},\qquad G_{ij}\prec n^{-1/2},\\ (\mathbf{u}\otimes I)^{*}G(\mathbf{u}^{\prime}\otimes I)-(\mathbf{u}\otimes I)^{*}(\tau^{\mathcal{D}}\otimes I)[\tilde{\mathsf{g}}](\mathbf{u}^{\prime}\otimes I)\prec n^{-1/2}.\end{gathered}$ (109)
In light of the bound $G_{ii}^{-1}\prec 1$ from (90), the form $(\tau^{\mathcal{D}}\otimes I)[\tilde{\mathsf{g}}]_{ii}=(B+\theta_{i}\Xi-z-M_{B}(z))^{-1}$ from Lemma 4.3(a) where $M_{B}(z)=(\tau\otimes 1)[\tilde{\mathsf{g}}(z)]$, and the identity $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$, the second statement here shows also
$G_{ii}^{-1}-\big{(}(\tau^{\mathcal{D}}\otimes I)[\tilde{\mathsf{g}}]_{ii}\big{)}^{-1}=G_{ii}^{-1}-(B+\theta_{i}\Xi-z-M_{B}(z))\prec n^{-1/2},$
hence
$G_{ii}^{-1}-G_{jj}^{-1}=(\theta_{i}-\theta_{j})\Xi+O_{\prec}(n^{-1/2}).$
As $\mathbb{P}[\mathcal{E}]>1-n^{-D}$ for any fixed $D>0$ and all $n\geq n_{0}(D)$, these statements then also hold unconditionally. In particular, by Remark 3.3 this implies $\|G_{ij}\|_{\mathrm{op}}\prec n^{-1/2}$ and $\|G_{ii}^{-1}-G_{jj}^{-1}-(\theta_{i}-\theta_{j})\Xi\|_{\mathrm{op}}\prec n^{-1/2}$. The argument for $G_{\alpha\alpha}$ and $G_{\alpha\beta}$ is symmetric, so this shows (2224). The bounds (26) are an immediate consequence of (24).
To prove the remaining statements (21), (25), and (28), we apply Theorem 4.2 again to the second tensor factor, with $\mathcal{X}=\mathcal{A}$, $R=\tilde{\mathsf{g}}$, $\mathsf{r}_{0}=\mathsf{g}$, $H=B$, $\mathsf{h}=\mathsf{b}$, $\mathsf{x}=1_{\mathcal{A}}$, and
[…]
This gives, uniformly in $\alpha\neq\beta$ and $\mathbf{v},\mathbf{v}^{\prime}\in\mathbb{C}^{n}$ with $\|\mathbf{v}\|_{2},\|\mathbf{v}^{\prime}\|_{2}\leq\upsilon$,
$\begin{gathered}(1\otimes n^{-1}\operatorname{Tr})[\tilde{\mathsf{g}}]-(1\otimes\tau)[\mathsf{g}]\prec n^{-1},\\ \tilde{\mathsf{g}}_{\alpha\alpha}-(1\otimes\tau^{\mathcal{D}})[\mathsf{g}]_{\alpha\alpha}\prec n^{-1/2},\qquad\tilde{\mathsf{g}}_{\alpha\beta}\prec n^{-1/2},\\ (1\otimes\mathbf{v})^{*}\tilde{\mathsf{g}}(1\otimes\mathbf{v}^{\prime})-(1\otimes\mathbf{v})^{*}(1\otimes\tau^{\mathcal{D}})[\mathsf{g}](1\otimes\mathbf{v}^{\prime})\prec n^{-1/2}.\end{gathered}$ (110)
We remark that if $T:\mathcal{A}\to\mathcal{B}$ and $T^{\prime}:\mathcal{A}^{\prime}\to\mathcal{B}^{\prime}$ are two linear maps between vector spaces, then for any $x\in\mathcal{A}\otimes\mathcal{A}^{\prime}$,
$(T\otimes T^{\prime})[x]=(T\otimes 1)(1\otimes T^{\prime})[x]=(1\otimes T^{\prime})(T\otimes 1)[x].$ (111)
Thus we may combine the first statements of (109) and (110) to get as desired
$m(z)$ $=(n^{-1}\operatorname{Tr}\,\otimes\,n^{-1}\operatorname{Tr})[G]=n^{-1}\operatorname{Tr}\big{[}(n^{-1}\operatorname{Tr}\otimes I)[G]\big{]}$
$=(\tau\otimes n^{-1}\operatorname{Tr})[\tilde{\mathsf{g}}]+n^{-1}\operatorname{Tr}[\Delta_{a}]=\tau\big{[}(1\otimes n^{-1}\operatorname{Tr})[\tilde{\mathsf{g}}]\big{]}+n^{-1}\operatorname{Tr}[\Delta_{a}]$
$=\tau\otimes\tau[\mathsf{g}]+n^{-1}\operatorname{Tr}[\Delta_{a}]+\tau[\Delta_{b}]=m_{0}(z)+n^{-1}\operatorname{Tr}[\Delta_{a}]+\tau[\Delta_{b}]$
where $\Delta_{a}\in\mathbb{C}^{n\times n}$ and $\Delta_{b}\in\mathcal{A}$ are errors satisfying $\Delta_{a},\Delta_{b}\prec n^{-1}$. We have $|n^{-1}\operatorname{Tr}\Delta_{a}|\leq\|\Delta_{a}\|_{1}\prec n^{-1}$ and $|\tau[\Delta_{b}]|\leq\|\Delta_{b}\|_{1}\prec n^{-1}$ (Lemma D.3), showing (21). Similarly, applying (111) with $(T,T^{\prime})=(\mathbf{u}^{*}[\cdot]\mathbf{u}^{\prime},\mathbf{v}^{*}[\cdot]\mathbf{v}^{\prime})$ and $(T,T^{\prime})=(\tau^{\mathcal{D}},\mathbf{v}^{*}[\cdot]\mathbf{v}^{\prime})$, we may combine the last statements of (109) and (110) to get
$(\mathbf{u}\otimes\mathbf{v})^{*}G(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})$ $=\mathbf{v}^{*}\big{[}(\mathbf{u}\otimes I)^{*}G(\mathbf{u}^{\prime}\otimes I)\big{]}\mathbf{v}^{\prime}$
$=(\mathbf{u}\otimes\mathbf{v})^{*}(\tau^{\mathcal{D}}\otimes I)[\tilde{\mathsf{g}}](\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})+\mathbf{v}^{*}\Delta_{1}^{\prime}\mathbf{v}^{\prime}$
$=\mathbf{u}^{*}\tau^{\mathcal{D}}[(1\otimes\mathbf{v})^{*}\tilde{\mathsf{g}}(1\otimes\mathbf{v}^{\prime})]\mathbf{u}^{\prime}+\mathbf{v}^{*}\Delta_{1}^{\prime}\mathbf{v}^{\prime}$
$=(\mathbf{u}\otimes\mathbf{v})^{*}(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[\mathsf{g}](\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})+\mathbf{v}^{*}\Delta_{1}^{\prime}\mathbf{v}^{\prime}+\mathbf{u}^{*}\tau^{\mathcal{D}}[\Delta_{2}^{\prime}]\mathbf{u}^{\prime}$
where $\Delta_{1}^{\prime}\in\mathbb{C}^{n\times n}$ and $\Delta_{2}^{\prime}\in\mathcal{A}$ satisfy $\Delta_{1}^{\prime},\Delta_{2}^{\prime}\prec n^{-1/2}$. Since $\tau^{\mathcal{D}}$ is a $L^{p}$-contraction for each fixed $p$ (Lemma D.4), this implies also $\tau^{\mathcal{D}}[\Delta_{2}^{\prime}]\prec n^{-1/2}$. Then, since $\Delta_{1}^{\prime},\tau^{\mathcal{D}}(\Delta_{2}^{\prime})\in\mathbb{C}^{n\times n}$, by Remark 3.3 we have $\|\Delta_{1}^{\prime}\|_{\mathrm{op}},\|\tau^{\mathcal{D}}(\Delta_{2}^{\prime})\|_{\mathrm{op}}\prec n^{-1/2}$ and hence $\mathbf{v}^{*}\Delta_{1}^{\prime}\mathbf{v}^{\prime}\prec n^{-1/2}$ and $\mathbf{u}^{*}\tau^{\mathcal{D}}[\Delta_{2}^{\prime}]\mathbf{u}^{\prime}\prec n^{-1/2}$, showing (28). Finally, specializing this to $\mathbf{u}=\mathbf{u}^{\prime}=\mathbf{e}_{i}$ and $\mathbf{v}=\mathbf{v}^{\prime}=\mathbf{e}_{\alpha}$ shows (25). ∎
In the remainder of this section, we show the final estimate (27) of Theorem 2.4, $G_{ij,\alpha\beta}\prec n^{-1}$ when $i\neq j$ and $\alpha\neq\beta$. Recalling $\tilde{\mathsf{g}}$ from (108), define the matrices in $\mathbb{C}^{n\times n}$
$M_{B}=(\tau\otimes I)[\tilde{\mathsf{g}}],\qquad M_{i}=(B+\theta_{i}\Xi-zI-M_{B})^{-1}\text{ for }i=1,\ldots,n$ (112)
so that (by Lemma 4.3(b)) $M_{B}=n^{-1}\sum_{i}M_{i}$. Symmetrically, letting
$\check{\mathsf{g}}=(A\otimes 1_{\mathcal{A}}+I\otimes\mathsf{b}+\Theta\otimes\Xi-z\,I\otimes 1_{\mathcal{A}})^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{A},$
define
$M_{A}=(I\otimes\tau)[\check{\mathsf{g}}],\qquad M_{\alpha}=(A+\xi_{\alpha}\Theta-zI-M_{A})^{-1}\text{ for }\alpha=1,\ldots,n$ (113)
so that $M_{A}=n^{-1}\sum_{\alpha}M_{\alpha}$. We denote the commuting projection operators
$\mathbb{E}_{i}=\mathbb{E}[\cdot|A^{(i)},B],\qquad\mathcal{Q}_{i}=1-\mathbb{E}_{i},\qquad\mathbb{E}_{\alpha}=\mathbb{E}[\cdot|A,B^{[\alpha]}],\qquad\mathcal{Q}_{\alpha}=1-\mathbb{E}_{\alpha}$
and write as before $\mathcal{Q}_{S}=\prod_{i\in S}\mathcal{Q}_{i}$. With slight abuse of notation, we will use the distinction between Greek and Roman indices to distinguish between $M_{\alpha}$ and $M_{i}$, $\mathbb{E}_{\alpha}$ and $\mathbb{E}_{i}$, and $\mathcal{Q}_{\alpha}$ and $\mathcal{Q}_{i}$.
Lemma 5.1.
In the setting of Theorem 2.4, for any distinct $i,j,k$ and any $\alpha\neq\beta$, define
$\Delta^{k}=G_{kj}E_{\beta\alpha}G_{ik}\in\mathbb{C}^{n\times n}.$
Fix any $l\geq 1$. Then uniformly over $S\subseteq\{1,\ldots,n\}$ with $|S|\leq l$, distinct $i,j,k\notin S$, and $\alpha\neq\beta$,
$\mathcal{Q}_{S\cup\{k\}}[\Delta^{k}]\prec n^{-1-|S|/2}.$
Proof.
The proof is the same as that of Lemmas 4.12 and 4.13, using Lemma 3.6(c) to expand $G_{kj},G_{ik}$ in the indices of $S$ and applying the estimates $\|E_{\beta\alpha}\|_{\mathrm{op}}\leq 1$ and $G_{pq}^{(S^{\prime})}\prec n^{-1/2}$ for and $p,q\notin S^{\prime}$ with $p\neq q$. We omit the details for brevity. ∎
Proof of Theorem 2.4, (27).
Throughout this proof, $G_{ij},G_{\alpha\beta}$ etc. are all matrices in $\mathbb{C}^{n\times n}$, so the stochastic domination notation $\prec$ may be understood in the operator norm sense, c.f. Remark 3.3. We recall our convention of Roman indices $i,j,k,\ldots$ for the first tensor factor and Greek indices $\alpha,\beta,\gamma,\ldots$ for the second tensor factor. In the following, we use a superscript ()($\cdot$)( ⋅ ) to denote a minor on $A$ and []delimited-[][$\cdot$][ ⋅ ] to denote a minor on $B$.
Fix any indices $i\neq j$ and $\alpha\neq\beta$. Applying Lemma 3.6(b) to the second tensor factor,
$G_{\alpha\beta}={-}G_{\alpha\alpha}\sum_{\gamma}^{[\alpha]}b_{\alpha\gamma}G_{\gamma\beta}^{[\alpha]}.$
Recall (c.f. (107)) that
$\sum_{\gamma}^{[\alpha]}b_{\alpha\gamma}G_{\gamma\beta}^{[\alpha]}\prec n^{-1/2}.$ (114)
By (45) applied to the second tensor factor, we have $G_{\alpha\alpha}=M_{\alpha}+O_{\prec}(n^{-1/2})$, where $M_{\alpha}$ depends only on $A$ and not on $B$. Then by Lemma 3.2(c), also $\mathbb{E}_{\alpha}G_{\alpha\alpha}=M_{\alpha}+O_{\prec}(n^{-1/2})$, so
$G_{\alpha\alpha}-\mathbb{E}_{\alpha}G_{\alpha\alpha}\prec n^{-1/2}.$ (115)
Then applying both (114) and (115),
$G_{\alpha\beta}={-}(\mathbb{E}_{\alpha}G_{\alpha\alpha})\sum_{\gamma}^{[\alpha]}b_{\alpha\gamma}G_{\gamma\beta}^{[\alpha]}+O_{\prec}(n^{-1}),$
so
$G_{ij,\alpha\beta}={-}\sum_{\gamma}^{[\alpha]}b_{\alpha\gamma}\mathbf{e}_{i}^{*}(\mathbb{E}_{\alpha}G_{\alpha\alpha})G_{\gamma\beta}^{[\alpha]}\mathbf{e}_{j}+O_{\prec}(n^{-1})$ $\prec\frac{1}{\sqrt{n}}\left(\sum_{\gamma}^{[\alpha]}\left|\mathbf{e}_{i}^{*}(\mathbb{E}_{\alpha}G_{\alpha\alpha})G_{\gamma\beta}^{[\alpha]}\mathbf{e}_{j}\right|^{2}\right)^{1/2}+\frac{1}{n}$
$\prec\frac{1}{\sqrt{n}}\,\mathbb{E}_{\alpha}\left(\sum_{\gamma}^{[\alpha]}\left|\mathbf{e}_{i}^{*}G_{\alpha\alpha}G_{\gamma\beta}^{[\alpha]}\mathbf{e}_{j}\right|^{2}\right)^{1/2}+\frac{1}{n},$ (116)
the first inequality applying independence of $(b_{\alpha\gamma})_{\gamma=1}^{n}$ with $(\mathbb{E}_{\alpha}G_{\alpha\alpha})G_{\gamma\beta}^{[\alpha]}$ and the scalar version of Lemma 3.4(a), and the second line applying Jensen’s inequality and convexity of the $\ell_{2}$-norm.
Fixing $i\neq j$ and $\alpha\neq\beta$, define $\Delta\in\mathbb{C}^{n\times n}$ as the matrix with entries
$\Delta_{\gamma\nu}=\mathbf{e}_{i}^{*}G_{\alpha\nu}G_{\gamma\beta}\mathbf{e}_{j}=\sum_{k}G_{ik,\alpha\nu}G_{kj,\gamma\beta}=\sum_{k}\underbrace{\mathbf{e}_{\gamma}^{*}G_{kj}\mathbf{e}_{\beta}\mathbf{e}_{\alpha}^{*}G_{ik}\mathbf{e}_{\nu}}_{=\Delta_{\gamma\nu}^{k}}.$ (117)
We claim that (uniformly over $i\neq j$ and $\alpha\neq\beta$)
$\|\Delta\|_{\mathrm{op}}\prec n^{-1/2}.$ (118)
Then, for every $\gamma\neq\alpha$ (including $\gamma=\beta$), applying
$G_{\alpha\alpha}G_{\gamma\beta}^{[\alpha]}=G_{\alpha\alpha}G_{\gamma\beta}-G_{\alpha\alpha}G_{\gamma\alpha}G_{\alpha\alpha}^{-1}G_{\alpha\beta},\quad G_{\alpha\alpha},G_{\alpha\alpha}^{-1}\prec 1,\quad G_{\gamma\alpha},G_{\alpha\beta}\prec n^{-1/2},$
which follow from Lemma 3.6(c), (90), and (46) of Theorem 4.2, we have
$\mathbf{e}_{i}^{*}G_{\alpha\alpha}G_{\gamma\beta}^{[\alpha]}\mathbf{e}_{j}=\Delta_{\gamma\alpha}-\mathbf{e}_{i}^{*}G_{\alpha\alpha}G_{\gamma\alpha}G_{\alpha\alpha}^{-1}G_{\alpha\beta}\mathbf{e}_{j}=\Delta_{\gamma\alpha}+O_{\prec}(n^{-1})$
and hence by (118),
$\left(\sum_{\gamma}^{[\alpha]}\left|\mathbf{e}_{i}^{*}G_{\alpha\alpha}G_{\gamma\beta}^{[\alpha]}\mathbf{e}_{j}\right|^{2}\right)^{1/2}\leq\|\Delta\|_{\mathrm{op}}+n^{-1/2}\prec n^{-1/2}.$
Applying this in (116) yields the desired bound $G_{ij,\alpha\beta}\prec n^{-1}$.
It remains to show (118). For this, defining $\Delta=\sum_{k}\Delta^{k}$ where $\Delta^{k}=G_{kj}E_{\beta\alpha}G_{ik}$ as in (117), observe that
$\|\Delta^{k}\|_{\mathrm{op}}\leq\|G_{kj}\|_{\mathrm{op}}\cdot\|G_{ik}\|_{\mathrm{op}}\prec\begin{cases}n^{-1/2}&\text{ if }k\in\{i,j\}\\ n^{-1}&\text{ if }k\notin\{i,j\}\end{cases}$
Then
$\Delta=\sum_{k}^{(ij)}(\mathbb{E}_{k}+\mathcal{Q}_{k})[\Delta^{k}]+O_{\prec}(n^{-1/2}).$
By Lemma 3.5(a) applied with the estimates of Lemma 5.1, we have $\sum_{k}^{(ij)}\mathcal{Q}_{k}[\Delta^{k}]\prec n^{-1/2}$, so
$\Delta=\sum_{k}^{(ij)}\mathbb{E}_{k}[\Delta^{k}]+O_{\prec}(n^{-1/2}).$
For any $k\notin\{i,j\}$, we have analogously to (114) and (115) that
$G_{kk}-M_{k}\prec n^{-1/2},\qquad\sum_{\ell}^{(k)}a_{k\ell}G_{\ell j}^{(k)}\prec n^{-1/2},\qquad\sum_{\ell}^{(k)}G_{i\ell}^{(k)}a_{\ell k}\prec n^{-1/2}.$
Hence, applying the resolvent identities of Lemma 3.6(b),
$G_{ik}$ $={-}\sum_{\ell}^{(k)}G_{i\ell}^{(k)}a_{\ell k}G_{kk}={-}\sum_{\ell}^{(k)}G_{i\ell}^{(k)}a_{\ell k}M_{k}+O_{\prec}(n^{-1})$
$G_{kj}$ $={-}\sum_{\ell}^{(k)}G_{kk}a_{k\ell}G_{\ell j}^{(k)}={-}\sum_{\ell}^{(k)}M_{k}a_{k\ell}G_{\ell j}^{(k)}+O_{\prec}(n^{-1}).$
Applying this into the definition of $\Delta^{k}$,
$\sum_{k}^{(ij)}\mathbb{E}_{k}[\Delta^{k}]$ $=\sum_{k}^{(ij)}\mathbb{E}_{k}\Big{[}G_{kj}E_{\beta\alpha}G_{ik}\Big{]}$
$=\sum_{k}^{(ij)}\sum_{\ell,m}^{(k)}\mathbb{E}_{k}[a_{km}a_{\ell k}]M_{k}G_{mj}^{(k)}E_{\beta\alpha}G_{i\ell}^{(k)}M_{k}+O_{\prec}(n^{-1/2})$
$=\frac{1}{n}\sum_{k}^{(ij)}\sum_{\ell}^{(ijk)}M_{k}G_{\ell j}^{(k)}E_{\beta\alpha}G_{i\ell}^{(k)}M_{k}+O_{\prec}(n^{-1/2}),$
the last equality using $\mathbb{E}[a_{km}a_{\ell k}]=n^{-1}\mathbf{1}\{\ell=m\}$ and then absorbing the summands with $\ell\in\{i,j\}$ into the $O_{\prec}(n^{-1/2})$ error. Now applying $G_{\ell j}-G_{\ell j}^{(k)},G_{i\ell}-G_{i\ell}^{(k)}\prec n^{-1}$ (as follows from Lemma 3.6(c)), $M_{k}\prec 1$, and $G_{\ell j}^{(k)},G_{i\ell}^{(k)}\prec n^{-1/2}$ for all $\ell\notin\{i,j,k\}$, we get
$\sum_{k}^{(ij)}\mathbb{E}_{k}[\Delta^{k}]$ $=\frac{1}{n}\sum_{k}^{(ij)}\sum_{\ell}^{(ijk)}M_{k}G_{\ell j}E_{\beta\alpha}G_{i\ell}M_{k}+O_{\prec}(n^{-1/2})$
$=\frac{1}{n}\sum_{k}\sum_{\ell}M_{k}G_{\ell j}E_{\beta\alpha}G_{i\ell}M_{k}+O_{\prec}(n^{-1/2}),$
the second line introducing an additional $O_{\prec}(n^{-1/2})$ errors upon including the summands with $\ell\in\{i,j,k\}$, followed by $k\in\{i,j\}$. Observing that
$\frac{1}{n}\sum_{k,\ell=1}^{n}M_{k}G_{\ell j}E_{\beta\alpha}G_{i\ell}M_{k}=\frac{1}{n}\sum_{k=1}^{n}M_{k}\left[\sum_{\ell=1}^{n}G_{\ell j}E_{\beta\alpha}G_{i\ell}\right]M_{k}=\frac{1}{n}\sum_{k=1}^{n}M_{k}\Delta M_{k},$
this gives
$\Delta=\sum_{k}^{(ij)}\mathbb{E}_{k}[\Delta^{k}]+O_{\prec}(n^{-1/2})=\frac{1}{n}\sum_{k=1}^{n}M_{k}\Delta M_{k}+O_{\prec}(n^{-1/2}).$
Thus $\mathcal{L}_{1}(\Delta)\prec n^{-1/2}$ where $\mathcal{L}_{1}$ is the linear operator of Lemma 4.5 (in the current setting with $\mathsf{m}_{0}=M_{B}$ and $(\mathsf{r}_{0})_{ii}=M_{i}$). By the quantitative invertibility of $\mathcal{L}_{1}$ shown in Lemma 4.5, this implies the claim (118), completing the proof of Theorem 2.4. ∎
6. Analysis of least-squares problem
In this section, we prove Theorem 2.5, Corollary 2.6, and Proposition 2.7. Recall the optimization objective from Section 2,
$f(X)=\frac{1}{2}\|XA+BX\|_{F}^{2}+\frac{1}{2}\sum_{i,j=1}^{n}\xi_{i}\theta_{j}x_{ij}^{2}$
and its minimizer under a linear constraint,
$\widehat{X}=\operatorname*{arg\,min}_{X\in\mathbb{R}^{n\times n}}f(X)\text{ subject to }\frac{1}{n}\mathbf{v}^{*}X\mathbf{u}=1.$ (119)
Proof of Theorem 2.5.
Consider the following vectorization of (119) (where we use the convention of vectorization by column, i.e. $\mathbf{x}=\sum_{k}\mathbf{e}_{k}\otimes X\mathbf{e}_{k}$)
$f(\mathbf{x})$ $=\frac{1}{2}\mathbf{x}^{*}[(A\otimes I+I\otimes B)^{2}+(\Theta\otimes\Xi)]\mathbf{x},$
$\hat{\mathbf{x}}$ $=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n^{2}}}f(\mathbf{x})\text{ subject to }n^{-1}(\mathbf{u}\otimes\mathbf{v})^{*}\mathbf{x}=1.$
Denote
$P$ $=[(A\otimes I+I\otimes B)^{2}+(\Theta\otimes\Xi)]^{-1}\in\mathbb{R}^{n\times n}\otimes\mathbb{R}^{n\times n},$
$\mathsf{p}$ $=[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+(\Theta\otimes\Xi)]^{-1}\in\mathcal{A}\otimes\mathcal{A},$
$P_{0}$ $=(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[\mathsf{p}].$
Simple calculus yields the explicit forms for $\hat{\mathbf{x}}$ and $f(\hat{\mathbf{x}})$ as
$\hat{\mathbf{x}}=\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P(\mathbf{u}\otimes\mathbf{v})}\cdot n^{-1}P(\mathbf{u}\otimes\mathbf{v}),\qquad f(\hat{\mathbf{x}})=\frac{1}{2}\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P(\mathbf{u}\otimes\mathbf{v})}.$ (120)
Consider the linearization of $P$ given by
$\widetilde{P}$ $=\begin{bmatrix}-i\,\Theta\otimes\Xi&A\otimes I+I\otimes B\\ A\otimes I+I\otimes B&-i\,I\otimes I\end{bmatrix}^{-1}$
$=\left(\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\otimes(A\otimes I+I\otimes B)-i\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes\Theta\otimes\Xi-i\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\otimes I\otimes I\right)^{-1}.$
Denote $\mathbf{e}_{1}=(1,0)\in\mathbb{C}^{2}$. Then by Schur’s complement,
$-i\widetilde{P}=\begin{bmatrix}\Theta\otimes\Xi&i(A\otimes I+I\otimes B)\\ i(A\otimes I+I\otimes B)&I\otimes I\end{bmatrix}^{-1}=\begin{bmatrix}[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}&*\\ *&*\end{bmatrix}$ (121)
so that $P=(\mathbf{e}_{1}\otimes I\otimes I)^{*}[-i\widetilde{P}](\mathbf{e}_{1}\otimes I\otimes I)$. Defining also
$\tilde{\mathsf{p}}$ $=\left(\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\otimes(\mathsf{a}\otimes 1+1\otimes\mathsf{b})-i\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes\Theta\otimes\Xi-i\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\otimes 1\otimes 1\right)^{-1},$ (122)
$\widetilde{P}_{0}$ $=(I_{2\times 2}\otimes\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[\tilde{\mathsf{p}}],$
we have similarly $\mathsf{p}=(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}[-i\tilde{\mathsf{p}}](\mathbf{e}_{1}\otimes 1\otimes 1)$. Then, by (111) applied with $(T,T^{\prime})=(\mathbf{e}_{1}^{*}[\cdot]\mathbf{e}_{1},\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})$, we have also
$P_{0}=(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}[-i\widetilde{P}_{0}](\mathbf{e}_{1}\otimes 1\otimes 1).$
By (47) of Theorem 4.2, uniformly over $\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$ with $\|\mathbf{u}\|_{2},\|\mathbf{v}\|_{2},\|\mathbf{u}^{\prime}\|_{2},\|\mathbf{v}^{\prime}\|_{2}\leq\upsilon\sqrt{n}$,
$\frac{1}{n^{2}}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}P(\mathbf{u}\otimes\mathbf{v})$ $={-i}\left(\mathbf{e}_{1}\otimes\frac{\mathbf{u}^{\prime}}{\sqrt{n}}\otimes\frac{\mathbf{v}^{\prime}}{\sqrt{n}}\right)^{*}\widetilde{P}\left(\mathbf{e}_{1}\otimes\frac{\mathbf{u}}{\sqrt{n}}\otimes\frac{\mathbf{v}}{\sqrt{n}}\right)$
$={-i}\left(\mathbf{e}_{1}\otimes\frac{\mathbf{u}^{\prime}}{\sqrt{n}}\otimes\frac{\mathbf{v}^{\prime}}{\sqrt{n}}\right)^{*}\widetilde{P}_{0}\left(\mathbf{e}_{1}\otimes\frac{\mathbf{u}}{\sqrt{n}}\otimes\frac{\mathbf{v}}{\sqrt{n}}\right)+O_{\prec}(n^{-1/2})$
$=\frac{1}{n^{2}}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})+O_{\prec}(n^{-1/2}).$ (123)
Here, in the second line, we have applied (47) of Theorem 4.2 twice as in the proof of Theorem 2.4 in Section 5, first to the second tensor factor conditional on $B$ and the event =$\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$ with $\mathcal{X}=\mathbb{C}^{2\times 2}\otimes\mathbb{C}^{n\times n}$ (the product of first and third factors),
$H=A,\quad\mathsf{x}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\otimes I,\quad\mathsf{z}=i\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes\Theta\otimes\Xi+i\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\otimes I\otimes I-\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\otimes I\otimes B,$
and then to the third tensor factor with $\mathcal{X}=\mathbb{C}^{2\times 2}\otimes\mathcal{A}$ (the product of the first and second factors),
$H=B,\quad\mathsf{x}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\otimes 1,\quad\mathsf{z}=i\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes\Theta\otimes\Xi+i\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\otimes 1\otimes I-\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\otimes\mathsf{a}\otimes I.$
Since $\mathsf{p}$ is a positive operator, satisfying $\mathsf{p}\geq\|\mathsf{p}^{-1}\|_{\mathrm{op}}^{-1}(1_{\mathcal{A}}\otimes 1_{\mathcal{A}})$, it follows from positivity of $\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}}$ (Lemma D.2) and the bounds $\|\mathsf{a}\|_{\mathrm{op}}=\|\mathsf{b}\|_{\mathrm{op}}=2$ and $\|\Theta\|_{\mathrm{op}},\|\Xi\|_{\mathrm{op}}\leq\upsilon$ that
$P_{0}=\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}}[\mathsf{p}]\geq\|\mathsf{p}^{-1}\|_{\mathrm{op}}^{-1}\geq(16+\upsilon^{2})^{-1}.$
Therefore, for $\mathbf{u}=\mathbf{u}^{\prime}$ and $\mathbf{v}=\mathbf{v}^{\prime}$ satisfying $\|\mathbf{u}\|_{2},\|\mathbf{v}\|_{2}\geq\upsilon^{-1}\sqrt{n}$, we have the constant lower bound
$\frac{1}{n^{2}}(\mathbf{u}\otimes\mathbf{v})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})\geq(16+\upsilon^{2})^{-1}\frac{1}{n^{2}}\|\mathbf{u}\|_{2}^{2}\|\mathbf{v}\|_{2}^{2}\geq(16\upsilon^{4}+\upsilon^{6})^{-1},$
so we may apply the approximation (123) to both the numerator and denominator of (120) to get
$f(\hat{\mathbf{x}})$ $=\frac{1}{2}\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})}+O_{\prec}(n^{-1/2}),$
$\frac{1}{n}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}\hat{\mathbf{x}}$ $=\frac{n^{-2}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})}+O_{\prec}(n^{-1/2}).$
Finally, since $P_{0}\in\mathcal{D}\otimes\mathcal{D}\subset\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}$ is a diagonal matrix, writing $\mathsf{u}=\operatorname{diag}(\mathbf{u})$ etc. we have
$n^{-2}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})$ $=n^{-2}(\operatorname{Tr}\otimes\operatorname{Tr})[(\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})^{*}P_{0}(\mathsf{u}\otimes\mathsf{v})]$
$=(n^{-1}\operatorname{Tr}\otimes n^{-1}\operatorname{Tr})(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})^{*}\mathsf{p}(\mathsf{u}\otimes\mathsf{v})]$
$=(\tau\otimes\tau)[(\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})^{*}\mathsf{p}(\mathsf{u}\otimes\mathsf{v})].$
Applying this identity to both the numerators and denominators above concludes the proof. ∎
Proof of Corollary 2.6.
Under the given assumptions, for any fixed non-commutative polynomial $p$, we have
$\lim_{n\to\infty}(\tau\otimes\tau)\Big{(}p(\mathsf{a}\otimes 1,1\otimes\mathsf{b},\Theta\otimes\Xi,\mathsf{u}\otimes\mathsf{v},\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})\Big{)}=(\tau\otimes\tau)\Big{(}p(\mathsf{a}\otimes 1,1\otimes\mathsf{b},\theta\otimes\xi,\mathsf{U}\otimes\mathsf{U},\mathsf{U}^{\prime}\otimes\mathsf{U}^{\prime})\Big{)}$
where $\Theta,\mathsf{u},\mathsf{u}^{\prime},\Xi,\mathsf{v},\mathsf{v}^{\prime}\in\mathcal{D}\subset\mathcal{A}$ on the left are real diagonal matrices, and $\theta,\mathsf{U},\mathsf{U}^{\prime},\xi,\mathsf{V},\mathsf{V}^{\prime}\in\mathcal{A}$ on the right are commuting, self-adjoint limiting operators, free of $(\mathsf{a},\mathsf{b})$ and for which $(\theta,\mathsf{U},\mathsf{U}^{\prime})$ and $(\xi,\mathsf{V},\mathsf{V}^{\prime})$ have joint laws under $\tau$ given by $\mathcal{P}$ and $\mathcal{Q}$, respectively. The assumptions $\|\Theta\|_{\mathrm{op}},\|\Xi\|_{\mathrm{op}}\leq\upsilon$ and $\Theta,\Xi\geq\delta$ imply that the spectrum of $(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi$ is contained in $[\delta^{2},16+\upsilon^{2}]$. The inverse function $x\mapsto x^{-1}$ may be approximated uniformly by polynomials on this interval, so the above convergence implies
$\lim_{n\to\infty}(\tau\otimes\tau)[(\mathsf{u}^{\prime}\otimes\mathsf{v}^{\prime})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathsf{u}\otimes\mathsf{v})]$
$=(\tau\otimes\tau)[(\mathsf{U}^{\prime}\otimes\mathsf{V}^{\prime})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\theta\otimes\xi]^{-1}(\mathsf{U}\otimes\mathsf{V})]$
where this limit depends only on the joint laws $\mathcal{P},\mathcal{Q}$. Defining this limit quantity as $T^{\prime}(\mathcal{P},\mathcal{Q})$, and the analogous limit with $\mathsf{u}=\mathsf{u}^{\prime}$ and $\mathsf{v}=\mathsf{v}^{\prime}$ as $T(\mathcal{P},\mathcal{Q})$, the corollary then follows from Theorem 2.5. ∎
Proof of Proposition 2.7.
Following the above proof of Corollary 2.6, we write $\theta,\mathsf{U},\mathsf{U}^{\prime},\xi,\mathsf{V},\mathsf{V}^{\prime}\in\mathcal{A}$ for the limiting self-adjoint operators, which are free of $(\mathsf{a},\mathsf{b})$ and such that $(\theta,\mathsf{U},\mathsf{U}^{\prime})$ and $(\xi,\mathsf{V},\mathsf{V}^{\prime})$ have joint laws under $\tau$ given by $\mathcal{P}$ and $\mathcal{Q}$. Thus $\|\mathsf{U}\|_{\infty}\equiv\|\mathsf{U}\|_{\mathrm{op}}$ in the statement of Proposition 2.7. In this proof, we denote by $\mathcal{D}\subset\mathcal{A}$ the von Neumann subalgebra generated by $\theta,\mathsf{U},\mathsf{U}^{\prime},\xi,\mathsf{V},\mathsf{V}^{\prime}$.
We set $\eta=\min\{\sqrt{x_{a}x_{b}}:x_{a}\in\mathrm{supp}(\theta),\,x_{b}\in\mathrm{supp}(\xi)\}$ and define a limiting linearized operator $\tilde{\mathsf{p}}$ analogous to (122),
[…]
It is direct to check as in (121) that we have
$[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\theta\otimes\xi]^{-1}=(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}[-i\eta^{-1}\tilde{\mathsf{p}}](\mathbf{e}_{1}\otimes 1\otimes 1).$ (124)
Introducing the shorthands
$\mathsf{d}_{a}=\theta^{-1/2},\qquad\mathsf{d}_{b}=\xi^{-1/2},$
$\tilde{\mathsf{d}}=\begin{bmatrix}\eta^{1/2}\mathsf{d}_{a}\otimes\mathsf{d}_{b}&0\\ 0&\eta^{-1/2}1\otimes 1\end{bmatrix},\qquad\tilde{\mathsf{a}}=\begin{bmatrix}0&\mathsf{d}_{a}\mathsf{a}\otimes\mathsf{d}_{b}\\ 0&0\end{bmatrix},\qquad\tilde{\mathsf{b}}=\begin{bmatrix}0&\mathsf{d}_{a}\otimes\mathsf{d}_{b}\mathsf{b}\\ 0&0\end{bmatrix}$
and fixing a real argument $z>1$, we write this as
[…]
$=\tilde{\mathsf{d}}\Big{(}\underbrace{\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*}}_{=\tilde{\mathsf{q}}}+\,i(z-1)-iz\Big{)}^{-1}\tilde{\mathsf{d}}.$
Note that
$\|\mathsf{d}_{a}\mathsf{a}\otimes\mathsf{d}_{b}+\mathsf{d}_{a}\otimes\mathsf{d}_{b}\mathsf{b}\|_{\mathrm{op}}\leq\|\mathsf{d}_{a}\|_{\mathrm{op}}\|\mathsf{d}_{b}\|_{\mathrm{op}}(\|\mathsf{a}\|_{\mathrm{op}}+\|\mathsf{b}\|_{\mathrm{op}})\leq 4\eta^{-1},$
so $\tilde{\mathsf{q}}=\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*}$ is self-adjoint with spectrum contained in $[-4\eta^{-1},4\eta^{-1}]$. Hence
$\|\tilde{\mathsf{q}}+i(z-1)\|_{\mathrm{op}}=\max\{|\lambda+i(z-1)|:\lambda\in\mathrm{spec}(\tilde{\mathsf{q}})\}\leq\sqrt{(z-1)^{2}+16\eta^{-2}}.$ (125)
Applying iteratively $(\tilde{\mathsf{q}}+i(z-1)-iz)^{-1}={-}(iz)^{-1}+(iz)^{-1}[\tilde{\mathsf{q}}+i(z-1)](\tilde{\mathsf{q}}+i(z-1)-iz)^{-1}$ to write a series expansion of $(\tilde{\mathsf{q}}+i(z-1)-iz)^{-1}$, we have
$\tilde{\mathsf{p}}$ $={-}\sum_{k=0}^{M-1}(iz)^{-(k+1)}\tilde{\mathsf{d}}[\tilde{\mathsf{q}}+i(z-1)]^{k}\tilde{\mathsf{d}}+\mathsf{r}_{M}$
$={-}\sum_{k=0}^{M-1}(iz)^{-(k+1)}\sum_{m=0}^{k}\binom{k}{m}[i(z-1)]^{k-m}\tilde{\mathsf{d}}\tilde{\mathsf{q}}^{m}\tilde{\mathsf{d}}+\mathsf{r}_{M}$
$=i\sum_{m=0}^{M-1}\underbrace{\left(\frac{1}{i^{m}z(z-1)^{m}}\sum_{k=m}^{M-1}\binom{k}{m}\left(\frac{z-1}{z}\right)^{k}\right)}_{=C_{m}(z)}\tilde{\mathsf{d}}\tilde{\mathsf{q}}^{m}\tilde{\mathsf{d}}+\mathsf{r}_{M}$
with remainder
$\|\mathsf{r}_{M}\|_{\mathrm{op}}$ $=\left\|z^{-M}\tilde{\mathsf{d}}[\tilde{\mathsf{q}}+i(z-1)]^{M}(\tilde{\mathsf{q}}-i)^{-1}\tilde{\mathsf{d}}\right\|_{\mathrm{op}}\leq\eta^{-1}\left(\frac{\sqrt{(z-1)^{2}+16\eta^{-2}}}{z}\right)^{M}.$
Here, we have applied (125), $\|(\tilde{\mathsf{q}}-i)^{-1}\|_{\mathrm{op}}\leq 1$, and $\|\tilde{\mathsf{d}}\|_{\mathrm{op}}\leq\eta^{-1/2}$ as follows from its definition. Then, recalling the Schur complement identity (124) and applying also norm contractivity of $\tau\otimes\tau$ (Lemma D.2), we obtain
$(\tau\otimes\tau)[(\mathsf{U}^{\prime}\otimes\mathsf{V}^{\prime})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\theta\otimes\xi]^{-1}(\mathsf{U}\otimes\mathsf{V})]$
$=\sum_{m=0}^{M-1}C_{m}(z)(\tau\otimes\tau)\Big{[}(\mathbf{e}_{1}\otimes\mathsf{U}^{\prime}\otimes\mathsf{V}^{\prime})^{*}[\eta^{-1}\tilde{\mathsf{d}}\tilde{\mathsf{q}}^{m}\tilde{\mathsf{d}}](\mathbf{e}_{1}\otimes\mathsf{U}\otimes\mathsf{V})\Big{]}+r_{M}$ (126)
where $r_{M}$ is an error satisfying (35). We remark that the left side is real because all elements are self-adjoint, while each summand on the right is real for even $m$ and pure imaginary for odd $m$ by the definition of $C_{m}(z)$. Thus, taking real parts, this identity also holds with the summation restricted to even $m$.
We now analyze the summand for each even $m\in\{0,\ldots,M-1\}$. Let $\mathcal{W}_{m}$ be the set of words in the letters $(\mathsf{A},\mathsf{A}^{*},\mathsf{B},\mathsf{B}^{*})$ starting with a letter in $\{\mathsf{A},\mathsf{B}\}$ and alternating between a letter $\{\mathsf{A},\mathsf{B}\}$ and a letter $\{\mathsf{A}^{*},\mathsf{B}^{*}\}$, understood as non-commutative monic monomials of four variables. Then, applying the definitions of $\tilde{\mathsf{d}},\tilde{\mathsf{a}},\tilde{\mathsf{b}}$, we have
$(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}\eta^{-1}\tilde{\mathsf{d}}(\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*})^{m}\tilde{\mathsf{d}}(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}$
$=\sum_{w\in\mathcal{W}_{m}}\mathsf{d}_{a}w(\mathsf{d}_{a}\mathsf{a},\mathsf{a}\mathsf{d}_{a},\mathsf{d}_{a},\mathsf{d}_{a})\mathsf{d}_{a}\otimes\mathsf{d}_{b}w(\mathsf{d}_{b},\mathsf{d}_{b},\mathsf{d}_{b}\mathsf{b},\mathsf{b}\mathsf{d}_{b})\mathsf{d}_{b}\in\mathcal{A}\otimes\mathcal{A}.$
Thus
$(\tau\otimes\tau)\Big{[}(\mathbf{e}_{1}\otimes\mathsf{U}^{\prime}\otimes\mathsf{V}^{\prime})^{*}[\eta^{-1}\tilde{\mathsf{d}}\tilde{\mathsf{q}}^{m}\tilde{\mathsf{d}}](\mathbf{e}_{1}\otimes\mathsf{U}\otimes\mathsf{V})\Big{]}$
$=\sum_{w\in\mathcal{W}_{m}}\tau\Big{(}\mathsf{U}^{\prime}\mathsf{d}_{a}w(\mathsf{d}_{a}\mathsf{a},\mathsf{a}\mathsf{d}_{a},\mathsf{d}_{a},\mathsf{d}_{a})\mathsf{d}_{a}\mathsf{U}\Big{)}\cdot\tau\Big{(}\mathsf{V}^{\prime}\mathsf{d}_{b}w(\mathsf{d}_{b},\mathsf{d}_{b},\mathsf{d}_{b}\mathsf{b},\mathsf{b}\mathsf{d}_{b})\mathsf{d}_{b}\mathsf{V}\Big{)}$
$=\sum_{w\in\mathcal{W}_{m}}\tau\Big{(}\mathsf{d}_{a}\mathsf{U}\mathsf{U}^{\prime}\mathsf{d}_{a}w(\mathsf{d}_{a}\mathsf{a},\mathsf{a}\mathsf{d}_{a},\mathsf{d}_{a},\mathsf{d}_{a})\Big{)}\cdot\tau\Big{(}\mathsf{d}_{b}\mathsf{V}\mathsf{V}^{\prime}\mathsf{d}_{b}w(\mathsf{d}_{b},\mathsf{d}_{b},\mathsf{d}_{b}\mathsf{b},\mathsf{b}\mathsf{d}_{b})\Big{)}.$
For any $\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1}\in\mathcal{D}$, by the free moment-cumulant relations [51, Proposition 11.4], we have
$\tau(\mathsf{d}_{1}\mathsf{a}\mathsf{d}_{2}\mathsf{a}\ldots\mathsf{d}_{k}\mathsf{a}\mathsf{d}_{k+1})=\sum_{\pi\in\operatorname{NC}(2k+1)}\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})$
where $\operatorname{NC}(2k+1)$ is the set of non-crossing partitions of $(1,2,\ldots,2k+1)$ and $\kappa_{\pi}$ is the free cumulant associated to each $\pi\in\operatorname{NC}(2k+1)$. Since $\mathsf{a}$ is free of $\mathcal{D}$ and has $2^{\text{nd}}$ free cumulant equal to 1 and remaining free cumulants 0 [51, Example 11.21], we have $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})=0$ unless the blocks of $\pi$ containing $\mathsf{a}$ constitute a non-crossing pairing and are disjoint from those containing $(\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1})$. Let $\operatorname{NC}_{2}(k)$ denote the set of all non-crossing pairings of $\{2,4,6,\ldots,2k\}$ corresponding to the locations of the $\mathsf{a}$’s, and for each $\rho\in\operatorname{NC}_{2}(k)$, let $K(\rho)\in\operatorname{NC}(k+1)$ be its complement in $\{1,2,\ldots,2k+1\}$, i.e. the coarsest non-crossing partition of the remaining elements $\{1,3,5,\ldots,2k+1\}$ for which $\rho\cup K(\rho)$ forms a non-crossing partition of $\{1,2,\ldots,2k+1\}$. Then each $\pi\in\operatorname{NC}(2k+1)$ for which $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})\neq 0$ is the union of some $\rho\in\operatorname{NC}_{2}(k)$ and some $\bar{\rho}\leq K(\rho)$ that refines $K(\rho)$, so we have
$\tau(\mathsf{d}_{1}\mathsf{a}\mathsf{d}_{2}\mathsf{a}\ldots\mathsf{d}_{k}\mathsf{a}\mathsf{d}_{k+1})$ $=\sum_{\rho\in\operatorname{NC}_{2}(k)}\sum_{\bar{\rho}\in\operatorname{NC}(k+1):\bar{\rho}\leq K(\rho)}\kappa_{\rho}(\mathsf{a},\ldots,\mathsf{a})\kappa_{\bar{\rho}}(\mathsf{d}_{1},\mathsf{d}_{2},\ldots,\mathsf{d}_{k+1})$
$=\sum_{\rho\in\operatorname{NC}_{2}(k)}\sum_{\bar{\rho}\leq K(\rho)}\kappa_{\bar{\rho}}(\mathsf{d}_{1},\mathsf{d}_{2},\ldots,\mathsf{d}_{k+1})=\sum_{\rho\in\operatorname{NC}_{2}(k)}\prod_{S\in K(\rho)}\tau\left(\prod_{i\in S}\mathsf{d}_{i}\right),$
the last equality applying that $\{\bar{\rho}:\bar{\rho}\leq K(\rho)\}$ is a product of non-crossing partitions of the blocks of $K(\rho)$ and applying the free moment-cumulant relation over each block of $K(\rho)$.
Applying this above, corresponding to each word $w\in\mathcal{W}_{m}$, let $\operatorname{NC}_{w,2}(\mathsf{A})$ be the set of all non-crossing pairings of the letters $\{A,A^{*}\}$ of $w$ (not necessarily pairing $A$ with $A^{*}$), and let $\operatorname{NC}_{w,2}(\mathsf{B})$ be those of the letters $\{B,B^{*}\}$ of $w$. Then
$\tau\Big{(}\mathsf{d}_{a}\mathsf{U}\mathsf{U}^{\prime}\mathsf{d}_{a}w(\mathsf{d}_{a}\mathsf{a},\mathsf{a}\mathsf{d}_{a},\mathsf{d}_{a},\mathsf{d}_{a})\Big{)}\cdot\tau\Big{(}\mathsf{d}_{b}\mathsf{V}\mathsf{V}^{\prime}\mathsf{d}_{b}w(\mathsf{d}_{b},\mathsf{d}_{b},\mathsf{d}_{b}\mathsf{b},\mathsf{b}\mathsf{d}_{b})\Big{)}\\ =\sum_{\rho_{a}\in\operatorname{NC}_{w,2}(\mathsf{A})}\sum_{\rho_{b}\in\operatorname{NC}_{w,2}(\mathsf{B})}\mathrm{val}(\rho_{a},\rho_{b})$
where $\mathrm{val}(\cdot)$ is precisely the quantity defined in Section 2.2.1. Finally, we observe that summing first over words $w\in\mathcal{W}_{m}$ of $(A,A^{*},B,B^{*})$ and then over pairings $\rho_{a}\in\operatorname{NC}_{w,2}(\mathsf{A})$ and $\rho_{b}\in\operatorname{NC}_{w,2}(\mathsf{B})$ is equivalent to summing over all disjoint non-crossing pairings $(\rho_{a},\rho_{b})\in\operatorname{NC}_{2,2}(m)$, and then identifying $w$ as the word with letters $\{\mathsf{A},\mathsf{A}^{*}\}$ for elements of $\rho_{a}$ and $\{\mathsf{B},\mathsf{B}^{*}\}$ for elements of $\rho_{b}$ and that alternates between $\{A,B\}$ and $\{A^{*},B^{*}\}$. Thus
$(\tau\otimes\tau)\Big{[}(\mathbf{e}_{1}\otimes\mathsf{U}^{\prime}\otimes\mathsf{V}^{\prime})^{*}[\eta^{-1}\tilde{\mathsf{d}}\tilde{\mathsf{q}}^{m}\tilde{\mathsf{d}}](\mathbf{e}_{1}\otimes\mathsf{U}\otimes\mathsf{V})\Big{]}=\sum_{(\rho_{a},\rho_{b})\in\operatorname{NC}_{2,2}(m)}\mathrm{val}(\rho_{a},\rho_{b}).$
A direct calculation shows that this identity holds also for $m=0$, upon defining $\operatorname{NC}_{2,2}(0)$ to have the single pair (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) with its value defined as in (34). Applying this back to (126) and using $i^{m}=(-1)^{m/2}$ in $C_{m}(z)$ concludes the proof. ∎
Appendix A Contour integral representation in the case $\Theta=\Xi=\eta I$
The optimization problem (1) with $\Theta=\Xi=\eta I$ and $\eta\sim 1/(\log n)^{C}$ was studied previously in [30]. In this setting, its solution is given explicitly by
$\widehat{X}=\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P(\mathbf{u}\otimes\mathbf{v})}\cdot n^{-1}P(\mathbf{u}\otimes\mathbf{v})\quad\text{ where}\quad P=[(A\otimes I+I\otimes B)^{2}+\eta^{2}I\otimes I]^{-1}.$
Applying the linearization
$[(A\otimes I+I\otimes B)^{2}+\eta^{2}I\otimes I]^{-1}=\frac{1}{\eta}\Im[A\otimes I+I\otimes B-i\eta I\otimes I]^{-1},$
the analyses of [30] rested on a deterministic approximation for the resolvent
$R(z)=(Q-z\,I\otimes I)^{-1},\qquad Q=A\otimes I+I\otimes B$
at spectral scales $\Im z\sim 1/(\log n)^{C}$ decaying slowly with $n$.
This setting is special because the matrices $A\otimes I$ and $I\otimes B$ constituting $Q$ commute. In this setting, writing the spectral decompositions $A=\sum_{j=1}^{n}\lambda_{j}\mathbf{u}_{j}\mathbf{u}_{j}^{*}$ and $B=\sum_{k=1}^{n}\mu_{k}\mathbf{v}_{k}\mathbf{v}_{k}^{*}$, the spectral decomposition of $Q$ is explicit and given by
$Q=\sum_{j,k=1}^{n}(\lambda_{j}+\mu_{k})(\mathbf{u}_{j}\otimes\mathbf{v}_{k})(\mathbf{u}_{j}\otimes\mathbf{v}_{k})^{*}.$
In particular, the limit eigenvalue distribution of $Q$ is the (classical) convolution of the semicircle law with itself. Furthermore, defining a contour $\Gamma$ enclosing $\{\lambda_{1},\ldots,\lambda_{n}\}$ and such that $|\Im w|\leq(\Im z)/2$ for all $w\in\Gamma$, by the Cauchy integral formula applied to $f_{k}(w)=(w+\mu_{k}-z)^{-1}$ (which is analytic inside $\Gamma$) the resolvent of $Q$ has the explicit contour integral representation
$R(z)$ $=\sum_{j,k=1}^{n}\frac{1}{\lambda_{j}+\mu_{k}-z}\mathbf{u}_{j}\mathbf{u}_{j}^{*}\otimes\mathbf{v}_{k}\mathbf{v}_{k}^{*}$
[…]
where $R_{A}(z)=(A-zI)^{-1}$ and $R_{B}(z)=(B-zI)^{-1}$. From this representation, resolvent estimates of the form in Theorem 2.4 may be deduced from known local laws for the resolvents of the Wigner matrices $A$ and $B$, see e.g. [27, 24], and analysis of this commutative case suggests that the estimates in Theorem 2.4 are also optimal (for fixed $z\in\mathbb{C}^{+}$).
We emphasize that this type of analysis and contour integral representation does not extend to models of the form $Q=A\otimes I+I\otimes B+\Theta\otimes\Xi$ when $\Theta\otimes\Xi$ does not commute with either $A\otimes I$ or $I\otimes B$, which is the focus of our current work.
Appendix B Operator concentration inequalities
In this section, we prove Lemma 3.4 using the following version of the non-commutative Rosenthal inequality of [38]. We recall that $\mathcal{X}$ is a von Neumann algebra with faithful, normal, tracial state $\phi$, and $L^{p}(\mathcal{X})$ is its associated non-commutative $L^{p}$ space (c.f. Appendix D) with norm $\|\mathsf{x}\|_{p}=\phi((\mathsf{x}^{*}\mathsf{x})^{p/2})^{1/p}$.
Lemma B.1 ([38], Theorem 2.1).
Let $\mathcal{Y}\subset\mathcal{X}$ be a von Neumann subalgebra with $\phi$-invariant conditional expectation $\phi^{\mathcal{Y}}:\mathcal{X}\to\mathcal{Y}$. Suppose $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}\in L^{p}(\mathcal{X})$ satisfy $\phi^{\mathcal{Y}}(\mathsf{x}_{i})=0$, and are independent over $\mathcal{Y}$ in the sense that for each $i$, each $\mathsf{x}$ in the von Neumann subalgebra generated by $\mathsf{x}_{i}$, and each $\mathsf{x}^{\prime}$ in the von Neumann subalgebra generated by $\{\mathsf{x}_{j}:j\neq i\}$, we have $\phi^{\mathcal{Y}}(\mathsf{x}\mathsf{x}^{\prime})=\phi^{\mathcal{Y}}(\mathsf{x})\phi^{\mathcal{Y}}(\mathsf{x}^{\prime})$.
Then for any $p\in[2,\infty)$ and a universal constant $C>0$,
$\bigg{\|}\sum_{i=1}^{n}\mathsf{x}_{i}\bigg{\|}_{p}\leq Cp\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\phi^{\mathcal{Y}}(\mathsf{x}_{i}\mathsf{x}_{i}^{*})\bigg{)}^{1/2}\bigg{\|}_{p},\;\bigg{\|}\bigg{(}\sum_{i=1}^{n}\phi^{\mathcal{Y}}(\mathsf{x}_{i}^{*}\mathsf{x}_{i})\bigg{)}^{1/2}\bigg{\|}_{p},\;\bigg{(}\sum_{i=1}^{n}\|\mathsf{x}_{i}\|_{p}^{p}\bigg{)}^{1/p}\bigg{\}}.$
Lemma B.2 (Decoupling).
[…]
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{i\neq j}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg{\|}^{p}\bigg{]}\leq C^{p+1}\mathbb{E}\bigg{[}\bigg{\|}\sum_{i\neq j}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg{\|}^{p}\bigg{]}$
Proof.
By [14, Theorem 1], for a universal constant $C>0$,
[…]
Proof of Lemma 3.4.
Throughout, $C_{p},C_{p}^{\prime},C_{p}^{\prime\prime}$ denote $p$-dependent constants that may change from instance to instance.
For (a), fix any $p\geq 2$. We apply Lemma B.1 in the setting of [38, Example 1.3]: Let $L^{\infty}(\Omega)$ be the von Neumann algebra of bounded scalar random variables over the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$, and consider $\mathcal{M}=L^{\infty}(\Omega)\otimes\mathcal{X}$ equipped with the state $\mathbb{E}\circ\phi$. Then $\alpha_{i}\mathsf{x}_{i}\in L^{p}(\mathcal{M},\mathbb{E}\circ\phi)$, $\mathbb{E}:\mathcal{M}\to\mathcal{X}$ coincides with the conditional expectation onto the subalgebra $\mathcal{X}\subset\mathcal{M}$, and $\{\alpha_{i}\mathsf{x}_{i}\}_{i=1}^{n}$ are independent over $\mathcal{X}$ in the sense of Lemma B.1, so Lemma B.1 shows
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{i=1}^{n}\alpha_{i}\mathsf{x}_{i}\bigg{\|}_{p}^{p}\bigg{]}\leq C_{p}\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathbb{E}[|\alpha_{i}|^{2}]\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\;\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathbb{E}[|\alpha_{i}|^{2}]\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\;\sum_{i=1}^{n}\mathbb{E}[|\alpha_{i}|^{p}]\|\mathsf{x}_{i}\|_{p}^{p}\bigg{\}}.$
Applying the bounds $\mathbb{E}[|\alpha_{i}|^{2}],\mathbb{E}[|\alpha_{i}|^{p}]\leq C_{p}$, operator monotonicity of the square-root $0\leq\mathsf{x}\leq\mathsf{y}\Rightarrow\mathsf{x}^{1/2}\leq\mathsf{y}^{1/2}$, and monotonicity of the $L^{p}$-norm on the positive cone (Lemma D.3), this implies
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{i=1}^{n}\alpha_{i}\mathsf{x}_{i}\bigg{\|}_{p}^{p}\bigg{]}\leq C_{p}^{\prime}\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\;\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\;\sum_{i=1}^{n}\|\mathsf{x}_{i}\|_{p}^{p}\bigg{\}}.$ (127)
Here, for $p\geq 2$, the third term is bounded by the second by the following argument (see also [38, Eq. (2.4)]): Consider $\hat{\mathsf{x}}=\sum_{i=1}^{n}E_{i1}\otimes\mathsf{x}_{i}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ equipped with the trace $n^{-1}\operatorname{Tr}\otimes\phi$, and the linear map $T:\mathbb{C}^{n\times n}\otimes\mathcal{X}\to\mathbb{C}^{n\times n}\otimes\mathcal{X}$ defined by $T(E_{ij}\otimes\mathsf{x})=E_{i,j+i-1}\otimes\mathsf{x}$ where $j+i-1$ is interpreted modulo $n$. Thus $\hat{\mathsf{x}}$ has $\mathcal{X}$-valued entries $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}$ along the first column, and $T(\hat{\mathsf{x}})$ has these entries instead along the main diagonal. We have
$\|\hat{\mathsf{x}}\|_{p}^{p}$ $=\|(\hat{\mathsf{x}}^{*}\hat{\mathsf{x}})^{1/2}\|_{p}^{p}=\frac{1}{n}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},$
$\|T(\hat{\mathsf{x}})\|_{p}^{p}$ $=\|(T(\hat{\mathsf{x}})^{*}T(\hat{\mathsf{x}}))^{1/2}\|_{p}^{p}=\frac{1}{n}\sum_{i=1}^{n}\|(\mathsf{x}_{i}^{*}\mathsf{x}_{i})^{1/2}\|_{p}^{p}=\frac{1}{n}\sum_{i=1}^{n}\|\mathsf{x}_{i}\|_{p}^{p}.$
For $p=2$, this shows $\|T(\hat{\mathsf{x}})\|_{2}=\|\hat{\mathsf{x}}\|_{2}$. For $p=\infty$, we have
$\|T(\hat{\mathsf{x}})\|_{\mathrm{op}}=\max_{i}\|\mathsf{x}_{i}\|_{\mathrm{op}}=\max_{i}\|(\mathsf{x}_{i}^{*}\mathsf{x}_{i})^{1/2}\|_{\mathrm{op}}\leq\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{op}}=\|\hat{\mathsf{x}}\|_{\mathrm{op}}$
by operator monotonicity of the square-root and monotonicity of the operator norm on the positive cone. Then $\|T(\hat{\mathsf{x}})\|_{p}\leq\|\hat{\mathsf{x}}\|_{p}$ for all $p\in[2,\infty]$ by the Riesz-Thorin interpolation (Lemma D.5). Thus, the third term of (127) is at most the first, yielding the claim of part (a).
Part (b) follows from a two-fold application of part (a): Let $\mathsf{x}_{i}=\sum_{j}\beta_{j}\mathsf{y}_{ij}$, so $\sum_{i}\alpha_{i}\mathsf{x}_{i}=\sum_{i,j}\alpha_{i}\beta_{j}\mathsf{y}_{ij}$. Then, applying part (a) conditional on $(\beta_{i})_{i=1}^{n}$, we have
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{i,j=1}^{n}\alpha_{i}\beta_{j}\mathsf{y}_{ij}\bigg{\|}_{p}^{p}\bigg{]}\leq C_{p}\max\bigg{\{}\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]},\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}\bigg{\}}$ (128)
To apply part (a) again on these errors, define $\hat{\mathsf{y}}_{ij}=E_{1i}\otimes\mathsf{y}_{ij}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ equipped with the trace $n^{-1}\operatorname{Tr}\otimes\phi$. It follows that $\hat{\mathsf{y}}_{ij}\hat{\mathsf{y}}_{kl}^{*}=\mathbf{1}_{i=k}E_{11}\otimes\mathsf{y}_{ij}\mathsf{y}_{il}^{*}$, so
$E_{11}\otimes\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}=\sum_{j,l=1}^{n}\beta_{j}\bar{\beta}_{l}\sum_{i=1}^{n}E_{11}\otimes\mathsf{y}_{ij}\mathsf{y}_{il}^{*}=\bigg{(}\sum_{j=1}^{n}\beta_{j}\sum_{i=1}^{n}\hat{\mathsf{y}}_{ij}\bigg{)}\bigg{(}\sum_{l=1}^{n}\beta_{l}\sum_{k=1}^{n}\hat{\mathsf{y}}_{kl}\bigg{)}^{*}.$
Then
$\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}=n\bigg{\|}\bigg{(}E_{11}\otimes\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}=n\bigg{\|}\sum_{j=1}^{n}\beta_{j}\sum_{i=1}^{n}\hat{\mathsf{y}}_{ij}\bigg{\|}_{p}^{p}.$
Applying part (a) with $\hat{\mathsf{x}}_{j}=\sum_{i}\hat{\mathsf{y}}_{ij}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ in place of $\mathsf{x}_{j}\in\mathcal{X}$, this is bounded as
$\mathbb{E}\bigg{[}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{]}$ $\leq C_{p}n\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{j=1}^{n}\hat{\mathsf{x}}_{j}\hat{\mathsf{x}}_{j}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\bigg{\|}\bigg{(}\sum_{j=1}^{n}\hat{\mathsf{x}}_{j}^{*}\hat{\mathsf{x}}_{j}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{\}}$
$=C_{p}n\max\bigg{\{}\bigg{\|}\bigg{(}E_{11}\otimes\sum_{i,j}\mathsf{y}_{ij}\mathsf{y}_{ij}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\bigg{\|}\bigg{(}\sum_{i,j,k}E_{ik}\otimes\mathsf{y}_{ij}^{*}\mathsf{y}_{kj}\bigg{)}^{1/2}\bigg{\|}_{p}^{p}\bigg{\}}.$
The first term is $n^{-1}\|(\sum_{i,j}\mathsf{y}_{ij}\mathsf{y}_{ij}^{*})^{1/2}\|_{p}^{p}$. For the second term, we identify $\sum_{i,j,k}E_{ik}\otimes\mathsf{y}_{ij}^{*}\mathsf{y}_{kj}={\mathsf{Y}^{\mathsf{t}}}^{*}{\mathsf{Y}^{\mathsf{t}}}$ where $\mathsf{Y}^{\mathsf{t}}=\sum_{i,j}E_{ji}\otimes\mathsf{y}_{ij}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$, and we apply $\|({\mathsf{Y}^{\mathsf{t}}}^{*}{\mathsf{Y}^{\mathsf{t}}})^{1/2}\|_{p}^{p}=\|\mathsf{Y}^{\mathsf{t}}\|_{p}^{p}$. We have analogously $\mathbb{E}\|(\sum_{i}\mathsf{x}_{i}^{*}\mathsf{x}_{i})^{1/2}\|_{p}^{p}\leq C_{p}n\max(n^{-1}\|(\sum_{i,j}\mathsf{y}_{ij}^{*}\mathsf{y}_{ij})^{1/2}\|_{p}^{p},\|\mathsf{Y}\|_{p}^{p})$, and combining these gives part (b).
Finally, part (c) follows from part (b) and the decoupling result of Lemma B.2: Since $\mathsf{y}_{ii}=0$ for all $i=1,\ldots,n$,
$\mathbb{E}\bigg{[}\bigg{\|}\sum_{i\neq j}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg{\|}_{p}^{p}\bigg{]}\leq C_{p}\mathbb{E}\bigg{[}\bigg{\|}\sum_{i\neq j}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg{\|}_{p}^{p}\bigg{]}=C_{p}\mathbb{E}\bigg{[}\bigg{\|}\sum_{i,j=1}^{n}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg{\|}_{p}^{p}\bigg{]}$
$\leq C_{p}^{\prime}\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{i\neq j}\mathsf{y}_{ij}\mathsf{y}_{ij}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\bigg{\|}\bigg{(}\sum_{i\neq j}\mathsf{y}_{ij}^{*}\mathsf{y}_{ij}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},n\|\mathsf{Y}\|_{p}^{p},n\|\mathsf{Y}^{\mathsf{t}}\|_{p}^{p}\bigg{\}}.$
Appendix C Fluctuation averaging
In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [29, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [29].
Lemma C.1 ([29], Lemma A.3).
Fix $l\geq 1$. For each $a=1,\ldots,l$, let $B^{a}=(B_{ij}^{a})_{i,j=1}^{n}\in\mathbb{R}^{n\times n}$ satisfy
$B_{ij}^{a}\geq 0,\quad\quad B_{ii}^{a}=0,\quad\quad\|B^{a}\|_{F}\leq 1\qquad\text{ for all }i,j=1,\ldots,n.$
For $(\mathbf{i},\mathbf{j})=(i_{1},...,i_{l},j_{1},...,j_{l})\in\{1,\ldots,n\}^{2l}$, let $T(\mathbf{i},\mathbf{j})$ be the set of elements of $\{1,\ldots,n\}$ that appear exactly once in $(\mathbf{i},\mathbf{j})$. Then for some constant $C_{l}>0$ and all $t\in\{0,\ldots,2l\}$,
$\sum_{(\mathbf{i},\mathbf{j})\in\{1,\ldots,n\}^{2l}:|T(\mathbf{i},\mathbf{j})|=t}\;\prod_{a=1}^{l}B_{i_{a}j_{a}}^{a}\leq C_{l}n^{t/2}$
Proof of Lemma 3.5.
For part (a), fix any $\epsilon,D>0$ and $p\in[1,\infty)$. Pick an even integer $l>p$ such that $\epsilon(l-1)>D+1$. Then by monotonicity of $\|\cdot\|_{p}$ in $p$ (Lemma D.3),
$\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}^{l}\leq\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{l}^{l}=\mathbb{E}\phi\left(\left[\left(\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right)\left(\sum_{i=1}^{n}\bar{u}_{i}\mathsf{x}_{i}^{*}\right)\right]^{l/2}\right)=\sum_{\mathbf{i}}u_{\mathbf{i}}\mathbb{E}\phi\left(\mathsf{x}_{\mathbf{i}}\right)$ (129)
where we denote
$\mathbf{i}=(i_{1},i_{2},...,i_{l}),\quad\sum_{\mathbf{i}}=\sum_{i_{1},...,i_{l}=1}^{n},\quad u_{\mathbf{i}}=u_{i_{1}}\bar{u}_{i_{2}}\cdots u_{i_{l-1}}\bar{u}_{i_{l}},\quad\mathsf{x}_{\mathbf{i}}=\mathsf{x}_{i_{1}}\mathsf{x}_{i_{2}}^{*}\cdots\mathsf{x}_{i_{l-1}}\mathsf{x}_{i_{l}}^{*}=\prod_{a=1}^{l}\widetilde{\mathsf{x}}_{i_{a}}$
and write $\widetilde{\mathsf{x}}_{i_{a}}=\mathsf{x}_{i_{a}}$ if $a$ is odd and $\widetilde{\mathsf{x}}_{i_{a}}=\mathsf{x}_{i_{a}}^{*}$ if $a$ is even. Here and below, the product over $a=1,\ldots,l$ is non-commutative, and should be understood in the ordered sense.
For fixed $\mathbf{i}$, let $T\equiv T(\mathbf{i})\subseteq\{1,\ldots,n\}$ be the indices that appear exactly once in $\mathbf{i}=(i_{1},\ldots,i_{l})$. Using that $\{\mathbb{E}_{j},\mathcal{Q}_{j}\}$ commute, we have the identity
$\mathsf{x}=\left(\prod_{j\in T}(\mathbb{E}_{j}+\mathcal{Q}_{j})\right)[\mathsf{x}]=\sum_{S\subseteq T}\mathbb{E}_{T\setminus S}\mathcal{Q}_{S}[\mathsf{x}].$
(In the case $T=\emptyset$, this is the trivial identity $\mathsf{x}=\mathsf{x}$.) Applying this to each $\widetilde{\mathsf{x}}_{i_{a}}$,
$\mathsf{x}_{\mathbf{i}}=\sum_{S_{1},\ldots,S_{l}\subseteq T}\mathsf{x}(S_{1},\ldots,S_{l}),\quad\quad\mathsf{x}(S_{1},\ldots,S_{l})=\prod_{a=1}^{l}\mathbb{E}_{T\setminus S_{a}}\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a}}].$
By Hölder’s inequality (Lemma D.3) and (37),
$|\phi(\mathsf{x}(S_{1},\ldots,S_{l}))|\leq\prod_{a=1}^{l}\|\mathbb{E}_{T\setminus S_{a}}\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]\|_{l}\leq\prod_{a=1}^{l}\mathbb{E}_{T\setminus S_{a}}\|\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]\|_{l}.$
Let us write $S\setminus i$ for $S$ removing $i$ if $i\in S$, or for $S$ itself if $i\notin S$. Since $\mathbb{E}_{i_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]=0$ by assumption, we have $\mathcal{Q}_{i_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]=\widetilde{\mathsf{x}}_{i_{a}}$, so $\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]=\mathcal{Q}_{S_{a}\setminus i_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]$. Then the given condition (39) and Lemma 3.2(c) (in the setting of scalar random variables) imply $\mathbb{E}_{T\setminus S_{a}}\|\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]\|_{l}\prec n^{-\alpha-\beta|S_{a}\setminus i_{a}|}$, for each fixed $l$ uniformly over $i_{a}\in\{1,\ldots,n\}$ and over $S_{a}\subseteq T\subseteq\{1,\ldots,n\}$ with $|T|\leq l$. Thus, multiplying across $a=1,\ldots,l$ and taking the full expectation, for all $n\geq n_{0}(l,\epsilon)$ we have
$|\mathbb{E}\phi(\mathsf{x}(S_{1},\ldots,S_{l}))|\leq\mathbb{E}|\phi(\mathsf{x}(S_{1},\ldots,S_{l}))|\leq n^{-\alpha l-\beta\sum_{a=1}^{l}|S_{a}\setminus i_{a}|+\epsilon}.$ (130)
Now consider any $a\in\{1,\ldots,l\}$ such that $i_{a}\in T$. Observe that
  • $\mathsf{x}(S_{1},\ldots,S_{l})=0$ unless $i_{a}\in S_{a}$. Indeed, if instead $i_{a}\in T\setminus S_{a}$, then the assumption $\mathbb{E}_{i_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]=0$ implies $\mathbb{E}_{T\setminus S_{a}}\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]=0$.
  • $\mathbb{E}\phi(\mathsf{x}(S_{1},\ldots,S_{l}))=0$ unless also $i_{a}\in S_{b}$ for some $b\neq a$. Indeed, if instead $i_{a}\in T\setminus S_{b}$ for every $b\neq a$, then $\mathbb{E}_{T\setminus S_{b}}\mathcal{Q}_{S_{b}}[\widetilde{\mathsf{x}}_{i_{b}}]=\mathbb{E}_{i_{a}}\mathbb{E}_{T\setminus(S_{b}\cup\{i_{a}\})}\mathcal{Q}_{S_{b}}[\widetilde{\mathsf{x}}_{i_{b}}]$ is $\mathscr{G}_{i_{a}}$-measurable, hence
    $\mathbb{E}_{i_{a}}\mathsf{x}(S_{1},\ldots,S_{l})=\prod_{b<a}\mathbb{E}_{T\setminus S_{b}}\mathcal{Q}_{S_{b}}[\widetilde{\mathsf{x}}_{i_{b}}]\cdot\underbrace{\mathbb{E}_{i_{a}}\mathbb{E}_{T\setminus S_{a}}\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a}}]}_{=0}\cdot\prod_{b>a}\mathbb{E}_{T\setminus S_{b}}\mathcal{Q}_{S_{b}}[\widetilde{\mathsf{x}}_{i_{b}}]=0$
    where the middle term is 0 because $\mathbb{E}_{i_{a}}\mathcal{Q}_{S_{a}}=0$ for $i_{a}\in S_{a}$. Then by linearity of $\phi$, $\mathbb{E}\phi(\mathsf{x}(S_{1},\ldots,S_{l}))=\phi(\mathbb{E}\mathsf{x}(S_{1},\ldots,S_{l}))=0$.
Thus, if $\mathbb{E}\phi(\mathsf{x}(S_{1},\ldots,S_{l}))\neq 0$, then each $i_{a}\in T$ must appear in both $S_{a}$ and some set $S_{b}$ for $b\neq a$, where $i_{a}\neq i_{b}$ because $i_{a}$ appears only once in $\mathbf{i}=(i_{1},\ldots,i_{l})$ by definition of $T$. So $\sum_{a=1}^{l}|S_{a}\setminus i_{a}|\geq|T|$. Applying this to (130) and then to (129), since for each fixed $\mathbf{i}$ and $T\equiv T(\mathbf{i})$ the number of choices of subsets $S_{1},\ldots,S_{l}\subseteq T$ is at most a constant $C_{l}$,
$\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}^{l}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{\mathbf{i}}|u_{\mathbf{i}}|n^{-\beta|\mathcal{T(\mathbf{i})}|}\leq C_{l}n^{-\alpha l+\epsilon}\|\mathbf{u}\|_{\infty}^{l}\sum_{t=1}^{l}n^{-\beta t}\cdot|\{\mathbf{i}:T(\mathbf{i})=t\}|.$ (131)
By scale invariance of the statement of the lemma, let us assume without loss of generality that $\|\mathbf{u}\|_{\infty}=n^{-1}$. Applying this and the bound |$|\{\mathbf{i}:T(\mathbf{i})=t\}|\leq C_{l}^{\prime}n^{t+(l-t)/2}$ for a constant $C_{l}^{\prime}>0$, we get (for different constants $C_{l},C_{l}^{\prime}>0$)
$\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}^{l}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{t=1}^{l}(n^{-\beta})^{t}(n^{-1/2})^{l-t}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{t=1}^{l}(n^{-\beta}+n^{-1/2})^{l}\leq C_{l}^{\prime}n^{-(\alpha+\beta^{\prime})l+\epsilon}$
where $\beta^{\prime}=\min\{1/2,\beta\}$. Then by Markov’s inequality under our choices $l>p$ and $\epsilon(l-1)>D+1$,
$\mathbb{P}\left(\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}\geq n^{-\alpha-\beta^{\prime}+\epsilon}\right)\leq n^{(\alpha+\beta^{\prime}-\epsilon)l}\,\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}^{l}\leq C_{l}^{\prime}n^{-\epsilon(l-1)}<n^{-D}$
for all $n\geq n_{0}(l,\epsilon,D)$. Here $l$ depends only on $(p,\epsilon,D)$, showing $\sum_{i}u_{i}\mathsf{x}_{i}\prec n^{-\alpha-\beta^{\prime}}$ as claimed in part (a).
The argument for part (b) is the same, until the analysis of (131) where we apply a different counting argument: By scale invariance, we may consider without loss of generality $\mathbf{u}\in\mathbb{C}^{n}$ with $\|\mathbf{u}\|_{2}=1$. Under the given condition (40), specializing to $\beta=1/2$, the first inequality of (131) becomes
$\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}^{l}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{\mathbf{i}}|u_{\mathbf{i}}|n^{-|\mathcal{T(\mathbf{i})}|/2}.$
Now let $\pi(\mathbf{i})$ be the partition of $\{1,\ldots,l\}$ induced by coincidence of indices in $\mathbf{i}$, i.e. $a,b$ belong to the same block of $\pi(\mathbf{i})$ if and only if $i_{a}=i_{b}$. Then $|T(\mathbf{i})|\equiv|T(\pi(\mathbf{i}))|$ is the number of singleton blocks of $\pi(\mathbf{i})$, depending on $\mathbf{i}$ only via $\pi(\mathbf{i})$, so we may write the above as
$\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}^{l}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{\pi}n^{-|T(\pi)|/2}\sum_{\mathbf{i}:\pi(\mathbf{i})=\pi}|u_{\mathbf{i}}|.$ (132)
Here $\sum_{\mathbf{i}:\pi(\mathbf{i})=\pi}|u_{\mathbf{i}}|$ is a sum over one (distinct) index $i\in\{1,\ldots,n\}$ for each block $P$ of $\pi$, for which we have $\sum_{\mathbf{i}:\pi(\mathbf{i})=\pi}|u_{\mathbf{i}}|\leq\prod_{P\in\pi}\sum_{i=1}^{n}|u_{i}|^{|P|}$. Under our assumed normalization $\|\mathbf{u}\|_{2}=1$, we have
$\sum_{i=1}^{n}|u_{i}|\leq\sqrt{n},\quad\quad\sum_{i=1}^{n}|u_{i}|^{k}\leq\|\mathbf{u}\|_{\infty}^{k-2}\sum_{i=1}^{n}|u_{i}|^{2}\leq 1\quad\text{ for any }k\geq 2.$
Thus $\sum_{\mathbf{i}:\pi(\mathbf{i})=\pi}|u_{\mathbf{i}}|\leq n^{|T(\pi)|/2}$. Applying this to (132), we have $\mathbb{E}\|\sum_{i}u_{i}\mathsf{x}_{i}\|_{p}^{l}\leq C_{l}^{\prime}n^{-\alpha l+\epsilon}$. The proof of (b) now follows by the same Markov inequality argument as in part (a).
Part (c) is also similar: By scale invariance, we may consider $U\in\mathbb{C}^{n\times n}$ such that $\sum_{i\neq j}|u_{ij}|^{2}=1$. Fixing $p\in[1,\infty)$ and picking an even integer $l>p$,
$\mathbb{E}\Bigg{\|}\sum_{i\neq j}u_{ij}\mathsf{x}_{ij}\Bigg{\|}_{p}^{l}\leq\mathbb{E}\phi\Bigg{(}\Bigg{[}\Bigg{(}\sum_{i\neq j}u_{ij}\mathsf{x}_{ij}\Bigg{)}\Bigg{(}\sum_{i\neq j}\bar{u}_{ij}\mathsf{x}_{ij}^{*}\Bigg{)}\Bigg{]}^{l/2}\Bigg{)}=\sum_{\mathbf{i},\mathbf{j}}u_{\mathbf{i},\mathbf{j}}\mathbb{E}\phi(\mathsf{x}_{\mathbf{i},\mathbf{j}})$
where
$(\mathbf{i},\mathbf{j})=(i_{1},\ldots,i_{l},j_{1},\ldots,j_{l}),\quad\quad\sum_{\mathbf{i},\mathbf{j}}=\sum_{i_{1}\neq j_{1}}\cdots\sum_{i_{l}\neq j_{l}}$
$u_{\mathbf{i},\mathbf{j}}=u_{i_{1}j_{1}}\bar{u}_{i_{2}j_{2}}\cdots u_{i_{l-1}j_{l-1}}\bar{u}_{i_{l}j_{l}},\quad\quad\mathsf{x}_{\mathbf{i},\mathbf{j}}=\mathsf{x}_{i_{1}j_{1}}\mathsf{x}_{i_{2}j_{2}}^{*}\cdots\mathsf{x}_{i_{l-1}j_{l-1}}\mathsf{x}_{i_{l}j_{l}}^{*}=\prod_{a=1}^{l}\widetilde{\mathsf{x}}_{i_{a}j_{a}}.$
Define $T\equiv T(\mathbf{i},\mathbf{j})$ as the indices that appear exactly once in the combined index list $(\mathbf{i},\mathbf{j})=(i_{1},\ldots,i_{l},j_{1},\ldots,j_{l})$. Then, expanding
$\mathsf{x}_{\mathbf{i},\mathbf{j}}=\sum_{S_{1},\ldots,S_{l}\subseteq T}\mathsf{x}(S_{1},\ldots,S_{l})=\sum_{S_{1},\ldots,S_{l}\subseteq T}\prod_{a=1}^{l}\mathbb{E}_{T\setminus S_{a}}\mathcal{Q}_{S_{a}}[\widetilde{\mathsf{x}}_{i_{a},j_{a}}],$
the same arguments as above using the conditions $\mathbb{E}_{i_{a}}[\widetilde{\mathsf{x}}_{i_{a},j_{a}}]=\mathbb{E}_{j_{a}}[\widetilde{\mathsf{x}}_{i_{a},j_{a}}]=0$ and (41) show
  • $|\mathbb{E}\phi(\mathsf{x}(S_{1},\ldots,S_{l}))|\leq n^{-\alpha l-\sum_{a=1}^{l}|S_{a}\setminus\{i_{a},j_{a}\}|/2+\epsilon}$.
  • If $i_{a}\in T$ (or $j_{a}\in T$), then $\mathsf{x}(S_{1},\ldots,S_{l})=0$ unless $i_{a}\in S_{a}$ (resp. $j_{a}\in S_{a}$).
  • If $i_{a}\in T$ (or $j_{a}\in T$), then $\mathbb{E}\phi(\mathsf{x}(S_{1},\ldots,S_{l}))=0$ unless furthermore $i_{a}\in S_{b}$ (resp. $j_{a}\in S_{b}$) for some $b\neq a$.
Thus $\sum_{a=1}^{l}|S_{a}\setminus\{i_{a},j_{a}\}|\geq|T|$, so we obtain similarly as part (b)
$\mathbb{E}\left\|\sum_{i\neq j}u_{ij}\mathsf{x}_{ij}\right\|_{p}^{l}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{\pi}n^{-|T(\pi)|/2}\sum_{\mathbf{i},\mathbf{j}:\pi(\mathbf{i},\mathbf{j})=\pi}|u_{\mathbf{i},\mathbf{j}}|$
where $\pi$ is the partition of $\{1,\ldots,2l\}$ induced by coincident indices of the combined list $(\mathbf{i},\mathbf{j})$, and $|T(\pi)|$ is the number of singleton blocks of $\pi$. Lemma C.1 applied with
$B_{ij}^{a}=\begin{cases}|u_{ij}|&\text{ if }i\neq j\\ 0&\text{ if }i=j\end{cases}\qquad\text{ for all }a=1,\ldots,l$
shows $\sum_{\pi:|T(\pi)|=t}\sum_{\mathbf{i},\mathbf{j}:\pi(\mathbf{i},\mathbf{j})=\pi}|u_{\mathbf{i},\mathbf{j}}|\leq C_{l}n^{t/2}$. Then the proof of part (c) is concluded by the same Markov inequality argument as in part (a-b). ∎
Appendix D von Neumann algebras and non-commutative $L^{p}$ spaces
We collect here several pieces of background on von Neumann algebras and non-commutative $L^{p}$ spaces that are needed in our main arguments. We refer to [49], [13, Section 1], and [57, Chapter 14] for additional discussion. Throughout, $\mathcal{X}$ is a (finite) von Neumann algebra with faithful, normal, tracial state $\phi$ as in Section 3.
Lemma D.1 ([34], Lemma 3.1).
Suppose $\mathsf{x}\in\mathcal{X}$ is such that for some $\epsilon>0$, either $\Im\mathsf{x}\geq\epsilon 1$ or $\Im\mathsf{x}\leq-\epsilon 1$. Then $\mathsf{x}$ is invertible and $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq 1/\epsilon$.
Lemma D.2 (Conditional expectation. [9], Lemma 1.5.11).
Let $\mathcal{B}\subseteq\mathcal{X}$ be a von Neumann subalgebra. Then there exists a unique linear map $\phi^{\mathcal{B}}:\mathcal{X}\to\mathcal{B}$ (the $\phi$-invariant conditional expectation) that satisfies the following:
  • $\phi^{\mathcal{B}}$ is normal, contractive in the operator norm, and completely positive.
  • For any $\mathsf{y}_{1},\mathsf{y}_{2}\in\mathcal{B}$ and $\mathsf{x}\in\mathcal{X}$, we have $\phi^{\mathcal{B}}[\mathsf{y}_{1}\mathsf{x}\mathsf{y}_{2}]=\mathsf{y}_{1}\phi^{\mathcal{B}}[\mathsf{x}]\mathsf{y}_{2}$.
  • For any $\mathsf{y}\in\mathcal{B}$, $\phi(\mathsf{y})=\phi(\phi^{\mathcal{B}}[\mathsf{y}])$.
Defining $\|\mathsf{x}\|_{p}=\phi(|x|^{p})^{1/p}$, the space $L^{p}(\mathcal{X})$ is the Banach space completion of $\mathcal{X}$ under $\|\cdot\|_{p}$. We set $\|\mathsf{x}\|_{\infty}\equiv\|\mathsf{x}\|_{\mathrm{op}}$ and $L^{\infty}(\mathcal{X})\equiv\mathcal{X}$. These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [49] or [13, Section 1] for this construction.
Lemma D.3 (Non-commutative $L^{p}$-spaces. [57] Theorem 14.1, [13] Proposition 1.1).
For each $p\in[1,\infty)$, $\|\mathsf{x}\|_{p}=\phi(|x|^{p})^{1/p}$ defines a complete norm on $L^{p}(\mathcal{X})$, satisfying
$\|\mathsf{x}\|_{p}\leq\|\mathsf{y}\|_{p}\text{ for all }\mathsf{x},\mathsf{y}\in L^{p}(\mathcal{X})\text{ with }0\leq\mathsf{x}\leq\mathsf{y},\qquad|\phi(x)|\leq\|\mathsf{x}\|_{1}\text{ for all }\mathsf{x}\in L^{1}(\mathcal{X}).$
  1. (a)
    (Hölder’s inequality) For any $1\leq p,q,r\leq\infty$ with $1/p+1/q=1/r$, these norms satisfy
    $\|\mathsf{x}\mathsf{y}\|_{r}\leq\|\mathsf{x}\|_{p}\|\mathsf{y}\|_{q}\text{ forall }\mathsf{x}\in L^{p}(\mathcal{X}),\;\mathsf{y}\in L^{q}(\mathcal{X}).$
    In particular, $\|\mathsf{x}\|_{p}\leq\|\mathsf{x}\|_{q}$ for any $1\leq p\leq q\leq\infty$ so $L^{q}(\mathcal{X})\subseteq L^{p}(\mathcal{X})$, and $\|\mathsf{x}\mathsf{y}\|_{p}\leq\|\mathsf{x}\|_{\mathrm{op}}\|\mathsf{y}\|_{p}$.
  2. (b)
    (Duality) For each $p\in[1,\infty)$, let $q$ be such that $\frac{1}{p}+\frac{1}{q}=1$. Then the map $\mathsf{y}\in L^{q}\mapsto\ell_{\mathsf{y}}\in(L^{p})^{*}$ given by $\ell_{\mathsf{y}}(\mathsf{x})=\phi(\mathsf{x}\mathsf{y})$ is a Banach space isomorphism between $L^{q}$ and the dual $(L^{p})^{*}$ of $L^{p}$.
Lemma D.4 ($L^{p}$-contractivity of conditional expectation).
Let $\mathcal{B}\subseteq\mathcal{X}$ be a von Neumann subalgebra, and let $\phi^{\mathcal{B}}:\mathcal{X}\to\mathcal{B}$ be the unique $\phi$-invariant conditional expectation. Then for any $p\in[1,\infty)$ and $\mathsf{x}\in\mathcal{X}$,
$\|\phi^{\mathcal{B}}(\mathsf{x})\|_{p}\leq\|\mathsf{x}\|_{p}.$
Proof.
For $p=1$, let $\mathsf{y}\in\mathcal{B}$ be the unitary operator for which $\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x})=|\phi^{\mathcal{B}}(\mathsf{x})|$. Then
$\|\phi^{\mathcal{B}}(\mathsf{x})\|_{1}=\phi(\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x}))=\phi(\mathsf{y}\mathsf{x})\leq\|\mathsf{y}\|_{\mathrm{op}}\|\mathsf{x}\|_{1}=\|\mathsf{x}\|_{1}.$
Similarly for $p\in(1,\infty)$, by the density of $\mathcal{X}$ in $L^{q}(\mathcal{X})$ and the above $L^{p}$-$L^{q}$ duality on $\mathcal{X}$ as well as on its subalgebra $\mathcal{B}$,
$\|\phi^{\mathcal{B}}(\mathsf{x})\|_{p}=\sup_{\mathsf{y}\in\mathcal{B}:\|\mathsf{y}\|_{q}=1}\phi(\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x}))=\sup_{\mathsf{y}\in\mathcal{B}:\|\mathsf{y}\|_{q}=1}\phi(\mathsf{y}\mathsf{x})\leq\sup_{\mathsf{y}\in\mathcal{X}:\|\mathsf{y}\|_{q}=1}\phi(\mathsf{y}\mathsf{x})=\|\mathsf{x}\|_{p}.$
Lemma D.5 (Riesz-Thorin interpolation. [13], Proposition 1.6).
Suppose, for some $p_{0},q_{0},p_{1},q_{1}\in[1,\infty]$, that $T:\mathcal{X}\to L^{q_{0}}(\mathcal{X},\phi)\cap L^{q_{1}}(\mathcal{X},\phi)$ is a linear map satisfying
$\|T\mathsf{x}\|_{q_{0}}\leq M_{0}\|\mathsf{x}\|_{p_{0}},\qquad\|T\mathsf{x}\|_{q_{1}}\leq M_{1}\|\mathsf{x}\|_{p_{1}}$
for all $\mathsf{x}\in\mathcal{X}$ and some $M_{0},M_{1}>0$. If $\frac{1}{p_{\theta}}=\frac{1-\theta}{p_{1}}+\frac{\theta}{p_{2}}$ and $\frac{1}{q_{\theta}}=\frac{1-\theta}{q_{1}}+\frac{\theta}{q_{2}}$, then for all $\mathsf{x}\in\mathcal{X}$,
$\|T\mathsf{x}\|_{q_{\theta}}\leq M_{0}^{1-\theta}M_{1}^{\theta}\|\mathsf{x}\|_{p_{\theta}}.$
Acknowledgments
Z.F. would like to thank Cheng Mao, Jiaming Xu, and Yihong Wu for the collaboration [30] that inspired this work. A preliminary version of these results were presented at the Random Matrices and Applications workshop at ICERM, and we would like to thank the workshop participants for their helpful feedback. This research is supported in part by NSF DMS-2142476 and a Sloan Research Fellowship.

### v2

@@
 Here, $\boldsymbol{\theta},\boldsymbol{\xi},\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ are deterministic ridge-regularization and linear constraint parameters, and we will study a setting of random inputs given by independent Wigner matrices $A,B\in\mathbb{R}^{n\times n}$.
-Notably, the optimization variable $X$ has dimension comparable to $A$ and $B$.
+Notably, the optimization variable $X=(x_{ij})_{i,j=1}^{n}$ (with $x_{ij}$ denoting the entries of $X$) has dimension comparable to $A$ and $B$.
 This problem (1) may be written equivalently in terms of the vectorization $\mathbf{x}=\operatorname{vec}(X)\in\mathbb{R}^{n^{2}}$ and diagonal matrices $\Theta=\operatorname{diag}(\boldsymbol{\theta})$ and $\Xi=\operatorname{diag}(\boldsymbol{\xi})$, as
@@
 where $G_{0}\in\mathbb{C}^{n^{2}\times n^{2}}$ is a diagonal deterministic-equivalent matrix.
-Some high-level aspects of this model that may be suggestive of properties to be expected also in nonlinear models such as (2) and (3) include a characterization of the large-$n$ limit by a pair of fixed-point equations in an operator algebra rather than over the scalar field, and a “two-tiered” mean-field structure as reflected in (6), which arises from separate mean-field approximations over $A$ and $B$.
+We remark that although this model (1) is a “toy” setting that simplifies many of the additional complexities of nonlinear models such as (2) and (3), some high-level aspects of this model may be suggestive of properties to be expected also in these nonlinear models, including a characterization of the large-$n$ limit by a pair of fixed-point equations in an operator algebra rather than over the scalar field, and a “two-tiered” mean-field structure as reflected in (6), which arises from separate mean-field approximations over $A$ and $B$.
 A special case of the optimization problem (1) with $\Theta=\Xi=\eta I_{n\times n}$ (and correlated Wigner inputs $A$ and $-B$) was previously analyzed in [31], in the context of the graph matching application.
@@
 In the first stage, conditioning on $B$, the independence structure of $A\otimes I$ suggests a Schur-complement analysis at the level of its $n\times n$ blocks: Denoting $G_{ij}=(\mathbf{e}_{i}\otimes I)^{*}G(\mathbf{e}_{j}\otimes I)\in\mathbb{C}^{n\times n}$ and applying standard resolvent identities, we have
-$G_{ii}=\Big{(}a_{ii}I+B+\theta_{i}\Xi-zI-\sum_{r,s}^{(i)}a_{ir}G_{rs}^{(i)}a_{si}\Big{)}^{-1},\qquad G_{ij}={-}G_{ii}\sum_{r}^{(i)}a_{ir}G_{rj}^{(i)}.$ (9)
+$G_{ii}=\Big{(}a_{ii}I+B+\theta_{i}\Xi-zI-\sum_{r,s}^{(i)}a_{ir}G_{rs}^{(i)}a_{si}\Big{)}^{-1},\qquad G_{ij}={-}G_{ii}\sum_{r}^{(i)}a_{ir}G_{rj}^{(i)},$ (9)
+where $G_{rs}^{(i)}$ is the version of $G_{rs}$ setting to 0 the $i^{\text{th}}$ row and column of $A$, and $\sum_{r}^{(i)}$ is the summation over indices $r\neq i$.
 Applying concentration of the quadratic form in the first expression and averaging over $i=1,\ldots,n$, we will eventually obtain an approximate fixed-point relation for the partial trace
@@
     $\mathcal{A}$ is a von Neumann algebra with unit $1_{\mathcal{A}}$, operator norm $\|\cdot\|_{\mathrm{op}}$, and faithful, normal, tracial state $\tau:\mathcal{A}\to\mathbb{C}$.
+(We review this definition at the start of Section 3.)
   2. (b)
@@
 We remark that $G_{0}(z)$ is a deterministic diagonal matrix in $\mathbb{C}^{n^{2}\times n^{2}}$.
-(We refer to [32, Theorem 4.4] for a previous example of this type of construction, in a different model.)
+(We refer to [33, Theorem 4.4] for a previous example of this type of construction, in a different model.)
 In contrast to more classical random matrix models, the above Stieltjes transform $m_{0}(z)$ in general does not seem to admit a simple characterization in terms of scalar-valued fixed-point equations.
-However, it may be characterized via a pair of fixed-point equations in the operator algebra $\mathcal{A}$, and $G_{0}(z)$ may be constructed from this characterization in the following way.
+In some previously studied models with variance profiles or correlated entries, deterministic equivalents for the Stietljes transform and resolvent may instead be defined via fixed-point equations in a vector or matrix space [5, 26, 4].
+For our model of interest, the most direct characterization of its asymptotic Stieltjes transform seems to be via the following fixed-point equation in the operator algebra $\mathcal{A}$, from which a deterministic equivalent $G_{0}(z)$ for its resolvent may also be constructed.
 Proposition 2.3.
 In the setting of Assumption 2.2, set $\mathcal{A}^{+}=\{\mathsf{x}\in\mathcal{A}:\Im\mathsf{x}\geq\epsilon\text{ forsome }\epsilon>0\}$.
-For any $z\in\mathbb{C}^{+}$, there exist unique elements $\mathsf{m}_{a}(z),\mathsf{m}_{b}(z)\in\mathcal{A}^{+}$ satisfying the fixed-point equations
-$\mathsf{m}_{a}(z)$ $=\frac{1}{n}\sum_{\alpha=1}^{n}(\mathsf{a}+\xi_{\alpha}\Theta-z-\mathsf{m}_{a}(z))^{-1},$ (19)
+For any $z\in\mathbb{C}^{+}$, there exists a unique element $\mathsf{m}_{a}(z)\in\mathcal{A}^{+}$ satisfying the fixed-point equation
+$\mathsf{m}_{a}(z)$ $=\frac{1}{n}\sum_{\alpha=1}^{n}(\mathsf{a}+\xi_{\alpha}\Theta-z-\mathsf{m}_{a}(z))^{-1}.$ (19)
+Similarly, there exists a unique element $\mathsf{m}_{b}(z)\in\mathcal{A}^{+}$ satisfying
 $\mathsf{m}_{b}(z)$ $=\frac{1}{n}\sum_{i=1}^{n}(\mathsf{b}+\theta_{i}\Xi-z-\mathsf{m}_{b}(z))^{-1}.$ (20)
@@
 $G_{0}(z)$ $=\sum_{\alpha=1}^{n}\tau^{\mathcal{D}}\big{[}(\mathsf{a}+\xi_{\alpha}\Theta-\mathsf{z}-\mathsf{m}_{a}(z))^{-1}\big{]}\otimes E_{\alpha\alpha}=\sum_{i=1}^{n}E_{ii}\otimes\tau^{\mathcal{D}}\big{[}(\mathsf{b}+\theta_{i}\Xi-z-\mathsf{m}_{b}(z))^{-1}\big{]}.$
-Our main results for this model are the following estimates for its Stieltjes transform $m(z)$ and resolvent $G(z)$.
+Qualitative approximations of the resolvent $G(z)$ and Stieltjes transform $m(z)$ by the above constructions $G_{0}(z)$ and $m_{0}(z)$ may be obtained using standard techniques of free probability theory, such as a moment expansion of $G(z)$ and $m(z)$ for large $\Im z$ and analytic continuation to the full upper half-plane.
+A primary purpose of our work is to obtain more quantitative estimates that are not readily deduced from these types of methods.
+The following theorem states these estimates, which represent optimal approximations for $G(z)$ and $m(z)$ on the global spectral scale.
 Here, the stochastic domination notation $\xi\prec\zeta$ (c.f.
@@
 The estimates (25), (26), and (27) indicate that the entries of the resolvent fall on three scales of orders 1, $n^{-1/2}$, and $n^{-1}$, as depicted in the numerical simulation of Figure 1.
+In a setting where $\Theta\otimes\Xi$ commutes with $A\otimes I$ and $I\otimes B$, the origins and optimality of these estimates may be understood via a contour integral representation of $G(z)$, as we discuss in Appendix A.
+The above theorem illustrates that various properties of $G(z)$ for general $\Theta\otimes\Xi$ are in fact similar to this commutative case.
 Figure 1.
@@
 3. Preliminaries
-In the remainder of the paper, we prove the preceding results.
+In the remainder of the paper, we prove the preceding results. We note that the bulk of our analysis is in fact carried out for a generalized resolvent operator
+$R=\bigg{(}H\otimes\mathsf{x}-\sum_{k=1}^{K}D_{k}\otimes\mathsf{x}_{k}\bigg{)}^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$
+in the setting of an abstract von Neumann algebra $\mathcal{X}$, which we define more precisely in Section 4.
+This analysis may be of independent interest to models beyond the ones we consider in our work.
 We summarize in this section the technical tools needed for our analyses.
-Section 4 contains the core of our main argument for analyzing a deformed Wigner-type matrix model in an operator-algebra setting.
+Section 4 contains the core of our main argument for analyzing the above generalized resolvent in an operator-algebra setting.
 Section 5 completes the analyses for the Kronecker deformed Wigner model (17), and Section 6 completes the analyses for the least-squares problem (30).
@@
 For a $\mathcal{X}$-valued random variable $\mathsf{x}$ satisfying $\mathbb{E}\|\mathsf{x}\|_{\mathrm{op}}<\infty$, we denote by $\mathbb{E}\mathsf{x}\in\mathcal{X}$ its expectation and $\mathbb{E}[\mathsf{x}\mid\mathscr{G}]$ its conditional expectation with respect to a sub-sigma-field $\mathscr{G}\subset\mathscr{F}$ (c.f.
-[57, Chapter 1]).
+[58, Chapter 1]).
 3.1. Stochastic domination
@@
 Proof.
-The argument is similar to the scalar setting (c.f. [28, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.3),
+The argument is similar to the scalar setting (c.f. [29, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.2),
 $\bigg{\|}\sum_{v\in V}\mathsf{x}(u,v)\bigg{\|}_{p}\leq\sum_{v\in V}\|\mathsf{x}(u,v)\|_{p},\qquad\|\mathsf{x}_{1}\mathsf{x}_{2}\|_{p}\leq\|\mathsf{x}_{1}\|_{2p}\|\mathsf{x}_{2}\|_{2p}.$
@@
 Then (c.f.
-[57, Proposition 1.12])
+[58, Proposition 1.12])
 $\|\mathbb{E}[\mathsf{x}\mid\mathscr{G}]\|_{p}\leq\mathbb{E}[\|\mathsf{x}\|_{p}\mid\mathscr{G}].$ (37)
@@
 The following statements may be deduced from the non-commutative Rosenthal inequalities of [39].
-In the setting of Rademacher variables, similar Khintchine inequalities have been shown in [56, Eq.
-(8.4.11)] and [59, Theorem 6.22].
+In the setting of Rademacher variables, similar Khintchine inequalities have been shown in [57, Eq.
+(8.4.11)] and [60, Theorem 6.22].
 Lemma 3.4.
@@
 with the conventions $\mathbb{E}_{\emptyset}[\mathsf{x}]=\mathcal{Q}_{\emptyset}[\mathsf{x}]=\mathsf{x}$.
-The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [23, Theorem 4.7] and [29, Lemma A.2].
+The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [24, Theorem 4.7] and [30, Lemma A.2].
 Lemma 3.5 (Fluctuation averaging).
@@
 Consider the generalized resolvent
-$R=(H\otimes\mathsf{x}-\mathsf{z})^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X},$
-which exists by Lemma D.1 as ${-}\Im(H\otimes\mathsf{x}-\mathsf{z})=\Im\mathsf{z}\geq\epsilon$ if $\Im\mathsf{z}_{i}\geq\epsilon$ for every $i$.
+$R=(H\otimes\mathsf{x}-\mathsf{z})^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}.$
+We note that this generalized resolvent exists by the following standard lemma, since $\Im(H\otimes\mathsf{x}-\mathsf{z})={-}\Im\mathsf{z}\leq{-}\epsilon$.
+Lemma 3.6 ([35], Lemma 3.1).
+Suppose $\mathsf{x}\in\mathcal{X}$ is such that for some $\epsilon>0$, either $\Im\mathsf{x}\geq\epsilon$ or $\Im\mathsf{x}\leq-\epsilon$.
+Then $\mathsf{x}$ is invertible and $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq 1/\epsilon$.
 For an index set $S\subseteq\{1,\ldots,n\}$, we define $H^{(S)}\in\mathbb{C}^{n\times n}$ and $R^{(S)}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ by
@@
 $H^{(S)}-\mathsf{z}=\begin{pmatrix}h_{11}^{(S)}\mathsf{x}-\mathsf{z}_{1}&h_{12}^{(S)}\mathsf{x}&\cdots&h_{1n}^{(S)}\mathsf{x}\\ h_{21}^{(S)}\mathsf{x}&h_{22}^{(S)}\mathsf{x}-\mathsf{z}_{2}&\cdots&h_{2n}^{(S)}\mathsf{x}\\ \vdots&\vdots&\ddots&\vdots\\ h_{n1}^{(S)}\mathsf{x}&h_{n2}^{(S)}\mathsf{x}&\cdots&h_{nn}^{(S)}\mathsf{x}-\mathsf{z}_{n}\end{pmatrix}\in\mathbb{C}^{n\times n}\otimes\mathcal{X},$
-which are purely algebraic and identical to those of the scalar setting, see e.g. [26, Lemma 4.2]. ∎
+which are purely algebraic and identical to those of the scalar setting, see e.g. [27, Lemma 4.2]. ∎
 3.5. Maximum modulus principle
-We will use the following quantitative version of the maximum modulus principle, following [60, Appendix A].
+We will use the following quantitative version of the maximum modulus principle, following [61, Appendix A].
 Lemma 3.8.
@@
 $r\mapsto\max_{z\in\mathbb{D}:|z|=e^{r}}\log|g(z)|$
-is increasing and convex [35]. Fixing $a>0$, consider the conformal mapping $\psi_{a}:\mathbb{D}\rightarrow\mathbb{C}^{+}$ given by $\psi_{a}(z)=ia\frac{1+z}{1-z}$. For any $z\in\mathbb{D}$,
+is increasing and convex [36, Section 1].
+Fixing $a>0$, consider the conformal mapping $\psi_{a}:\mathbb{D}\rightarrow\mathbb{C}^{+}$ given by $\psi_{a}(z)=ia\frac{1+z}{1-z}$.
+For any $z\in\mathbb{D}$,
 $\left|\psi_{a}(z)-ia\frac{1+|z|^{2}}{1-|z|^{2}}\right|=\left|\frac{2a(z-|z|^{2})}{(1-z)(1-|z|^{2})}\right|=\frac{2a|z|}{1-|z|^{2}}\left|\frac{1-\bar{z}}{1-z}\right|=\frac{2a|z|}{1-|z|^{2}},$
@@
 4.1. Operator fixed-point equations
-Lemma 4.3.
-Let $\mathcal{D}\subset\mathcal{M}\subset\mathcal{A}$ and $\mathcal{X}$ be as in Assumption 4.1.
-Let $\mathsf{x}\in\mathcal{X}$ be self-adjoint, and let $\mathsf{h}\in\mathcal{A}$ be a semicircular element free of $\mathcal{M}$.
-  1. (a)
-    ($\mathcal{D}\otimes\mathcal{X}$-valued fixed point) For any $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}=\{\mathsf{z}\in\mathcal{D}\otimes\mathcal{X}:\Im\mathsf{z}\geq\epsilon\text{ for some }\epsilon>0\}$, there exists a unique solution $\mathsf{s}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ to the fixed-point equation
-    $\mathsf{s}=\left({-}\mathsf{z}-I_{n\times n}\otimes[\mathsf{x}\left(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]\right)\mathsf{x}]\right)^{-1}.$ (48)
-    This solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$.
-  2. (b)
-    ($\mathcal{X}$-valued fixed point) For any $\mathsf{z}_{1},\ldots,\mathsf{z}_{n}\in\mathcal{X}^{+}$, there exists a unique solution $\mathsf{m}\in\mathcal{X}^{+}$ to the fixed-point equation
-    $\mathsf{m}=\frac{1}{n}\sum_{i=1}^{n}\left({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}\mathsf{x}\right)^{-1}.$ (49)
-    Setting $\mathsf{z}=\sum_{i=1}^{n}E_{ii}\otimes\mathsf{z}_{i}\in\mathcal{D}\otimes\mathcal{X}$, this solution is given by $\mathsf{m}_{0}=(\tau\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$.
-Proof.
-For existence and uniqueness in part (a), we apply the general result of [37, Theorem 2.1]: Let $\mathcal{B}$ be any $C^{*}$-algebra and denote its right operator half-plane
-=$\mathcal{B}^{\prime}=\{\mathsf{a}\in\mathcal{B}:\Re\mathsf{a}=\tfrac{1}{2}(\mathsf{a}+\mathsf{a}^{*})\geq\epsilon\text{ for some }\epsilon>0\}.$
-Then for any $\mathsf{b}\in\mathcal{B}^{\prime}$ and any analytic mapping $\eta:\mathcal{B}^{\prime}\to\mathcal{B}^{\prime}$ that is bounded on bounded domains of $\mathcal{B}^{\prime}$, there exists a unique solution $\mathsf{s}\in\mathcal{B}^{\prime}$ to the equation $\mathsf{s}=[\mathsf{b}+\eta(\mathsf{s})]^{-1}$.
-Fixing $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, we take $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$, $\mathsf{b}={-}i\mathsf{z}/2$, and $\eta(\mathsf{s})={-}i\mathsf{z}/2+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$, noting that $\Re[\mathsf{b}]=\Im\mathsf{z}/2\geq\epsilon>0$ and $\Re\eta(\mathsf{s})=\Im z/2+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\Re s])\mathsf{x}\geq\epsilon>0$ for some $\epsilon>0$ whenever $\Re s\geq 0$ and $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, because $\mathsf{x}$ is self-adjoint and $n^{-1}\operatorname{Tr}\otimes 1$ is a positive map (Lemma D.2).
-Thus, there exists a unique solution $\mathsf{s}^{\prime}\in(\mathcal{D}\otimes\mathcal{X})^{\prime}$ to
-$\mathsf{s}^{\prime}=\big{(}{-}i\mathsf{z}+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}^{\prime}])\mathsf{x}\big{)}^{-1}.$
-Multiplying by $i$, there exists a unique solution $\mathsf{s}=i\mathsf{s}^{\prime}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ to (48).
-To show that this solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$, we claim that the map $\mathsf{s}\mapsto I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$ is the $\mathcal{D}\otimes\mathcal{X}$-valued R-transform of $\mathsf{h}\otimes\mathsf{x}$, and (48) is the Cauchy-R relation over $\mathcal{D}\otimes\mathcal{X}$.
-We recall a few relevant notions from operator-valued free probability theory [47, Chapter 9.2]: Let $\operatorname{NC}(m)$ be the lattice of non-crossing partitions of $\{1,\ldots,m\}$.
+In the following we recall a few relevant notions from operator-valued free probability theory [48, Chapter 9.2]: Let $\operatorname{NC}(m)$ be the lattice of non-crossing partitions of $\{1,\ldots,m\}$.
 Associated to any von Neumann subalgebra $\mathcal{B}$ of $\mathcal{Y}=\mathcal{A}\otimes\mathcal{X}$ and its conditional expectation $\tau^{\mathcal{B}}:\mathcal{Y}\to\mathcal{B}$ (c.f.
@@
 $G_{\mathsf{y}}^{\mathcal{B}}(\mathsf{b})=(\mathsf{b}-\mathcal{R}_{\mathsf{y}}^{\mathcal{B}}(G_{\mathsf{y}}^{\mathcal{B}}(\mathsf{b})))^{-1}.$ (50)
-Specializing to our setting with $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$, we claim that for any $\mathsf{s}\in\mathcal{B}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
+Lemma 4.3.
+Let $\mathcal{D}\subset\mathcal{M}\subset\mathcal{A}$ and $\mathcal{X}$ be as in Assumption 4.1.
+Let $\mathsf{x}\in\mathcal{X}$ be self-adjoint, and let $\mathsf{h}\in\mathcal{A}$ be a semicircular element free of $\mathcal{M}$.
+  1. (a)
+    ($\mathcal{D}\otimes\mathcal{X}$-valued fixed point) For any $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}=\{\mathsf{z}\in\mathcal{D}\otimes\mathcal{X}:\Im\mathsf{z}\geq\epsilon\text{ for some }\epsilon>0\}$, there exists a unique solution $\mathsf{s}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ to the fixed-point equation
+    $\mathsf{s}=\left({-}\mathsf{z}-I_{n\times n}\otimes[\mathsf{x}\left(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]\right)\mathsf{x}]\right)^{-1}.$ (51)
+    This solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$.
+  2. (b)
+    ($\mathcal{X}$-valued fixed point) For any $\mathsf{z}_{1},\ldots,\mathsf{z}_{n}\in\mathcal{X}^{+}$, there exists a unique solution $\mathsf{m}\in\mathcal{X}^{+}$ to the fixed-point equation
+    $\mathsf{m}=\frac{1}{n}\sum_{i=1}^{n}\left({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}\mathsf{x}\right)^{-1}.$ (52)
+    Setting $\mathsf{z}=\sum_{i=1}^{n}E_{ii}\otimes\mathsf{z}_{i}\in\mathcal{D}\otimes\mathcal{X}$, this solution is given by $\mathsf{m}_{0}=(\tau\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$.
+Proof.
+For existence and uniqueness in part (a), we apply the general result of [38, Theorem 2.1]: Let $\mathcal{B}$ be any $C^{*}$-algebra and denote its right operator half-plane
+=$\mathcal{B}^{\prime}=\{\mathsf{a}\in\mathcal{B}:\Re\mathsf{a}=\tfrac{1}{2}(\mathsf{a}+\mathsf{a}^{*})\geq\epsilon\text{ for some }\epsilon>0\}.$
+Then for any $\mathsf{b}\in\mathcal{B}^{\prime}$ and any analytic mapping $\eta:\mathcal{B}^{\prime}\to\mathcal{B}^{\prime}$ that is bounded on bounded domains of $\mathcal{B}^{\prime}$, there exists a unique solution $\mathsf{s}\in\mathcal{B}^{\prime}$ to the equation $\mathsf{s}=[\mathsf{b}+\eta(\mathsf{s})]^{-1}$.
+Fixing $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, we take $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$, $\mathsf{b}={-}i\mathsf{z}/2$, and $\eta(\mathsf{s})={-}i\mathsf{z}/2+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$, noting that $\Re[\mathsf{b}]=\Im\mathsf{z}/2\geq\epsilon>0$ and $\Re\eta(\mathsf{s})=\Im z/2+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\Re s])\mathsf{x}\geq\epsilon>0$ for some $\epsilon>0$ whenever $\Re s\geq 0$ and $\mathsf{z}\in(\mathcal{D}\otimes\mathcal{X})^{+}$, because $\mathsf{x}$ is self-adjoint and $n^{-1}\operatorname{Tr}\otimes 1$ is a positive map (Lemma D.1).
+Thus, there exists a unique solution $\mathsf{s}^{\prime}\in(\mathcal{D}\otimes\mathcal{X})^{\prime}$ to
+$\mathsf{s}^{\prime}=\big{(}{-}i\mathsf{z}+I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}^{\prime}])\mathsf{x}\big{)}^{-1}.$
+Multiplying by $i$, there exists a unique solution $\mathsf{s}=i\mathsf{s}^{\prime}\in(\mathcal{D}\otimes\mathcal{X})^{+}$ to (51).
+To show that this solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$, we claim that the map $\mathsf{s}\mapsto I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$ is the $\mathcal{D}\otimes\mathcal{X}$-valued R-transform of $\mathsf{h}\otimes\mathsf{x}$, and (51) is the Cauchy-R relation over $\mathcal{D}\otimes\mathcal{X}$.
+Namely, for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
 $\mathcal{R}_{\mathsf{h}\otimes\mathsf{x}}^{\mathcal{D}\otimes\mathcal{X}}(\mathsf{s})=I_{n\times n}\otimes\mathsf{x}\big{(}n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}]\big{)}\mathsf{x}.$ (53)
 Indeed, since $\mathsf{h}$ is free of $\mathcal{M}$ and hence of $\mathcal{D}\subset\mathcal{M}$, it is readily checked by definition (c.f.
-[47, Definition 9.5]) that $\mathsf{h}\otimes\mathsf{x}$ is free of $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$ with amalgamation over $1\otimes\mathcal{X}$ under the conditional expectation $\tau\otimes 1:\mathcal{A}\otimes\mathcal{X}\to\mathcal{X}\cong 1\otimes\mathcal{X}$.
-Then by [50, Theorem 3.6], for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
+[48, Definition 9.5]) that $\mathsf{h}\otimes\mathsf{x}$ is free of $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$ with amalgamation over $1\otimes\mathcal{X}$ under the conditional expectation $\tau\otimes 1:\mathcal{A}\otimes\mathcal{X}\to\mathcal{X}\cong 1\otimes\mathcal{X}$.
+Then by [51, Theorem 3.6], for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
 $\mathcal{R}_{\mathsf{h}\otimes\mathsf{x}}^{\mathcal{D}\otimes\mathcal{X}}(\mathsf{s})$ $=\sum_{m\geq 1}\kappa_{m}^{\mathcal{D}\otimes\mathcal{X}}((\mathsf{h}\otimes\mathsf{x})\mathsf{s},\ldots,(\mathsf{h}\otimes\mathsf{x})\mathsf{s},\mathsf{h}\otimes\mathsf{x})$
@@
 where $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$ are the scalar-valued free cumulants of $\mathsf{h}$: This may be verified by expressing $\kappa_{m}^{1\otimes\mathcal{X}}(\mathsf{h}\otimes 1,\ldots,\mathsf{h}\otimes 1)$ in terms of moments under the conditional expectation $\tau\otimes 1$ using Möbius-inversion of (48) (c.f.
-[47, Eq.
+[48, Eq.
 (9.19)]), applying the identity $(\tau\otimes 1)[\mathsf{a}\otimes 1]=\tau(\mathsf{a})(1\otimes 1)$ for each moment term, and re-applying (48) to express the result back in terms of the scalar cumulants $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$.
 Here, $\mathsf{h}\in\mathcal{A}$ is semi-circular, so $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})=1$ if $m=2$ and 0 otherwise (c.f.
-[51, Example 11.21]).
+[52, Example 11.21]).
 Thus only the $m=2$ term above remains, and we obtain the claim (53).
@@
 $\leq\|\mathsf{r}(\mathsf{z}+\Delta)\|_{\mathrm{op}}\|\Delta\|_{p}\|\mathsf{r}(\mathsf{z})\|_{\mathrm{op}}\leq 2\delta^{-2}\left\|\Delta\right\|_{p}$
-where we used $L^{p}$-contractivity of $\tau^{\mathcal{D}}\otimes 1$ (Lemma D.4), Hölder’s inequality (Lemma D.3), the given conditions $\Im\mathsf{z}\geq\delta$ and $\Im(\mathsf{z}+\Delta)\geq\delta/2$, and Lemma D.1.
+where we used $L^{p}$-contractivity of $\tau^{\mathcal{D}}\otimes 1$ (Lemma D.3), Hölder’s inequality (Lemma D.2), the given conditions $\Im\mathsf{z}\geq\delta$ and $\Im(\mathsf{z}+\Delta)\geq\delta/2$, and Lemma 3.6.
 Similarly for part (b), defining $\mathsf{m}_{0}(\mathsf{z})=(n^{-1}\operatorname{Tr}\otimes 1)[\mathsf{r}(\mathsf{z})]$ and setting $\mathsf{m}^{\prime}=\mathsf{m}-\Delta$, we have $\mathsf{m}^{\prime}=n^{-1}\sum_{i}({-}\mathsf{z}_{i}-\mathsf{x}\Delta\mathsf{x}-\mathsf{x}\mathsf{m}^{\prime}\mathsf{x})^{-1}$ so Lemma 4.3(a) implies that $\mathsf{m}^{\prime}=\mathsf{m}_{0}(\mathsf{z}+I\otimes\mathsf{x}\Delta\mathsf{x})$.
@@
 $\leq\gamma^{2}\|\mathsf{a}\|_{p}+\|\mathsf{r}\|_{\mathrm{op}}^{2}\|1\otimes\mathsf{a}\|_{p}\leq(\gamma^{2}+\delta^{-2})\|\mathsf{a}\|_{p}.$ (57)
-Here, the second line uses Hölder’s inequality (Lemma D.3), $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$, contractivity of the conditional expectation $\tau\otimes 1$ in $L^{p}$ (Lemma D.4), and $\|\mathsf{r}\|_{\mathrm{op}}\leq\delta^{-1}$ by the assumption $\Im\mathsf{z}\geq\delta$ and Lemma D.1.
+Here, the second line uses Hölder’s inequality (Lemma D.2), $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$, contractivity of the conditional expectation $\tau\otimes 1$ in $L^{p}$ (Lemma D.3), and $\|\mathsf{r}\|_{\mathrm{op}}\leq\delta^{-1}$ by the assumption $\Im\mathsf{z}\geq\delta$ and Lemma 3.6.
 On the other hand, differentiating (56) at $\omega_{0}=\omega(0)$ and using $({-}\mathsf{z}_{i}-\mathsf{x}\omega_{0}\mathsf{x})^{-1}=({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}=(\mathsf{r}_{0})_{ii}$ by Lemma 4.3(a), we have
@@
 Proof.
-See [33, Lemma 3.5]. ∎
+See [34, Lemma 3.5]. ∎
 Lemma 4.7.
@@
 $M^{(iS)}(z)=\left(H^{(iS)}\otimes\mathsf{x}-\mathsf{z}+i(\delta/2)-z\right)^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}.$
-Note that $\Im(\mathsf{z}-i(\delta/2)+z)>\delta/2$ for all $z\in\mathbb{C}^{+}$, so by Lemma D.1, this inverse is well-defined and
+Note that $\Im(\mathsf{z}-i(\delta/2)+z)>\delta/2$ for all $z\in\mathbb{C}^{+}$, so by Lemma 3.6, this inverse is well-defined and
 $\|M^{(iS)}(z)\|_{\mathrm{op}}\leq 2/\delta\text{ for all }z\in\mathbb{C}^{+}.$ (70)
@@
 for some constant $\alpha\in(0,1/2)$ depending only on $\gamma,\upsilon,\delta$, and for all $n\geq n_{0}(\gamma,\upsilon,\delta)$. Thus (68) holds on the event $\mathcal{E}$ for all such $n$.
-We now check that $\mathbb{P}[\mathcal{E}]\geq 1-n^{-D}$ for all $n\geq n_{0}(l,K,D)$: By a standard tail bound for the operator norm (see e.g. [23, Theorem 7.3]), the event
+We now check that $\mathbb{P}[\mathcal{E}]\geq 1-n^{-D}$ for all $n\geq n_{0}(l,K,D)$: By a standard tail bound for the operator norm (see e.g. [24, Theorem 7.3]), the event
 $\mathcal{E}_{0}$ $=\bigcap_{S\subset\{1,\ldots,n\}:|S|\leq l}\;\bigcap_{i\notin S}\left\{\|H^{(iS)}\|_{\mathrm{op}}\leq 3\right\}\subset\{\|H\|_{\mathrm{op}}\leq 3\}$
@@
 $\|R_{ij}^{(S)}\|_{\mathrm{op}}\leq\|R_{ii}^{(S)}\|_{\mathrm{op}}\|\mathsf{x}\|_{\mathrm{op}}\left\|\sum_{r}^{(iS)}h_{ir}R_{rj}^{(iS)}\right\|_{\mathrm{op}}\leq\gamma\delta^{-1}n^{-\alpha}$ (78)
-where we have used also $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma D.1 and the assumption $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$.
+where we have used also $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma 3.6 and the assumption $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$.
 Adjusting the value of $\alpha$ yields the second statement of (77).
@@
 On $\mathcal{E}$, we have $\|\mathrm{I}\|_{\mathrm{op}}\leq n^{-\alpha}$ by (68).
-For $\mathrm{II}$ and $\mathrm{IV}$, Lemma D.1 implies $\|\widetilde{R}_{ii}^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ for both $i\in S$ and $i\notin S$, the latter because $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ and the former because $\Im R^{(S)}\geq 0$ so $\Im R_{jj}^{(S)}\geq 0$ and $\Im\mathsf{z}_{i}\geq\delta$ in the definition (79).
+For $\mathrm{II}$ and $\mathrm{IV}$, Lemma 3.6 implies $\|\widetilde{R}_{ii}^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ for both $i\in S$ and $i\notin S$, the latter because $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ and the former because $\Im R^{(S)}\geq 0$ so $\Im R_{jj}^{(S)}\geq 0$ and $\Im\mathsf{z}_{i}\geq\delta$ in the definition (79).
 Then
@@
 $\frac{1}{n}\sum_{i=1}^{n}R_{ii}=\frac{1}{n}\operatorname{Tr}\otimes 1[\mathsf{r}_{0}]+O_{\prec}(n^{-1/2})=\mathsf{m}_{0}+O_{\prec}(n^{-1/2}).$
-Applying this, the identity $\mathsf{a}^{-1}-\mathsf{b}^{-1}=\mathsf{a}^{-1}(\mathsf{b}-\mathsf{a})\mathsf{b}^{-1}$, and $\|\widehat{R}_{ii}\|_{\mathrm{op}},\|(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma D.1, we may approximate $\widehat{R}_{ii}$ in (104) as
+Applying this, the identity $\mathsf{a}^{-1}-\mathsf{b}^{-1}=\mathsf{a}^{-1}(\mathsf{b}-\mathsf{a})\mathsf{b}^{-1}$, and $\|\widehat{R}_{ii}\|_{\mathrm{op}},\|(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma 3.6, we may approximate $\widehat{R}_{ii}$ in (104) as
 $\widehat{R}_{ii}=(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x}+O_{\prec}(n^{-1/2}))^{-1}=(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}+O_{\prec}(n^{-1/2})=(\mathsf{r}_{0})_{ii}+O_{\prec}(n^{-1/2}),$
@@
 $=\sum_{w\in\mathcal{W}_{m}}\tau\Big{(}\mathsf{d}_{a}\mathsf{U}\mathsf{U}^{\prime}\mathsf{d}_{a}w(\mathsf{d}_{a}\mathsf{a},\mathsf{a}\mathsf{d}_{a},\mathsf{d}_{a},\mathsf{d}_{a})\Big{)}\cdot\tau\Big{(}\mathsf{d}_{b}\mathsf{V}\mathsf{V}^{\prime}\mathsf{d}_{b}w(\mathsf{d}_{b},\mathsf{d}_{b},\mathsf{d}_{b}\mathsf{b},\mathsf{b}\mathsf{d}_{b})\Big{)}.$
-For any $\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1}\in\mathcal{D}$, by the free moment-cumulant relations [51, Proposition 11.4], we have
+For any $\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1}\in\mathcal{D}$, by the free moment-cumulant relations [52, Proposition 11.4], we have
 $\tau(\mathsf{d}_{1}\mathsf{a}\mathsf{d}_{2}\mathsf{a}\ldots\mathsf{d}_{k}\mathsf{a}\mathsf{d}_{k+1})=\sum_{\pi\in\operatorname{NC}(2k+1)}\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})$
 where $\operatorname{NC}(2k+1)$ is the set of non-crossing partitions of $(1,2,\ldots,2k+1)$ and $\kappa_{\pi}$ is the free cumulant associated to each $\pi\in\operatorname{NC}(2k+1)$.
-Since $\mathsf{a}$ is free of $\mathcal{D}$ and has $2^{\text{nd}}$ free cumulant equal to 1 and remaining free cumulants 0 [51, Example 11.21], we have $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})=0$ unless the blocks of $\pi$ containing $\mathsf{a}$ constitute a non-crossing pairing and are disjoint from those containing $(\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1})$.
+Since $\mathsf{a}$ is free of $\mathcal{D}$ and has $2^{\text{nd}}$ free cumulant equal to 1 and remaining free cumulants 0 [52, Example 11.21], we have $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})=0$ unless the blocks of $\pi$ containing $\mathsf{a}$ constitute a non-crossing pairing and are disjoint from those containing $(\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1})$.
 Let $\operatorname{NC}_{2}(k)$ denote the set of all non-crossing pairings of $\{2,4,6,\ldots,2k\}$ corresponding to the locations of the $\mathsf{a}$’s, and for each $\rho\in\operatorname{NC}_{2}(k)$, let $K(\rho)\in\operatorname{NC}(k+1)$ be its complement in $\{1,2,\ldots,2k+1\}$, i.e.
@@
 Lemma B.1 ([39], Theorem 2.1).
-Let $\mathcal{Y}\subset\mathcal{X}$ be a von Neumann subalgebra with $\phi$-invariant conditional expectation $\phi^{\mathcal{Y}}:\mathcal{X}\to\mathcal{Y}$.
+Let $\mathcal{Y}\subset\mathcal{X}$ be a von Neumann subalgebra with $\phi$-invariant conditional expectation (c.f.
+Lemma D.1) $\phi^{\mathcal{Y}}:\mathcal{X}\to\mathcal{Y}$.
 Suppose $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}\in L^{p}(\mathcal{X})$ satisfy $\phi^{\mathcal{Y}}(\mathsf{x}_{i})=0$, and are independent over $\mathcal{Y}$ in the sense that for each $i$, each $\mathsf{x}$ in the von Neumann subalgebra generated by $\mathsf{x}_{i}$, and each $\mathsf{x}^{\prime}$ in the von Neumann subalgebra generated by $\{\mathsf{x}_{j}:j\neq i\}$, we have $\phi^{\mathcal{Y}}(\mathsf{x}\mathsf{x}^{\prime})=\phi^{\mathcal{Y}}(\mathsf{x})\phi^{\mathcal{Y}}(\mathsf{x}^{\prime})$.
@@
 Proof.
-By [14, Theorem 1], for a universal constant $C>0$,
+By [15, Theorem 1], for a universal constant $C>0$,
 […]
@@
 For (a), fix any $p\geq 2$.
-We apply Lemma B.1 in the setting of [38, Example 1.3]: Let $L^{\infty}(\Omega)$ be the von Neumann algebra of bounded scalar random variables over the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$, and consider $\mathcal{M}=L^{\infty}(\Omega)\otimes\mathcal{X}$ equipped with the state $\mathbb{E}\circ\phi$.
+We apply Lemma B.1 in the setting of [39, Example 1.3]: Let $L^{\infty}(\Omega)$ be the von Neumann algebra of bounded scalar random variables over the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$, and consider $\mathcal{M}=L^{\infty}(\Omega)\otimes\mathcal{X}$ equipped with the state $\mathbb{E}\circ\phi$.
 Then $\alpha_{i}\mathsf{x}_{i}\in L^{p}(\mathcal{M},\mathbb{E}\circ\phi)$, $\mathbb{E}:\mathcal{M}\to\mathcal{X}$ coincides with the conditional expectation onto the subalgebra $\mathcal{X}\subset\mathcal{M}$, and $\{\alpha_{i}\mathsf{x}_{i}\}_{i=1}^{n}$ are independent over $\mathcal{X}$ in the sense of Lemma B.1, so Lemma B.1 shows
@@
 $\mathbb{E}\bigg{[}\bigg{\|}\sum_{i=1}^{n}\alpha_{i}\mathsf{x}_{i}\bigg{\|}_{p}^{p}\bigg{]}\leq C_{p}^{\prime}\max\bigg{\{}\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}\mathsf{x}_{i}^{*}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\;\bigg{\|}\bigg{(}\sum_{i=1}^{n}\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg{)}^{1/2}\bigg{\|}_{p}^{p},\;\sum_{i=1}^{n}\|\mathsf{x}_{i}\|_{p}^{p}\bigg{\}}.$ (127)
-Here, for $p\geq 2$, the third term is bounded by the second by the following argument (see also [38, Eq.
-(2.4)]): Consider $\hat{\mathsf{x}}=\sum_{i=1}^{n}E_{i1}\otimes\mathsf{x}_{i}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ equipped with the trace $n^{-1}\operatorname{Tr}\otimes\phi$, and the linear map $T:\mathbb{C}^{n\times n}\otimes\mathcal{X}\to\mathbb{C}^{n\times n}\otimes\mathcal{X}$ defined by $T(E_{ij}\otimes\mathsf{x})=E_{i,j+i-1}\otimes\mathsf{x}$ where $j+i-1$ is interpreted modulo $n$.
+Here, for $p\geq 2$, the third term is bounded by the second by the following argument (see also [39, Eq.
+(2.4)]): Consider $\hat{\mathsf{x}}=\sum_{i=1}^{n}E_{i1}\otimes\mathsf{x}_{i}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ equipped with the trace $n^{-1}\operatorname{Tr}\otimes\phi$, and the linear map $T:\mathbb{C}^{n\times n}\otimes\mathcal{X}\to\mathbb{C}^{n\times n}\otimes\mathcal{X}$ defined by
+$T(E_{ij}\otimes\mathsf{x})=\begin{cases}E_{i,j+i-1}\otimes\mathsf{x}&\text{if $j=1$}\\ 0&\text{otherwise}\end{cases}$
+where $j+i-1$ is interpreted modulo $n$.
 Thus $\hat{\mathsf{x}}$ has $\mathcal{X}$-valued entries $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}$ along the first column, and $T(\hat{\mathsf{x}})$ has these entries instead along the main diagonal.
@@
 Appendix C Fluctuation averaging
-In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [29, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [29].
+In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [30, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [30].
 Lemma C.1 ([30], Lemma A.3).
@@
 We collect here several pieces of background on von Neumann algebras and non-commutative $L^{p}$ spaces that are needed in our main arguments.
-We refer to [49], [13, Section 1], and [57, Chapter 14] for additional discussion.
+We refer to [50], [14, Section 1], and [58, Chapter 14] for additional discussion.
 Throughout, $\mathcal{X}$ is a (finite) von Neumann algebra with faithful, normal, tracial state $\phi$ as in Section 3.
-Lemma D.1 ([34], Lemma 3.1).
-Suppose $\mathsf{x}\in\mathcal{X}$ is such that for some $\epsilon>0$, either $\Im\mathsf{x}\geq\epsilon 1$ or $\Im\mathsf{x}\leq-\epsilon 1$.
-Then $\mathsf{x}$ is invertible and $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq 1/\epsilon$.
 Lemma D.1 (Conditional expectation. [10], Lemma 1.5.11).
@@
 We set $\|\mathsf{x}\|_{\infty}\equiv\|\mathsf{x}\|_{\mathrm{op}}$ and $L^{\infty}(\mathcal{X})\equiv\mathcal{X}$.
-These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [49] or [13, Section 1] for this construction.
+These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [50] or [14, Section 1] for this construction.
 Lemma D.2 (Non-commutative $L^{p}$-spaces. [58] Theorem 14.1, [14] Proposition 1.1).
@@
 Proof.
-For $p=1$, let $\mathsf{y}\in\mathcal{B}$ be the unitary operator for which $\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x})=|\phi^{\mathcal{B}}(\mathsf{x})|$. Then
+For $p=1$, let $\mathsf{y}\in\mathcal{B}$ be the projection operator for which $\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x})=|\phi^{\mathcal{B}}(\mathsf{x})|$. Then
 $\|\phi^{\mathcal{B}}(\mathsf{x})\|_{1}=\phi(\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x}))=\phi(\mathsf{y}\mathsf{x})\leq\|\mathsf{y}\|_{\mathrm{op}}\|\mathsf{x}\|_{1}=\|\mathsf{x}\|_{1}.$

### v3

@@
 and the bilinear-form estimates
-$(\mathbf{u}\otimes\mathbf{v})^{*}[G-G_{0}](\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})\prec n^{-1/2}\text{ for deterministic unit vectors }\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{C}^{n},$ (7)
+[…]
 where $G_{0}\in\mathbb{C}^{n^{2}\times n^{2}}$ is a diagonal deterministic-equivalent matrix.
 We remark that although this model (1) is a “toy” setting that simplifies many of the additional complexities of nonlinear models such as (2) and (3), some high-level aspects of this model may be suggestive of properties to be expected also in these nonlinear models, including a characterization of the large-$n$ limit by a pair of fixed-point equations in an operator algebra rather than over the scalar field, and a “two-tiered” mean-field structure as reflected in (6), which arises from separate mean-field approximations over $A$ and $B$.
-A special case of the optimization problem (1) with $\Theta=\Xi=\eta I_{n\times n}$ (and correlated Wigner inputs $A$ and $-B$) was previously analyzed in [31], in the context of the graph matching application.
+A special case of the optimization problem (1) with $\Theta=\Xi=\eta I_{n\times n}$ (and correlated Wigner inputs $A$ and B-B) was previously analyzed in [31], in the context of the graph matching application.
 We review a central idea of this previous analysis in Appendix A, which, however, is special to a commutative setting and does not extend more generally.
@@
 As the implicit dependence of $M_{B}$ on $B$ in (10) is rather complex and makes a direct Schur-complement analysis difficult, we introduce an explicit operator algebra representation and write
-[…]
+$M_{B}=(\tau\otimes I)\Big[\underbrace{(\mathsf{a}\otimes I+1\otimes B+\Theta\otimes\Xi-z\,1\otimes I)^{-1}}_{=\tilde{\mathsf{g}}}\Big$] (11)
 for a ($n$-dependent) von Neumann algebra $\mathcal{A}$ with trace $\tau$, containing a semicircular variable $\mathsf{a}$ and the subalgebra $\mathbb{C}^{n\times n}$ free of $\mathsf{a}$ (Lemma 4.3).
@@
 This type of operator associated to Dyson fixed-point equations has been studied previously in [2, 4], and we adapt a Perron-Frobenius argument of [2] to show quantitative invertibility of $\mathcal{L}_{2}$ for any fixed $z\in\mathbb{C}^{+}$.
-In our context, we require invertibility of $\mathcal{L}_{2}$ in the $L^{p}$-norm for each $p\in[1,\infty)$, and in infinite-dimensional settings where $\mathcal{L}_{2}$ may not have an exact Perron-Frobenius eigenvector — we thus carry out this analysis in the $L^{1}$-$L^{\infty}$ duality rather than the Hilbert-space setting of [2], and appeal to the Riesz-Thorin interpolation to obtain invertibility for all $p$ (Lemma 4.7).
+In our context, we require invertibility of $\mathcal{L}_{2}$ in the $L^{p}$-norm for each $p\in[1,\infty)$, and in infinite-dimensional settings where $\mathcal{L}_{2}$ may not have an exact Perron-Frobenius eigenvector — we thus carry out this analysis in the $^{1}-LL^{\infty}$ duality rather than the Hilbert-space setting of [2], and appeal to the Riesz-Thorin interpolation to obtain invertibility for all $p$ (Lemma 4.7).
 This shows $n^{-1}\sum_{r}G_{rj}^{(i)}G_{rj}^{(i)*}\prec n^{-1}+n^{-3\alpha}$, which applied to (14) shows the implication
@@
   3. (c)
-[…]
+    $\mathcal{A}$ has two semicircular elements $\mathsf{a},\mathsf{b}$ (i.e.
+satisfying $\tau(\mathsf{a}^{k})=\tau(\mathsf{b}^{k})=\int_{-2}^{2}x^{k}\,\frac{1}{2\pi}\sqrt{4-x^{2}}\,dx$ for each $k\geq 1$) which are free of $\mathcal{M}$ with respect to $\tau$.
 Let $\mathcal{D}\subset\mathcal{M}$ be the subalgebra of diagonal matrices, generated by the diagonal basis elements $\{E_{ii}\}_{i=1}^{n}$, and denote by $\tau^{\mathcal{D}}:\mathcal{A}\to\mathcal{D}$ the diagonal projection
@@
 For spectral arguments $z\in\mathbb{C}^{+}$, we define its (deterministic, $n$-dependent) resolvent and Stieltjes transform by
-[…]
+$\mathsf{g}(z)$ $=(\mathsf{q}-z\,1\otimes 1)^{-1},$
 $m_{0}(z)$ $=\tau\otimes\tau[\mathsf{g}(z)]\in\mathbb{C}^{+}.$
@@
 $\widehat{X}=\operatorname*{arg\,min}_{X\in\mathbb{R}^{n\times n}}f(X)\qquad\text{ subject to }\qquad\frac{1}{n}\mathbf{v}^{*}X\mathbf{u}=1.$ (30)
-Our main results for this optimization problem (30) are the following asymptotic characterization of the minimum objective value $f(\widehat{X})$ and the values of linear forms $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$ for deterministic test vectors $\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
+Our main results for this optimization problem (30) are the following asymptotic characterization of the minimum objective value f(X^)f($\widehat${X}) and the values of linear forms $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$ for deterministic test vectors $\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
 Theorem 2.5.
@@
 $n^{-2}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}(\mathbf{u}\otimes\mathbf{v}),$ (32)
-where $(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}\in\mathbb{R}^{n^{2}\times n^{2}}$ is a deterministic-equivalent approximation for the matrix $[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}$ arising in the vectorization of the objective (29), and these expressions (31) and (32) coincide because this deterministic-equivalent matrix is diagonal.
-In an asymptotic setting, if the empirical distributions of coordinates of $\Theta,\Xi,\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}$ converge to deterministic limits, then a qualitative implication of Theorem 2.5 is a characterization of the almost-sure limit values of $f(\widehat{X})$ and $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$, as summarized in the following corollary.
+where $(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}\in\mathbb{R}^{n^{2}\times n^{2}}$ is a deterministic-equivalent approximation for the matrix $(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}$ arising in the vectorization of the objective (29), and these expressions (31) and (32) coincide because this deterministic-equivalent matrix is diagonal.
+In an asymptotic setting, if the empirical distributions of coordinates of $\Theta,\Xi,\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}$ converge to deterministic limits, then a qualitative implication of Theorem 2.5 is a characterization of the almost-sure limit values of f(X^)f($\widehat${X}) and $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$, as summarized in the following corollary.
 Corollary 2.6.
@@
 Figure 2.
-Values of $1/(2f(\widehat{X}))$ obtained from solving (30) across 10 independent realizations (solid dots, with vertical lines indicating 1 standard deviation) versus the theoretical prediction $T(\mathcal{P},\mathcal{Q})$ computed from Proposition 2.7 with $M=12$ (dashed lines).
+Values of $)1/(2f(\widehat{X}$)) obtained from solving (30) across 10 independent realizations (solid dots, with vertical lines indicating 1 standard deviation) versus the theoretical prediction $T(\mathcal{P},\mathcal{Q})$ computed from Proposition 2.7 with $M=12$ (dashed lines).
 Here $A,B$ are GOE matrices of size $n=1000$, we take $\theta_{i}^{-1},\xi_{\alpha}^{-1},u_{i}^{2},v_{\alpha}^{2}\sim\text{Uniform}(0.05,0.5)/k$ with $k\in\{1,1.3,3\}$, and the horizontal axis indicates the correlation between coordinate pairs $(\theta_{i}^{-1},u_{i}^{2})$ and between $(\xi_{\alpha}^{-1},v_{\alpha}^{2})$.
@@
 Proof.
-The argument is similar to the scalar setting (c.f. [29, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.2),
+The argument is similar to the scalar setting (c.f. [30, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.2),
 $\bigg\|\sum_{v\in V}\mathsf{x}(u,v)\bigg\|_{p}\leq\sum_{v\in V}\|\mathsf{x}(u,v)\|_{p},\qquad\|\mathsf{x}_{1}\mathsf{x}_{2}\|_{p}\leq\|\mathsf{x}_{1}\|_{2p}\|\mathsf{x}_{2}\|_{2p}.$
@@
 with the conventions $\mathbb{E}_{\emptyset}[\mathsf{x}]=\mathcal{Q}_{\emptyset}[\mathsf{x}]=\mathsf{x}$.
-The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [24, Theorem 4.7] and [30, Lemma A.2].
+The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [24, Theorem 4.7] and [29, Lemma A.2].
 Lemma 3.5 (Fluctuation averaging).
@@
 For existence and uniqueness in part (a), we apply the general result of [38, Theorem 2.1]: Let $\mathcal{B}$ be any $C^{*}$-algebra and denote its right operator half-plane
-=$\mathcal{B}^{\prime}=\{\mathsf{a}\in\mathcal{B}:\Re\mathsf{a}=\tfrac{1}{2}(\mathsf{a}+\mathsf{a}^{*})\geq\epsilon\text{ for some }\epsilon>0\}.$
+$\mathcal{B}^{\prime}=\{\mathsf{a}\in\mathcal{B}:\Re\mathsf{a}=\tfrac{1}{2}(\mathsf{a}+\mathsf{a}^{*})\geq\epsilon\text{ for some }\epsilon>0\}.$
 Then for any $\mathsf{b}\in\mathcal{B}^{\prime}$ and any analytic mapping $\eta:\mathcal{B}^{\prime}\to\mathcal{B}^{\prime}$ that is bounded on bounded domains of $\mathcal{B}^{\prime}$, there exists a unique solution $\mathsf{s}\in\mathcal{B}^{\prime}$ to the equation $\mathsf{s}=[\mathsf{b}+\eta(\mathsf{s})]^{-1}$.
@@
 Lemma 4.6 below).
-For reasons that will be clear in the proof, the approximation error must be controlled in $L^{1}$ rather than $L^{2}$, and thus we implement a version of this argument in the $L^{1}$-$L^{\infty}$ duality rather than in a Hilbert space setting, to obtain a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{\infty}\to L^{\infty}}$.
+For reasons that will be clear in the proof, the approximation error must be controlled in $L^{1}$ rather than $L^{2}$, and thus we implement a version of this argument in the $^{1}-LL^{\infty}$ duality rather than in a Hilbert space setting, to obtain a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{\infty}\to L^{\infty}}$.
 Second, as we will also require a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ for each $p\in[1,\infty)$, we carry out a dual version of this argument to obtain also a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{1}\to L^{1}}$, and hence deduce a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ via the Riesz-Thorin interpolation.
@@
 Applying these bounds above and then taking $\epsilon\to 0$, we obtain
-$01-r(\mathcal{F})\geq\delta c(\gamma,\upsilon,\delta)>$ 0 . (63)
+$1-r(\mathcal{F})\geq\delta c(\gamma,\upsilon,\delta)>0.$ (63)
 Since $\mathcal{F}^{\prime}$ is the adjoint of $\mathcal{F}$ and thus shares its spectrum, this shows $r(\mathcal{F}^{\prime})=r(\mathcal{F})<1$.
@@
 Thus, in place of (60) we have
-$(\mathrm{Id}-\mathcal{G}^{\prime})[\mathsf{x}(\Im\mathsf{m}_{0})\mathsf{x}]=\mathsf{w},\qquad\text{ for }\mathsf{w}=\frac{1}{n}\sum_{i=1}^{n}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}[\Im\mathsf{z}_{i}](\mathsf{r}_{0})_{ii}\mathsf{x}.$ (66)
+[…]
 In place of (61) and (62), we may apply $\phi(\mathsf{v}\mathsf{x}[\Im\mathsf{m}_{0}]\mathsf{x})\leq\gamma^{2}\delta^{-1}$ and $\mathsf{w}\geq\gamma^{-2}c(\gamma,\upsilon,\delta)$, where the first inequality uses $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$ and the second uses $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$ and $c(\gamma,\upsilon,\delta)$ as defined in (62).
-Then, multiplying (66) by $\mathsf{v}$, taking the trace, and then taking the limit $\epsilon\to 0$, we obtain similarly to (63) that $01-r(\mathcal{G})\geq\delta\gamma^{-4}c(\gamma,\upsilon,\delta)>$ 0.
+Then, multiplying (66) by $\mathsf{v}$, taking the trace, and then taking the limit $\epsilon\to 0$, we obtain similarly to (63) that $1-r(\mathcal{G})\geq\delta\gamma^{-4}c(\gamma,\upsilon,\delta)>0$.
 This implies that $\mathrm{Id}-\mathcal{G}^{\prime}:L^{\infty}\to L^{\infty}$ is invertible with positivity-preserving inverse, and repeating the preceding arguments gives, for any $\mathsf{a}\in L^{\infty}$,
@@
 $\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(i\delta/2)|$ $\leq\sup_{z\in S_{r_{1}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|$
-$\leq\frac{|r_{1}-r_{0}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{0}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|+\frac{|r_{1}-r_{2}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{2}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|$
+$\leq\frac{|r_{1}-r_{2}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{0}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|+\frac{|r_{1}-r_{0}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{2}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|$
 $\leq-\alpha\log n$
@@
 $\mathbf{h}_{i}^{(iS)*}M\mathbf{h}_{i}^{(iS)}-n^{-1}\operatorname{Tr}^{(iS)}M\prec n^{-1}\left(\sum_{r,s}^{(iS)}|M_{rs}|^{2}\right)^{1/2}\prec n^{-1/2}\|M\|_{\mathrm{op}}.$ (75)
-Note that $w(D^{(iS)})$ is independent of $\mathbf{h}_{i}^{(iS)}$, and |$|\{S\subset\{1,\ldots,n\}:|S|\leq l\}|\leq n^{l}$ and $|\mathcal{W}_{t}|=(K+2)^{t}\leq n^{C\log(K+2)}$ for all $t\leq T(n)\leq C\log n$.
+Note that $w(D^{(iS)})$ is independent of $\mathbf{h}_{i}^{(iS)}$, and $|\{S\subset\{1,\ldots,n\}:|S|\leq l\}|\leq n^{l}$ and $|\mathcal{W}_{t}|=(K+2)^{t}\leq n^{C\log(K+2)}$ for all $t\leq T(n)\leq C\log n$.
 Then, fixing $\epsilon^{\prime}=0.05$, applying (75) to each matrix $M=w(D^{(iS)})$ conditional on $H^{(iS)}$, and taking a union bound, the event
@@
 $={-}h_{ii}\mathsf{x}+\mathsf{x}\Bigg[\underbrace{\sum_{r}^{(iS)}\left(|h_{ir}|^{2}-\frac{1}{n}\right)R_{rr}^{(iS)}+\sum_{r\neq s}^{(iS)}h_{ir}R_{rs}^{(iS)}h_{si}}_{:=\mathrm{I}}-\underbrace{\frac{1}{n}\sum_{j\in S\cup\{i\}}\widetilde{R}_{jj}^{(S)}}_{:=\mathrm{II}}+\underbrace{\frac{1}{n}\sum_{j}^{(iS)}(R_{jj}^{(iS)}-R_{jj}^{(S)})}_{:=\mathrm{III}}\Bigg]\mathsf{x}$
-[…]
+$\text{ for all }i\notin S$ (81)
 and
@@
 Collecting these bounds and applying also $|h_{ii}|\leq n^{-\alpha}$ on $\mathcal{E}$, we have for a constant $C(\gamma,\delta)>0$ and every $i=1,\ldots,n$ that
-$\|\Delta_{i}^{(S)}\|_{\mathrm{op}}\leq C(\gamma,\delta)n^{-\alpha}\text{ on }\mathcal{E}.$ (85)
+[…]
 Taking the Kronecker product with $E_{ii}$ on both sides of (80), summing over $i=1,\ldots,n$, recalling $\mathsf{z}=\sum_{i}E_{ii}\otimes\mathsf{z}_{i}$, and setting $\Delta^{(S)}=\sum_{i}E_{ii}\otimes\Delta_{i}^{(S)}$, we have
@@
 $G(z)$ $=(Q-z\,I\otimes I)^{-1},\qquad m(z)=n^{-2}\operatorname{Tr}G(z),$
-[…]
+$\mathsf{g}(z)$ $=(\mathsf{q}-z\,1\otimes 1)^{-1},\qquad G_{0}(z)=(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[\mathsf{g}(z)],\qquad m_{0}(z)=\tau\otimes\tau[\mathsf{g}(z)].$
 We first show Proposition 2.3 and all statements of Theorem 2.4 except the estimate $G_{ij,\alpha\beta}\prec n^{-1}$ of (27) using the analyses of Section 4.
@@
 and abbreviate $G=G(z)$, $\tilde{\mathsf{g}}=\tilde{\mathsf{g}}(z)$, and $\mathsf{g}=\mathsf{g}(z)$.
-Fix any $\epsilon>0$ and consider the event =$\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$.
+Fix any $\epsilon>0$ and consider the event $\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$.
 We apply Theorem 4.2 conditional on $B$ and this event $\mathcal{E}$, with $\mathcal{X}=\mathbb{C}^{n\times n}$, $R=G$, $\mathsf{r}_{0}=\tilde{\mathsf{g}}$, $H=A$, $\mathsf{h}=\mathsf{a}$, $\mathsf{x}=I$, and
@@
 To prove the remaining statements (21), (25), and (28), we apply Theorem 4.2 again to the second tensor factor, with $\mathcal{X}=\mathcal{A}$, $R=\tilde{\mathsf{g}}$, $\mathsf{r}_{0}=\mathsf{g}$, $H=B$, $\mathsf{h}=\mathsf{b}$, $\mathsf{x}=1_{\mathcal{A}}$, and
-[…]
+$\mathsf{z}={-}\mathsf{a}\otimes I-\Theta\otimes\Xi+z\,1_{\mathcal{A}}\otimes I.$
 This gives, uniformly in $\alpha\neq\beta$ and $\mathbf{v},\mathbf{v}^{\prime}\in\mathbb{C}^{n}$ with $\|\mathbf{v}\|_{2},\|\mathbf{v}^{\prime}\|_{2}\leq\upsilon$,
@@
 We recall our convention of Roman indices $i,j,k,\ldots$ for the first tensor factor and Greek indices $\alpha,\beta,\gamma,\ldots$ for the second tensor factor.
-In the following, we use a superscript ()($\cdot$)( ⋅ ) to denote a minor on $A$ and []delimited-[][$\cdot$][ ⋅ ] to denote a minor on $B$.
+In the following, we use a superscript $(\cdot)$ to denote a minor on $A$ and $\cdot$] to denote a minor on $B$.
 Fix any indices $i\neq j$ and $\alpha\neq\beta$. Applying Lemma 3.7(b) to the second tensor factor,
@@
 $P_{0}$ $=(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[\mathsf{p}].$
-Simple calculus yields the explicit forms for $\hat{\mathbf{x}}$ and $f(\hat{\mathbf{x}})$ as
+[…]
 $\hat{\mathbf{x}}=\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P(\mathbf{u}\otimes\mathbf{v})}\cdot n^{-1}P(\mathbf{u}\otimes\mathbf{v}),\qquad f(\hat{\mathbf{x}})=\frac{1}{2}\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P(\mathbf{u}\otimes\mathbf{v})}.$ (120)
@@
 Denote $\mathbf{e}_{1}=(1,0)\in\mathbb{C}^{2}$. Then by Schur’s complement,
-$-i\widetilde{P}=\begin{bmatrix}\Theta\otimes\Xi&i(A\otimes I+I\otimes B)\\ i(A\otimes I+I\otimes B)&I\otimes I\end{bmatrix}^{-1}=\begin{bmatrix}[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}&*\\ *&*\end{bmatrix}$ (121)
+$i\widetilde{P}=\begin{bmatrix}\Theta\otimes\Xi&i(A\otimes I+I\otimes B)\\ i(A\otimes I+I\otimes B)&I\otimes I\end{bmatrix}^{-1}=\begin{bmatrix}[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}&*\\ *&*\end{bmatrix}$ (121)
 so that $P=(\mathbf{e}_{1}\otimes I\otimes I)^{*}[-i\widetilde{P}](\mathbf{e}_{1}\otimes I\otimes I)$. Defining also
@@
 $=\frac{1}{n^{2}}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})+O_{\prec}(n^{-1/2}).$ (123)
-Here, in the second line, we have applied (47) of Theorem 4.2 twice as in the proof of Theorem 2.4 in Section 5, first to the second tensor factor conditional on $B$ and the event =$\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$ with $\mathcal{X}=\mathbb{C}^{2\times 2}\otimes\mathbb{C}^{n\times n}$ (the product of first and third factors),
+Here, in the second line, we have applied (47) of Theorem 4.2 twice as in the proof of Theorem 2.4 in Section 5, first to the second tensor factor conditional on $B$ and the event $\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$ with $\mathcal{X}=\mathbb{C}^{2\times 2}\otimes\mathbb{C}^{n\times n}$ (the product of first and third factors),
 $H=A,\quad\mathsf{x}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\otimes I,\quad\mathsf{z}=i\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes\Theta\otimes\Xi+i\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\otimes I\otimes I-\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\otimes I\otimes B,$
@@
 so we may apply the approximation (123) to both the numerator and denominator of (120) to get
-$f(\hat{\mathbf{x}})$ $=\frac{1}{2}\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})}+O_{\prec}(n^{-1/2}),$
+[…]
 $\frac{1}{n}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}\hat{\mathbf{x}}$ $=\frac{n^{-2}(\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})}+O_{\prec}(n^{-1/2}).$
@@
 where $\Theta,\mathsf{u},\mathsf{u}^{\prime},\Xi,\mathsf{v},\mathsf{v}^{\prime}\in\mathcal{D}\subset\mathcal{A}$ on the left are real diagonal matrices, and $\theta,\mathsf{U},\mathsf{U}^{\prime},\xi,\mathsf{V},\mathsf{V}^{\prime}\in\mathcal{A}$ on the right are commuting, self-adjoint limiting operators, free of $(\mathsf{a},\mathsf{b})$ and for which $(\theta,\mathsf{U},\mathsf{U}^{\prime})$ and $(\xi,\mathsf{V},\mathsf{V}^{\prime})$ have joint laws under $\tau$ given by $\mathcal{P}$ and $\mathcal{Q}$, respectively.
-The assumptions $\|\Theta\|_{\mathrm{op}},\|\Xi\|_{\mathrm{op}}\leq\upsilon$ and $\Theta,\Xi\geq\delta$ imply that the spectrum of $(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi$ is contained in $[\delta^{2},16+\upsilon^{2}]$.
+The assumptions $\|\Theta\|_{\mathrm{op}},\|\Xi\|_{\mathrm{op}}\leq\upsilon$ and $\Theta,\Xi\geq\delta$ imply that the spectrum of $(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi$ is contained in $\delta^{2},16+\upsilon^{2}]$.
 The inverse function $x\mapsto x^{-1}$ may be approximated uniformly by polynomials on this interval, so the above convergence implies
@@
 We set $\eta=\min\{\sqrt{x_{a}x_{b}}:x_{a}\in\mathrm{supp}(\theta),\,x_{b}\in\mathrm{supp}(\xi)\}$ and define a limiting linearized operator $\tilde{\mathsf{p}}$ analogous to (122),
-[…]
+$\tilde{\mathsf{p}}=\begin{bmatrix}-i\eta^{-1}\,\theta\otimes\xi&\mathsf{a}\otimes 1+1\otimes\mathsf{b}\\ \mathsf{a}\otimes 1+1\otimes\mathsf{b}&-i\eta\,1\otimes 1\end{bmatrix}^{-1}.$
 It is direct to check as in (121) that we have
-$[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\theta\otimes\xi]^{-1}=(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}[-i\eta^{-1}\tilde{\mathsf{p}}](\mathbf{e}_{1}\otimes 1\otimes 1).$ (124)
+$(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\theta\otimes\xi]^{-1}=(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}[-i\eta^{-1}\tilde{\mathsf{p}}](\mathbf{e}_{1}\otimes 1\otimes 1).$ (124)
 Introducing the shorthands
@@
 and fixing a real argument $z>1$, we write this as
-[…]
+$\tilde{\mathsf{p}}$ $=\left(\tilde{\mathsf{d}}^{-1}\begin{bmatrix}-i\,1\otimes 1&\mathsf{d}_{a}\mathsf{a}\otimes\mathsf{d}_{b}+\mathsf{d}_{a}\otimes\mathsf{d}_{b}\mathsf{b}\\ \mathsf{a}\mathsf{d}_{a}\otimes\mathsf{d}_{b}+\mathsf{d}_{a}\otimes\mathsf{b}\mathsf{d}_{b}&-i\,1\otimes 1\end{bmatrix}\tilde{\mathsf{d}}^{-1}\right)^{-1}$
 $=\tilde{\mathsf{d}}\Big(\underbrace{\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*}}_{=\tilde{\mathsf{q}}}+\,i(z-1)-iz\Big)^{-1}\tilde{\mathsf{d}}$.
@@
 $\|\mathsf{d}_{a}\mathsf{a}\otimes\mathsf{d}_{b}+\mathsf{d}_{a}\otimes\mathsf{d}_{b}\mathsf{b}\|_{\mathrm{op}}\leq\|\mathsf{d}_{a}\|_{\mathrm{op}}\|\mathsf{d}_{b}\|_{\mathrm{op}}(\|\mathsf{a}\|_{\mathrm{op}}+\|\mathsf{b}\|_{\mathrm{op}})\leq 4\eta^{-1},$
-so $\tilde{\mathsf{q}}=\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*}$ is self-adjoint with spectrum contained in $[-4\eta^{-1},4\eta^{-1}]$. Hence
+so $\tilde{\mathsf{q}}=\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*}$ is self-adjoint with spectrum contained in $4\eta^{-1},4\eta^{-1}]$. Hence
 $\|\tilde{\mathsf{q}}+i(z-1)\|_{\mathrm{op}}=\max\{|\lambda+i(z-1)|:\lambda\in\mathrm{spec}(\tilde{\mathsf{q}})\}\leq\sqrt{(z-1)^{2}+16\eta^{-2}}.$ (125)
@@
 $(\tau\otimes\tau)\Big[(\mathbf{e}_{1}\otimes\mathsf{U}^{\prime}\otimes\mathsf{V}^{\prime})^{*}[\eta^{-1}\tilde{\mathsf{d}}\tilde{\mathsf{q}}^{m}\tilde{\mathsf{d}}](\mathbf{e}_{1}\otimes\mathsf{U}\otimes\mathsf{V})\Big]=\sum_{(\rho_{a},\rho_{b})\in\operatorname{NC}_{2,2}(m)}\mathrm{val}(\rho_{a},\rho_{b}).$
-A direct calculation shows that this identity holds also for $m=0$, upon defining $\operatorname{NC}_{2,2}(0)$ to have the single pair (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) with its value defined as in (34).
+A direct calculation shows that this identity holds also for $m=0$, upon defining $\operatorname{NC}_{2,2}(0)$ to have the single pair $(\emptyset,\emptyset)$ with its value defined as in (34).
 Applying this back to (126) and using $i^{m}=(-1)^{m/2}$ in $C_{m}(z)$ concludes the proof.
@@
 Applying the linearization
-$[(A\otimes I+I\otimes B)^{2}+\eta^{2}I\otimes I]^{-1}=\frac{1}{\eta}\Im[A\otimes I+I\otimes B-i\eta I\otimes I]^{-1},$
+$(A\otimes I+I\otimes B)^{2}+\eta^{2}I\otimes I]^{-1}=\frac{1}{\eta}\Im[A\otimes I+I\otimes B-i\eta I\otimes I]^{-1},$
 the analyses of [31] rested on a deterministic approximation for the resolvent
@@
 $R(z)$ $=\sum_{j,k=1}^{n}\frac{1}{\lambda_{j}+\mu_{k}-z}\mathbf{u}_{j}\mathbf{u}_{j}^{*}\otimes\mathbf{v}_{k}\mathbf{v}_{k}^{*}$
-[…]
-[…]
+$=\frac{1}{2\pi i}\oint_{\Gamma}\frac{1}{w-\lambda_{j}}\frac{1}{w+\mu_{k}-z}\mathbf{u}_{j}\mathbf{u}_{j}^{*}\otimes\mathbf{v}_{k}\mathbf{v}_{k}^{*}\,dw$
+$={-}\frac{1}{2\pi i}\oint_{\Gamma}R_{A}(w)\otimes R_{B}(z-w)\,dw$
 where $R_{A}(z)=(A-zI)^{-1}$ and $R_{B}(z)=(B-zI)^{-1}$.
@@
 Lemma B.2 (Decoupling).
-[…]
+Let $(\alpha_{i})_{i=1}^{n}$ be a sequence of independent $\mathbb{C}$-valued random variables, let $(\alpha_{i}^{\prime})_{i=1}^{n}$ be an independent copy of $(\alpha_{i})_{i=1}^{n}$, and let $(\mathsf{y}_{ij}:i,j=1,\ldots,n)$ be elements of a Banach space with norm $\|\cdot\|$.
+Then for a universal constant $C>0$,
 $\mathbb{E}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg\|^{p}\bigg]\leq C^{p+1}\mathbb{E}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg\|^{p}\bigg]$
 Proof.
-By [15, Theorem 1], for a universal constant $C>0$,
-[…]
-[…]
+By [16, Theorem 1], for a universal constant $C>0$,
+$\frac{1}{p}\,\mathbb{E}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg\|^{p}\bigg]$ $=\int_{0}^{\infty}t^{p-1}\mathbb{P}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{j}\mathsf{y}_{ij}\bigg\|\geq t\bigg]\;dt$
+$\leq\int_{0}^{\infty}Ct^{p-1}\mathbb{P}\bigg[C\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg\|\geq t\bigg]\;dt=\frac{C^{p+1}}{p}\mathbb{E}\bigg[\bigg\|\sum_{i\neq j}\alpha_{i}\alpha_{i}^{\prime}\mathsf{y}_{ij}\bigg\|^{p}\bigg].$
 Proof of Lemma 3.4.
@@
 $T(E_{ij}\otimes\mathsf{x})=\begin{cases}E_{i,j+i-1}\otimes\mathsf{x}&\text{if $j=1$}\\ 0&\text{otherwise}\end{cases}$
-where $j+i-1$ is interpreted modulo $n$.
-Thus $\hat{\mathsf{x}}$ has $\mathcal{X}$-valued entries $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}$ along the first column, and $T(\hat{\mathsf{x}})$ has these entries instead along the main diagonal.
-We have
+[…]
 $\|\hat{\mathsf{x}}\|_{p}^{p}$ $=\|(\hat{\mathsf{x}}^{*}\hat{\mathsf{x}})^{1/2}\|_{p}^{p}=\frac{1}{n}\bigg\|\bigg(\sum_{i=1}^{n}\mathsf{x}_{i}^{*}\mathsf{x}_{i}\bigg)^{1/2}\bigg\|_{p}^{p},$
@@
 Appendix C Fluctuation averaging
-In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [30, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [30].
+In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [29, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [29].
 Lemma C.1 ([29], Lemma A.3).
@@
 By scale invariance of the statement of the lemma, let us assume without loss of generality that $\|\mathbf{u}\|_{\infty}=n^{-1}$.
-Applying this and the bound |$|\{\mathbf{i}:T(\mathbf{i})=t\}|\leq C_{l}^{\prime}n^{t+(l-t)/2}$ for a constant $C_{l}^{\prime}>0$, we get (for different constants $C_{l},C_{l}^{\prime}>0$)
+Applying this and the bound $|\{\mathbf{i}:T(\mathbf{i})=t\}|\leq C_{l}^{\prime}n^{t+(l-t)/2}$ for a constant $C_{l}^{\prime}>0$, we get (for different constants $C_{l},C_{l}^{\prime}>0$)
 $\mathbb{E}\left\|\sum_{i=1}^{n}u_{i}\mathsf{x}_{i}\right\|_{p}^{l}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{t=1}^{l}(n^{-\beta})^{t}(n^{-1/2})^{l-t}\leq C_{l}n^{-\alpha l+\epsilon}\sum_{t=1}^{l}(n^{-\beta}+n^{-1/2})^{l}\leq C_{l}^{\prime}n^{-(\alpha+\beta^{\prime})l+\epsilon}$
@@
 We collect here several pieces of background on von Neumann algebras and non-commutative $L^{p}$ spaces that are needed in our main arguments.
-We refer to [50], [14, Section 1], and [58, Chapter 14] for additional discussion.
+We refer to [50], [15, Section 1], and [58, Chapter 14] for additional discussion.
 Throughout, $\mathcal{X}$ is a (finite) von Neumann algebra with faithful, normal, tracial state $\phi$ as in Section 3.
@@
 We set $\|\mathsf{x}\|_{\infty}\equiv\|\mathsf{x}\|_{\mathrm{op}}$ and $L^{\infty}(\mathcal{X})\equiv\mathcal{X}$.
-These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [50] or [14, Section 1] for this construction.
+These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [50] or [15, Section 1] for this construction.
 Lemma D.2 (Non-commutative $L^{p}$-spaces. [58] Theorem 14.1, [15] Proposition 1.1).
 For each $p\in[1,\infty)$, $\|\mathsf{x}\|_{p}=\phi(|x|^{p})^{1/p}$ defines a complete norm on $L^{p}(\mathcal{X})$, satisfying
-$\|\mathsf{x}\|_{p}\leq\|\mathsf{y}\|_{p}\text{ for all }\mathsf{x},\mathsf{y}\in L^{p}(\mathcal{X})\text{ with }0\leq\mathsf{x}\leq\mathsf{y},\qquad|\phi(x)|\leq\|\mathsf{x}\|_{1}\text{ for all }\mathsf{x}\in L^{1}(\mathcal{X}).$
+[…]
   1. (a)